### v1

# LOLA: LLM-Assisted Online Learning Algorithm for Content Experiments

Abstract
In the rapidly evolving digital content landscape, media firms and news publishers require automated and efficient methods to enhance user engagement. This paper introduces the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery. Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests aimed at evaluating the performance of various headlines associated with the same article content, we first investigate three broad pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs. Our findings indicate that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline among two options. In contrast, OpenAI-embedding-based classification models and fine-tuned Llama-3-8b models achieve comparable accuracy, around 82-84%, though still falling short of the performance of experimentation with sufficient traffic. We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to adaptively allocate traffic and maximize clicks. Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B testing method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic or numerous arms. Our approach is both scalable and broadly applicable to content experiments across a variety of digital settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
Keywords: LLMs, Content Experiments, Bandits, News
1 Introduction
Digital content consumption has seen unprecedented growth, leading to a proliferation of content across various platforms. In today’s real-time digital environment, media firms and news publishers need automated and efficient methods to determine which content generates high user engagement and platform growth. This includes identifying the most appealing articles, the most attractive headlines, and the catchiest cover images (NYTHeadline2017; NYTimes2019). Traditionally, media firms and publishers have relied on experimentation-based approaches to address this problem. Broadly speaking, there are two general types of experimentation styles—(1) Standard A/B tests and (2) Online learning algorithms or bandits. [^1] We use the terms adaptive experimentation, online learning algorithm, and bandits interchangeably throughout the paper. We now discuss each briefly.
A/B testing is the most straightforward experimentation method, where a firm allocates a fixed portion of traffic to different treatment arms, assesses the results, and then goes with the best-performing arm. This is also known as the Explore and Commit strategy, or E&C, where the firm explores for a fixed period using A/B tests and then commits to one treatment based on the results of those tests. This approach is widely used in the publishing industry. For instance, The New York Times recently built a centralized internal A/B test platform, ABRA (A/B Reporting and Allocation architecture), allowing different teams to experiment with their content; see (NYTimes2024) for more details. Another example is Upworthy, which has implemented extensive A/B testing to refine its headlines (matias2021upworthy). Upworthy editors create several headlines for the same article and randomly assign readers to see different headlines with equal probability for a fixed amount of traffic, and then show the best-performing headline to all the users. The main advantage of this approach is that it is trustworthy—with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased and accurate. However, a major drawback of this approach is the wastage of traffic—by assigning equal traffic to all the treatment arms during the exploration phase (including the poorly performing ones), the firm incurs high regret or low profits. This is especially problematic in the news industry since news articles tend to have short lifetimes and become stale quickly (typically within a day or two). Therefore, if a firm wastes a lot of traffic to learn the best article/headline, the article itself might become irrelevant by the end of the A/B test.
The second approach, online learning algorithms, or bandits, are able to address some of these problems. These algorithms dynamically adjust traffic allocation to articles/headlines based on their features, past performance, and associated uncertainty. This method provides a more efficient way to optimize content delivery by continually learning and adapting. For instance, both Yahoo News and The New York Times have successfully implemented contextual bandit algorithms to identify content that maximizes user engagement. Yahoo was one of the earliest platforms to use the linear contextual bandit for news article recommendation (li2010contextual). More recently, The New York Times has also adopted a similar algorithm to power its article recommendation system (NYTimes2019).
Online learning algorithms typically outperform A/B tests in terms of overall regret or reward. These algorithms reduce the extent of exploration over time by moving traffic away from poorly performing treatments dynamically. As a result, they tend to switch to the best-performing arms quickly and incur lower regret compared to standard E&C strategies. This has been established both theoretically (lattimore2020bandit) as well empirically. For example, (li2010contextual) demonstrate that the linear contextual bandit algorithm improves the total clicks on news articles over the random policy (i.e., A/B testing with equal probability for arms) at Yahoo News using off-policy evaluation on field data. [^2] Similarly, in the context of digital advertising, (schwartz2017customer) use field experiments to demonstrate that adaptive algorithms outperform the standard balanced A/B test.
Nevertheless, these online algorithms have notable drawbacks. In particular, in standard content experiments, these algorithms typically suffer from a cold-start problem. That is, firms typically start these adaptive experiments under the assumption that all have an equal probability of being the best-performing one. This can lead to a situation where more-than-necessary traffic is allocated to sub-optimal arms and higher regret as less appealing content is shown to users during the exploration phase. Additionally, due to their probabilistic nature, these algorithms can inadvertently converge on a suboptimal arm, especially when the experimental traffic is small, increasing the probability of identifying a suboptimal arm as the best one. These issues arise because online algorithms are designed for large traffic regimes, where the experimenter does not care too much about initialization since it has little impact on the asymptotic regret. However, as discussed earlier, in the news industry, there is limited traffic available for experimentation, given that news content becomes stale quickly. Therefore, the negative effects of the “equally effective” initialization do not diminish sufficiently during the time horizon of interest, leading to higher regret.
Given these issues with content experimentation, the main question that this paper asks and answers is whether we can leverage the recently developed Large Language Models (LLMs) to enhance current content experimentation practices and simplify the process of identifying appealing content. LLMs, trained on extensive human-generated data, have demonstrated the ability to mimic human preferences and behavior in various contexts. For example, (li2024frontiers) investigate whether LLM-generated survey responses produce perceptual maps similar to those generated by human survey responses in car branding analysis. Their study shows a high agreement rate of over 75%. More specifically, in the context of news articles, (yoganarasimhan2024feeds) show that LLMs are effective at understanding the polarization of news content. Hence, a relevant question is whether news publishers can directly use LLMs to predict the appeal of different content, potentially replacing traditional experimentation methods. This approach could significantly reduce the costs associated with experimentation while simplifying the content delivery processes.
To address this question, we leverage a large-scale dataset on content experiments by the publishing firm Upworthy conducted in the 2013–2015 time period. The dataset consists of 17,618 headline tests spanning 77,245 headlines, over 277 million impressions, and over 3.7 million clicks. The goal of each A/B test was to identify the best headline among a set of candidate headlines (for a given article). This dataset is ideal for our study since it was collected from A/B tests where each headline received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance. Such a high-quality dataset provides a trustworthy ground for comparing different approaches. Additionally, since the dataset was sourced from the real media industry, our methods can reflect real-world scenarios and offer practical guidance to both researchers and practitioners.
In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can predict headline attractiveness accurately. For this exercise, we focus on pairs of headlines (within a given A/B test) whose CTRs are significantly different (based on the observed data). This gives us a total of 39,158 pairs of headlines and allows us to recast the prediction task as a binary classification task, where a random guess gives 50% accuracy, and the perfect classifier reaches 100% accuracy. Thus, this serves as a preliminary testbed to evaluate whether pure-LLM-based approaches can predict which of the two headlines is catchier (or clickable).
We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM-based text-embedding approaches, and (3) Fine-tuning approaches. We discuss each approach and its performance next. First, for prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning. Zero-shot prompting is a technique in which an LLM generates responses without being explicitly trained on specific examples. This is the most common way in which most users interact with the LLM. In-context learning involves providing the LLM with several demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation). We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches, with accuracy hovering around 50%. GPT-4 performs marginally better—it achieves an accuracy of 57.26% for zero-shot prompting and 60-65% for in-context learning. Overall, this suggests that prompt-based approaches cannot aid/replace content experiments in any meaningful way.
Next, we turn to the second approach: text embeddings. Specifically, we transform headlines into embedding vectors using the most recent embedding model from OpenAI. We then concatenate two embeddings as inputs, along with the winner with higher CTR as the output, to train binary classification models, including logistic regression and multilayer perceptron (a type of neural network). We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82% accuracy, while more complex neural networks do not improve performance over the simple logistic model. Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora). This approach performs the best, with 83% accuracy, which is a small improvement over the embedding-based approach. [^3] On the other hand, it also marginally increases training costs and engineering complexity. Overall, we find that prompt-based approaches, which have been the standard method in recent marketing literature for using LLMs so far, perform poorly. In contrast, models based on LLM embeddings or fine-tuned open-source LLMs offer significantly more accurate predictions. The primary advantage of LLM-based methods is that predictions can be made without wasting any traffic. Nonetheless, even the best LLM-based models do not match the accuracy of experimentation-based approaches (with sufficient traffic).
Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm). Our key insight is to use predictions from fine-tuned LLM models as priors in online algorithms to optimize experimentation. The approach involves two steps. First, we train an LLM-based CTR prediction model, which can be an LLM-embedding-based model or a fine-tuned open-source LLM model, or indeed any other LLM-based approach. The second step integrates these predictions into online learning algorithms. LOLA is a general framework, and depending on the specific goals of the experimentation, it can be adapted for different prediction models in the first step and different online algorithms in the second step. For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs), detailed in Section 4.1. The idea of LLM-2UCBs is intuitive: we can view LLM-based CTR predictions as auxiliary samples before the start of the online algorithm (e.g., a 1% CTR can be viewed as 100 impressions with 1 click or 1,000 impressions with 10 clicks). After fine-tuning this auxiliary impression size hyperparameter, we incorporate these imaginary samples into the standard UCB algorithm to get the reward estimator and shrink the upper confidence bound.
LOLA combines the advantages of both experimentation and LLMs, making it accurate and efficient. By using LLMs, we reduce the waste of impressions on poor-performing headlines, especially in the initial stage when we face a cold-start problem and lack good prior information about the headlines’ performance. On the other hand, through experimentation, we can correct the errors in LLMs. Even though LLMs achieve around 82% accuracy, there are still 18% of incorrect cases. This hybrid approach ensures that while LLMs provide reasonable initial predictions, ongoing experimentation refines and improves overall performance, making LOLA an effective solution for content optimization.
We present a comprehensive evaluation of LOLA and compare its performance with three natural benchmarks using the Upworthy dataset: (1) Explore and Commit, (2) pure online learning algorithm, and (3) pure LLM-based model. The first benchmark, E&C, was the status quo at Upworthy during the time of data collection, wherein the editors first ran an A/B test on a set of headlines for a fixed set of impressions, and then displayed the winner for the rest of the traffic. For the pure online learning algorithm, we use the standard Upper Confidence Bound (UCB) algorithm, which is initialized with uniform CTRs. For the pure LLM-based approach we use the linear CTR prediction model in conjunction with OpenAI embeddings. We compute CTRs for all headlines and select the one with the highest CTR for all impressions, bypassing the experimentation phase. [^4] We chose the embedding-based approach for the evaluation instead of the fine-tuned model, given its ease of use and almost comparable performance to the fine-tuned open-source LLM. In LOLA, we use the same CTR prediction model as in the pure LLM approach and employ LLM-2UCBs to maximize accumulated clicks. Our numerical comparison primarily considers the regret minimization framework, as minimizing regret in Upworthy’s context equates to maximizing total clicks over all impressions, including those for experimentation. This metric is pertinent to business profitability since Upworthy, as a free online publisher without subscription fees, relies heavily on advertising revenue. Increased user clicks enable Upworthy to display more advertisements, thereby increasing revenue.
In our empirical evaluation, we find that LOLA outperforms all the baseline approaches. We find that for the earlier impressions, both LOLA and the pure LLM-based approach beat the pure experimentation-based approaches (i.e., the standard A/B test and UCB algorithm) due to relatively accurate LLM-based CTR predictions. However, as the number of impressions grows, the performance of the pure LLM-based approach declines rapidly, as it does not learn anything from past impressions and, as such, cannot improve its performance. Thus, its accuracy falls short compared to other experimentation-based methods. Thus, as discussed earlier, our experiments demonstrate that LOLA is able to leverage the power of LLMs early in the horizon and then build on it to take advantage of experimentation over time, thereby bringing together the strengths of both methods.
Finally, we observe that the performance gap between LOLA and the standard UCB narrows, given that both methods are asymptotically optimal and the benefit provided by LLM diminishes over time. Thus, we expect that LOLA will be especially valuable in settings where the traffic available for experimentation is limited and/or the number of arms is large. Nevertheless, LOLA continues to outperform standard UCB, even after very long horizons.
In summary, our paper makes two key contributions to the literature. First, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings. To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques. We demonstrate the value of our approach using a large-scale dataset from Upworthy, showcasing significant improvements in total clicks and profitability. Second, from a managerial perspective, LOLA can be used in various settings where firms need to decide which content to show, such as news articles, headlines, or even advertisements and website content. This method is based on open-source LLM models, making it cost-effective and easily deployable across different domains once fine-tuned on relevant datasets. Compared to the substantial benefits gained from improved content optimization, the cost of fine-tuning these models is negligible. Because this fine-tuning process on the open-sourced LLMs can be efficiently executed even on entry-level GPUs, making it accessible for a wide range of applications. This accessibility ensures that even firms with limited budgets can implement advanced LLM-based techniques to enhance their content strategy. More broadly, while we focused mainly on the news/publishing industry, our framework is general and can easily extend to other settings where firms need to evaluate which type of content maximizes user engagement, including online/digital advertising and social media platforms.
2 Upworthy Data and Experiments
We now discuss data sourced from Upworthy, a U.S. media publisher known for its innovative use of A/B test in online media publishing. In a bid to increase clicks to articles, Upworthy conducted randomized experiments with each article published, exploring different combinations of headlines and images to determine which elements most effectively engaged viewers. The archival record from January 24, 2013, to April 30, 2015, detailed in (matias2021upworthy), demonstrates how strategic packaging of headlines, subheadings, and images played a crucial role in Upworthy’s growth strategy. Such A/B testing practices are also widely used in the media and publishing industry, technology industry, and political campaigns. A/B testing allows firms to assess and refine the impact of various editorial and content choices (e.g., headlines or text of an article/email) systematically and automatically.
Test ID Headline Impressions Clicks
1 New York’s Last Chance To Preserve Its Water Supply 2,675 15
1 How YOU Can Help New York Stay Un-Fracked In Under 5 Minutes 2,639 19
1 Why Yoko Ono Is The Only Thing Standing Between New York And Catastrophic Gas Fracking 2,734 34
2 If You Know Anyone Who Is Afraid Of Gay People, Here’s A Cartoon That Will Ease Them Back To Reality 4,155 120
2 Hey Dude. If You Have An Older Brother, There’s A Bigger Chance You’re Gay 4,080 41
Table 1: Samples of Upworthy experiments’ results for headlines. We display the key columns for our analysis, including the ID for the A/B tests, the text of headlines, and the number of impressions and clicks in the test.
During the experiment period, Upworthy’s editorial team created several versions of headlines and/or images for each article (internally called “package” by Upworthy). A package is defined as one treatment or arm for an article and consists of a headline, image, or a combination of both. Editors would first choose several of what they believed were the most promising packages to be tested. Then, they A/B tested these packages or treatments to identify which one resonated the best with their audience. Specifically, Upworthy’s A/B testing system recorded how many impressions and clicks each package received. Key metrics from these tests were shown via a dashboard, which guided decisions on whether to adjust the content further or finalize a particular version. Table 1 shows two examples of A/B tests, with the relevant columns. The full details of all the A/B tests and their results are available from the Upworthy Research Archive.
We now describe the details of this dataset and discuss our pre-processing procedure. In the original dataset, there are 150,817 tested packages from 32,487 deployed A/B tests. The dataset records a total of 538,272,878 impressions and 8,182,674 clicks. All these statistics are also listed in the top panel of Table 2. These summary statistics indicate that during this period, each A/B test had an average of 4.64 packages, with each package receiving an average of 3,569 impressions and 54.26 clicks, resulting in an average click-through rate (CTR) of 1.52%. Within an A/B test, each package/treatment arm had the same probability of receiving an impression; so the number of impressions received by all the packages within a test is approximately the same. However, the actual number of impressions varies across tests, with the first quartile at 2,745, the median at 3,117, and the third quartile at 4,089. This is due to the relatively straightforward implementation of A/B tests, i.e., Upworthy did not conduct any power analysis in advance to determine the traffic for a given test, as confirmed by (matias2021upworthy).
Original data # of tests 32,487
# of packages 150,817
# of impression in tests 538,272,878
# of clicks in tests 8,182,674
Headline tests data # of tests 17,681
# of packages 77,245
# of impression in tests 277,338,713
# of clicks in tests 3,741,517
# of headline pairs 140,655
# of headline pairs with different CTRs at 0.05 significant level 39,158
# of headline pairs with different CTRs at 0.01 significant level 23,682
# of impressions (both test and non-test impressions) 2,351,171,402
Table 2: Basic summary statistics of Upworthy data during the time window between January 24, 2013 and April 30, 2015
Among these tests, some were conducted to test different images. However, the dataset does not allow us to trace back the actual images used in the tests. Therefore, we focus on the headline tests and filter out all the image tests. After this filtering, there are 17,681 tests comparing headlines, totaling 77,245 packages. On average, each article was tested with 4.37 headlines. Note that after filtering, each package or treatment simply refers to a unique headline; and therefore, we use the phrases package and headline interchangeably in the rest of the paper. The distribution of the number of packages/headlines for each article is shown in Table 3. After filtering, the dataset includes 277,338,713 impressions and 3,741,517 clicks; see the middle panel of Table 2. We also report the total number of impressions for the entire Upworthy website, as recorded in their data from Google Analytics. It is noteworthy that the impressions attributable to the tests constitute only 22.9% of the total traffic. This relatively small proportion of traffic can be attributed to Upworthy’s policy of conducting no more than one test per webpage to avoid interference effects.
# of headlines in one test # of tests % of samples
2 1,619 9.16
3 939 5.31
4 8,836 49.97
5 2,964 16.76
6 2,685 15.19
7 or more 638 3.61
Table 3: Distribution of number of headlines tested, across A/B tests.
To examine whether these A/B tests were informative, or not, we generate pairs of headlines and conduct pairwise t-tests on the CTRs of the headlines. For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^5] It is important to note that the issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives. In our dataset, tests involving 10 pairs or less (equivalent to no more than 5 packages) account for over 80% of all tests. The risk of multiple testing is relatively low when conducting fewer than 10 tests, and conventional significance levels, e.g., $\alpha=0.05$ are generally adequate to account for this risk. To ensure the robustness of our findings, we conducted an additional analysis by excluding tests with more than 5 packages. The comparisons among pure-LLM-based methods remained consistent because all methods were conducted on the same dataset. It is important to note that a higher risk of multiple testing results in a larger number of actually insignificant pairs within our significant subset, which may only decrease the prediction accuracy. In other words, the results we present using the significant pairs may underestimate the prediction accuracy. Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in Section 4, we utilize all available data without concern for multiple testing. see Table 4 for examples of such tests for the two A/B tests highlighted earlier in Table 1. After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level, and 23,682 at a 0.01 level. We do not observe a fundamental difference when conducting our experiments with 0.05 or 0.01 significant level pairs; therefore, for simplicity, we primarily use 0.05 level data in the main analysis in Section 3. In summary, we find that only 28% of the tests were significant at the 0.05 level, and a mere 17% are significant at the 0.01 level. Overall, these results suggest that most pairs of headlines tend to perform similarly, and the task of learning which headline is catchier is non-trivial.
Test ID Headline 1 Headline 2 CTR 1 CTR 2 p value Winner
1 How YOU Can Help… Why Yoko Ono Is… 0.0071 0.0124 0.054 2
1 Why Yoko Ono Is… New York’s Last… 0.0124 0.0056 0.009 1
1 How YOU Can Help… New York’s Last… 0.0071 0.0056 0.496 1
2 If You Know Anyone Who… Hey Dude. If… 0.0289 0.010 0.000 1
Table 4: Example of pairwise comparisons for two A/B tests from the Upworthy data.
3 Pure-LLM-Based Methods
In this section, we analyze 39,158 headline pairs, representing 28.05% of the total, at a 0.05 significance level. We focus on these significant pairs for three reasons: First, insignificant pairs suggest that decisions in final headline selection do not substantially affect users’ click behavior, implying that accuracy in predicting outcomes for “equally good” samples is not critically important from a business perspective. Second, to avoid information overload in our survey study, we limit our focus to significant pairs for a fair comparison between LLMs and humans. Third, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in Section 4. As such, relative comparison across methods is more important as long as we use the same dataset for all methods. Furthermore, we note that the methods discussed in this section can be readily adapted from binary to multi-class classification tasks. Indeed, in Section 4, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
In the rest of this section, we investigate the performance of three widely used LLM-based approaches—(1) GPT prompt-based approaches, including zero-shot prompting and in-context learning, (2) embedding-based classification models, and (3) fine-tuning of LLMs.
3.1 Prompt-based Approaches
In this section, we consider prompt-based approaches to evaluate which headline among a pair of headlines is catchier. These approaches simply use GPT prompts to learn which headline is catchier, and there is no further modeling or explicit fine-tuning involved. Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far .
We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning. We discuss both in detail below.
3.1.1 Zero-shot Prompting
Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly trained on specific examples. In this study, we evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models. We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses. The key parameters include specifying which GPT model to use (such as GPT-3 or GPT-4) and the system parameters, which guide the type of responses generated. The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation. Additionally, we set the default temperature parameter of 1.666The temperature parameter scales and controls the randomness of the model’s responses. GPT model generates responses token by token. Tokens are common sequences of characters found in a set of text. A helpful rule of thumb is that one token generally corresponds to about four characters of text in common English, translating to roughly three-quarters of a word (so 100 tokens $\approx$ 75 words), with each token sampled from a token distribution provided by the model. The temperature parameter determines whether this distribution is flattened or sharpened. A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused. For simplicity, we use the default setting with temperature as 1.0. We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy/performance.
Figure 1: Zero-Shot Prompting for Headline Selection.
In our setting, we design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1. We test this prompt using three OpenAI large language models, [^7] As recommended by OpenAI, see this page, accessed on May 24, 2024, one can default to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o. OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence., each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters, enhancing its capabilities. The goal here is to investigate whether improvements in model architecture and parameter size can improve the accuracy of our predictions via standard prompts.
3.1.2 In-context Learning
In-context learning is a technique where a model is provided with demonstrations within the input prompt to help it perform a task. Demonstrations are sample inputs and outputs that illustrate the desired behavior or response. This method does not change the underlying parameters of the GPT models; rather, it leverages the model’s ability to recognize patterns and apply them to new, similar tasks by conditioning on input-output examples. Previous studies have shown that in-context learning can significantly boost a model’s accuracy and relevance in responses compared to zero-shot prompting. (SangMichaelXieandSewonMin_2022) explains this using a Bayesian inference framework, suggesting that in-context learning involves inferring latent concepts from pre-training data, utilizing all components of the prompt (inputs, outputs, formatting) to locate these concepts, even when training examples have random outputs. (xie2021explanation) further elucidate that in-context learning can emerge from models pre-trained on documents with long-range coherence, requiring the inference of document-level concepts to generate coherent text. They provide empirical evidence and theoretical proofs showing that both Transformers and LSTMs can exhibit in-context learning on synthetic datasets. [^8] More generally, beyond LLMs, (garg2022can) investigate the capacity of Transformers to in-context learn various function classes, demonstrating that these models can be trained from scratch to learn linear, sparse linear, two-layer neural networks, and decision trees, even under distribution shifts between training data and inference-time prompts.
We design the in-context learning prompt shown in Figure 2. Our prompts include demonstrations of catchier headlines and ask the model to determine which headline is catchier for a subsequent pair of headlines. We use the headline with the higher CTR from the test, as indicated in the “Winner” column of Table 4, as the correct answer in our demonstrations. A few demonstrations are typically sufficient to guide the model. Improvement tends to have diminishing returns quickly; adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking). In our numerical experiments, we use two or five demonstrations to investigate the impact of the number of examples on performance. [^9] We did run experiments with a larger number of demonstrations, but did not find any significant improvements from doing so; and hence do not report them here.
Figure 2: In-Context Learning Prompt for Headline Selection.
An interesting aspect of in-context learning is the model’s robustness to incorrect labels in the demonstrations. (min2022rethinking) show that randomized labels in demonstrations barely affect performance and highlight other critical aspects like the label space, input text distribution, and sequence format. This is likely because the model relies more on the distribution of input text, label space, and format rather than specific input-label mappings. That is, in-context learning seems to benefit from recognizing general patterns and structures in the data, which the model infers from its pre-training rather than the actual labels. To examine if this is the case in our setting as well, we also consider experiments where we intentionally flip the labels in the demonstrations provided to the model.
3.1.3 Performance Analysis of Prompt-based Approaches
Before presenting the results, we make note of two points. First, we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity. Second, the GPT models can produce different answers even when given the exact same prompt. To mitigate this variability, we randomly select 1,000 significant pairs and conduct all our experiments on this same test set. Similarly, all the demonstrations are also randomly sampled and fixed for all experiments to reduce the variance of prompt outputs.
The results of all the prompt-based experiments are shown in Table 5. A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix A.1. The second column of 5, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting. Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with flipped outcome labels. The third column, “Flipped Label” indicates whether the labels were flipped or not.
GPT Model $n_{demo}$ Flipped Label Accuracy [0.025 0.975]
GPT-3.5 0 0 54.20% 51.10% 57.10%
GPT-3.5 2 0 48.70% 45.50% 51.50%
GPT-3.5 2 1 50.85% 47.60% 53.70%
GPT-3.5 5 0 51.30% 48.20% 54.30%
GPT-3.5 5 1 50.10% 46.80% 53.20%
GPT-4 0 0 57.26% 54.20% 60.00%
GPT-4 2 0 63.60% 60.50% 66.40%
GPT-4 2 1 62.90% 59.70% 65.70%
GPT-4 5 0 60.80% 57.70% 63.40%
GPT-4 5 1 61.80% 58.50% 64.80%
Table 5: Results for prompt-based approaches with different GPT models, the number of demonstrations, and whether the demo’s label is flipped or not. Accuracy is defined as the ratio of correct answers divided by the number of prompts (1000 in our analysis). The 95% confidence interval is calculated by performing 5,000 bootstrap resamples of the entire 1,000 pairs. To save on the monetary cost of prompts and not introduce additional variance, we did not rerun the GPT prompts for the same samples; instead, we used the responses/predictions from the original run.
We find that prompt-based approaches using GPT-3.5, except for the zero-shot prompting, are not significantly better than random predictions (accuracy rates hover around 50%). Neither the number of demonstrations used nor the label type have any significant impact on the accuracy of the GPT-3.5 predictions. Interestingly, we find zero-shot prompting using GPT-3.5 is significantly better than the random prediction, although by only 4.20%, and marginally outperforms the in-context learning. In comparison, experiments with GPT-4 have accuracy rates around 60-65%, which is a significant improvement in performance compared to GPT-3.5. For GPT-4, in-context learning significantly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance. There is also no significant difference between using correct labels and incorrect labels in the demonstrations for in-context learning. This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
Finally, we also report the monetary cost and total run-time of the prompt-based experiments. The cost is \$1.13 for GPT-3.5 with 1,000 pairs and \$20.11 for GPT-4, which is approximately 20 times the cost of GPT-3.5. The total run-time is 121.5 minutes for the entire experiment, which involved testing a total of 10,000 prompts. [^10] Note that the run-time may depend on the API usage tiers, which provide different rate limits.
In summary, we find that prompt-based methods are relatively slow (compared to other methods introduced later in this section), have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), and high monetary cost. All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
3.2 Classification Models with LLM-based Text Embeddings
In this section, we consider another LLM-based method for this prediction task. Specifically, we leverage the text embeddings from OpenAI and utilize classification models, including Logistic Regression (Logit) and Multilayer Perceptron (MLP), to predict which of the headlines in a pair is likely to generate a higher CTR. The Logit model is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables. On the other hand, MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data. We refer interested readers to Chapter 5 in (zhang2021dive) for more details.
3.2.1 Text Embedding Technique
Text embedding is an advanced technique in natural language processing (NLP) that transforms larger chunks of text, such as sentences, paragraphs, or documents, into numerical vectors. Such vector forms a representation of the semantic information of the input text, and can be further utilized by downstream machine learning models.
Models like BERT (devlin2018bert), Sentence-BERT (reimers2019sentence), and OpenAI’s embedding models (openai-embedding) (text-embedding-3-large for example) are viewed as state-of-the-art embedding models at the date of their release. Essentially, text embedding is a numerical representation of text that captures the rich information in the text. These embedding vectors are used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which determines the emotional tone of a piece of text.
It is important to distinguish between embedding models and standard LLMs used for prompts, e.g., GPT. While they are based on similar underlying Transformer architectures and extensive pretraining on large text corpora, they are optimized for different tasks. Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts, while embedding models focus on producing high-quality embeddings for downstream tasks. For a more detailed survey of the most recent text embedding models (patil2023survey).
In the remainder of this section, we leverage the embedding model from OpenAI to convert headline texts into numerical embedding vectors, which are then used to train the downstream binary classification models to identify which headline is better.
3.2.2 Implementation Details
In our numerical study, we employ the embedding model from OpenAI to facilitate a fair comparison with GPT-based prompts. At the time of this study, the latest and best-performing embedding model available was text-embedding-3-large. This model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021. A notable feature of this model is its flexibility, allowing users to tailor the embedding dimension to their specific needs. [^11] Other models capable of text embedding, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained. These dimensions cannot be changed or resized for different applications. However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka). This allows a single embedding to adapt to the computational constraints of downstream tasks. More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used. In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream classification models. [^12] A higher dimension does not necessarily mean better performance. While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples. This sparsity can increase the possibility of over-fitting.
Figure 3 illustrates the conceptual framework of our approach here. First, we call the OpenAI embedding API in Python to obtain text embeddings for the headlines based on text-embedding-3-large. We then concatenate the embeddings of two headline embeddings and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR. To assess the impact of model complexity on performance, we test two classification models: Logit and MLP. The embedding model parameters remain fixed during training, while only the classification model parameters are updated.
Figure 3: The pipeline of the classification with the LLM-based text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
As discussed earlier, we chose all headline pairs that are significant at 0.05 level for our analysis. It is worth noting that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment. To avoid information leakage, we first randomly split the data into 80% training and 20% test sets based on the test ID (the experiment ID for Upworthy) to ensure all headline pairs under the same experiment are either in the training set or the test set. Notice that Upworthy allowed additional experiments for the same headline under a new experiment ID because the editor may require an extra test to make an informed decision. It means that we may have the same headlines in both the training and test sets, even if they were in totally different experiments. In such cases, we discard headlines that appeared in the training set from the test set to avoid information leakage. [^13] There are alternative ways to do the train-test split without discarding any samples. One approach is assigning one headline pair at one time to training or test groups and then adding all pairs with the same headline in the same group until all samples are allocated. However, in our study, after conducting several experiments, we found that the splitting procedure has minimal impact on the final results. Hence, to maintain clarity and simplicity in our exposition, we simply discard duplicated testing samples, which constitute around 25% of the original test set.
One may still be concerned if this Upworthy data has been used as part of OpenAI’s training corpus, which could potentially invalidate our analysis. However, we argue this is unlikely for several reasons. First, it is improbable that OpenAI would have specifically downloaded this Upworthy csv format dataset from a subfolder in Upworthy Research Archive. Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than paired headlines with winner information. Therefore, during the training of OpenAI models, it is unlikely they were exposed to the winner information since the training objective is to predict the next word. Third, we conducted an experiment to determine if this dataset was indeed a part of OpenAI’s corpus. We used embeddings from two headlines with significantly different impressions (possibly from different tests) for the downstream task of predicting which headline received higher impressions. If the embedding model was trained on this corpus, its embeddings should reflect the number of impressions and perform well on this task. However, the observed accuracy was approximately 50%, indicating that the embeddings do not contain such information, confirming our assumption.
Finally, as discussed earlier, we consider both a Logit model (with linear features) and a MLP for the classification task (i.e., to predict which headline is catchier). The MLP model employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting. The MLP network is trained using the stochastic gradient descent algorithm to minimize binary cross-entropy. We perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration. For the OpenAI-256E model, we used the following MLP architecture: Inputs(512)-fc(256)-ReLU-dropout(0.5)-fc(1)-sigmoid, and the training hyperparameters were set with an $L_{2}$ penalty of $10^{-4}$ and 74 training epochs. For the OpenAI-3072E model, the MLP architecture was Inputs(6144)-fc(2048)-ReLU-dropout(0.5)-fc(1)-sigmoid, with an $L_{2}$ penalty of $10^{-4}$ and 32 training epochs. Please see Web Appendix A.2 for additional details of this model.
3.2.3 Performance Analysis of Embedding-based Approaches
We now present the performance of the different embedding-based classification models in Table 6. We also include a benchmark using logistic regression with Word2Vec word embeddings for a comprehensive comparison. We use Word2Vec as it was one of the best embedding models before the wide adoption of Transformer in NLP. Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information. We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec. To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
Embedding Model Classification Model Accuracy
OpenAI-256E Logit 77.56%
MLP 78.29%
OpenAI-3072E Logit 82.33%
MLP 82.06%
Word2Vec-256E Logit 68.58%
Word2Vec-3072E Logit 68.84%
Table 6: The performance of different embedding and classification models on test data. All models use the same training data, and accuracy is evaluated on the test data. Note that Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents. Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average of the accuracy from three reruns: 78.27%, 78.29% and 78.32% for OpenAI-256E + MLP and 82.43%, 81.67% and 82.09% for OpenAI-3072E + MLP.
We find that classification models using the 3072-dimensional text embedding from OpenAI achieve more than 80% accuracy, demonstrating superior performance compared to prompt-based approaches. This performance can be attributed to the fact that embedding-based classification models are trained on a customized dataset, allowing it to learn decision rules tailored precisely fine-tuned to the headline selection task. Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust. In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations, resulting in sub-optimal performance compared to a task-specific trained classifier. Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus. Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
Next, we discuss the detailed comparison between models using the embedding technique. We first compare the results under different embedding dimensions. Typically, more dimensions give greater quality encoding and, thus, higher prediction accuracy. However, there is some limit beyond which there are diminishing returns. In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 4.3%. It is notable that OpenAI embeddings with Logit outperform Word2Vec significantly by around 9.0% and 13.5% at the same dimensions of 256 and 3072, respectively. These observations are unsurprising since OpenAI’s LLMs are state-of-the-art language models that naturally perform better than the classic Word2Vec embeddings.
A more striking observation is that classification model complexity does not improve the prediction (conditional on the same inputs). Specifically, there is no significant accuracy difference between MLP and Logit when using OpenAI-256E or OpenAI-3072E. This observation suggests that a linear classifier on top of fixed LLM embeddings could effectively capture the attractiveness of content. An interesting question is why text embeddings obtained from models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for our downstream classification task. (saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text. Thus, the embeddings generated from a next-word prediction task inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
We further investigate two specific aspects of embedding-based approaches. First, we examine the conditions under which these approaches fail, and second, we quantify the role of the number of training samples on predictive accuracy. For both these analyses, we use logistic regression with linear features as our classification model (since this is the model with the highest accuracy and there is no significant improvement in the accuracy of MLP).
To address the first question, we first calculate the cosine similarity based on the text embeddings for each pair of headlines. This metric captures the similarity between two vectors in inner product space and is defined as the dot product of the vectors divided by the product of their lengths, i.e., ${e_{1}\cdot e_{2}}/(\|e_{1}\|\|e_{2}\|)$, where $e_{1}\cdot e_{2}$ is the dot product of the embedding vectors, and $\|e_{1}\|$ and $\|e_{2}\|$ are the their Euclidean norms. Figure 5(a) shows the distribution of cosine similarity across all headline pairs, whereas Figure 5(b) shows the distributions of cosine similarity separately for two cases: (1) cases where our embedding-based classification correctly predicts which headline is catchier, and (2) cases where our embedding-based classification incorrectly predicts which headline is catchier. We observe a rightward shift in the distribution of incorrect predictions (compared to the distribution of correct predictions). This suggests that incorrect predictions are more common when the headlines are very similar to each other (i.e., higher cosine similarity). There are two possible reasons for this. First, in cases where the headline pairs are very similar, there is a higher chance of Type 1 error (in the t-tests based on our A/B test data), i.e., the headlines may not be significantly different in fact. [^14] Conducting multiple tests in our data can inflate the overall Type I error rate, known as the multiple testing problem. This means there is a higher likelihood of encountering at least one Type I error. Similar headlines are more likely to lead to these errors because the actual difference in their CTRs is likely to be small. This increases the probability that variation in randomly sampled impressions and clicks in A/B tests might lead to significant results by chance, even when there is no true effect. Second, text embeddings may not be able to capture the differences between two headlines when they are similar, i.e., the embeddings may also be very similar. As a result, it is hard for the classification model to differentiate between them. [^15] Note that it is unlikely that this phenomenon is due to problems/issues with the classification model because both Logit and MLP achieve nearly identical performance.
(a) Performance of OpenAI-256E
(b) Performance of OpenAI-3072E
Figure 4: Performance of Logit classification models with OpenAI embeddings under varying training dataset sizes and different embedding sizes.
Next, to address the second question, we conduct experiments varying the proportion of data used for training vs. test. We use the Logit model for these experiments since it offers the best performance and is easy to train. The results from this exercise are shown in Figure 4. As we can see, when we use the 256-dimensional embeddings, approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high performance. However, when we use the 3072E-dimensional embedding, there is a clear over-fitting issue even if we use the simplest Logit model and 100% of the data for training. This suggests that for larger embedding vectors, we need much larger training datasets. [^16] It is unlikely that this is due to issues with the classification model since the Logit model is inherently parsimonious. To further investigate this issue, we experimented with both $L_{1}$ and $L_{2}$ regularization to control over-fitting. [^17] This was done using the penalty and C control parameters in the LogisticRegression function from sklearn package. Unfortunately, regularization did not improve the accuracy on the test data, suggesting that over-fitting is unlikely to be due to learning noise in the training data. Recall that the average headline in the Upworthy tests received 3,569 impressions and 54.26 clicks. Our analysis suggests that we need much larger A/B tests to learn the relationship between a headline’s text embeddings and its attractiveness effectively. In summary, these comparisons suggest that data samples from Upworthy are insufficient to unlock the potential of LLMs fully. Nevertheless, diverting more and more traffic to A/B tests is costly, and this observation is one of the motivations for our approach later in Section 4.
(a) Cosine similarity distribution over all pairs in training and test sets
(b) Cosine similarity over incorrect/correct predicted test pairs
Figure 5: The left panel shows the distribution of cosine similarity for all pairs. The right panel shows the distributions of cosine similarity for both incorrect and correct predictions using the logit classification model with OpenAI-3072E embeddings.
Finally, we discuss the time and compute costs of embedding-based approaches, and contrast them with prompt-based approaches. We find that obtaining text embeddings through the OpenAI API is extremely efficient in terms of time usage and is also more economical compared to direct prompt-based interactions. When using the embedding API, data transmission involves smaller, fixed-size vectors instead of variable-length text, which reduces data overhead and accelerates processing speed. We spent less than \$1 to obtain the text embedding for all 77,245 headlines, while the prompt experiment with only 10,000 pairs cost us around \$20.181818At the time of this study (May 2024), GPT-4 (turbo version) prompt costs \$10 per 1M input tokens and \$30 per 1M output tokens. Further, the embeddings used in our study, text-embedding-3-large, only cost \$0.13 per 1M input tokens and there is no output token cost associated with embeddings. Our entire dataset comprises 77,245 headlines, with an average of 19 tokens per headline, resulting in a total embedding cost of approximately \$0.18. In addition, we found that it is much faster to process embeddings (more than 3,000 pairs per minute), while the prompt experiment is much slower (can only test around 83 pairs per minute). This efficiency can translate into lower operational costs and help with easier integration into real-time applications. Another key advantage of the embedding-based approach is its transparency and robustness. Using simple models like logistic regression with text embeddings from OpenAI can give researchers and managers insight into the training process and source of performance improvements. This transparency is crucial in business settings, where a clear understanding of a model’s behavior and performance is often necessary to mitigate potential risks and biases. In contrast, in-context learning is a black-box approach where the user/manager cannot obtain any insight into the source of the performance improvement.
3.3 Fine-Tuning Open-Sourced LLMs with LoRA
In Section 3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself. The performance of these embedding-based approaches reached 82% accuracy without modifying any parameters in the LLM. This raises a natural question: if we can train LLMs directly, can we achieve better performance? However, training LLMs from scratch is costly and resource-intensive. Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLMs to improve performance while keeping the computational cost manageable. Note that to fine-tune LLMs, we need to be able to access the underlying model parameters. This is only feasible for open-source LLMs, like BERT and LLama, and not for proprietary LLMs like OpenAI’s GPT and Google’s Gemini. The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
For the fine-tuning task, we use Meta’s Llama-3 (meta_llama_3) as our primary base model because it is regarded as one of the most advanced open-source LLMs to date (llama3_model_card). zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance. [^19] Before the widespread adoption of the term “LLM”, BERT was considered among the best language models available, and is still widely used as a benchmark. To provide a comparison with these earlier models, we also fine-tuned DistilBERT (sanh2019distilbert)—a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (devlin2018bert). Details on the implementation and performance of the fine-tuned DistilBERT are shown in Web Appendix A.3. Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b). We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see Section 3.3.4 for details). Hence, we do not consider fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated. Our numerical experiment shows LoRA fine-tuned Llama-3-8b achieves nearly 83% accuracy.
The rest of this section is organized as follows. We first introduce the Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 in Section 3.3.1. Next, we present the LoRA method for efficient fine-tuning in Section 3.3.2. The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning. Finally, we detail our implementation in Section 3.3.3 and report our results in Section 3.3.4.
3.3.1 Transformer and Llama Architecture
The Transformer model, introduced by (vaswani2017attention), has revolutionized the field of NLP with its innovative attention mechanism and highly parallelizable architecture. Unlike traditional recurrent neural network (RNN) based models, the Transformer uses self-attention mechanisms to process input data. This shift allows for greater computational efficiency and significantly improved performance across a range of NLP tasks.
As illustrated in the right portion of Figure 6, the Transformer block comprises multiple layers, each featuring two primary components: a multi-head self-attention mechanism and a position-wise fully connected feed-forward network. Key enhancements include layer normalization (ba2016layer) and the use of residual connections (he2016deep), which improve training stability and help prevent issues such as vanishing gradients. While some specifics are omitted to maintain brevity and readability, it’s important to note that such architectures are subject to rapid evolution and vary across different LLMs.
Figure 6: Illustration of the Transformer block and architecture typical in most LLMs.
The left portion of the figure displays the overall architecture of an LLM, highlighting the layers of Transformer blocks and the connections that represent the attention mask. Lines between Transformer blocks in different layers indicate where attention is applied; the absence of a line signifies no attention interaction. Specifically, a causal attention mask is depicted, ensuring that predictions for each token depend only on preceding tokens. Note that this figure represents one type of attention mask; other types are also employed in various LLMs but are not depicted here. Additionally, only two layers are shown for simplicity; actual models may contain many more layers and connections, which are not depicted here to maintain clarity and focus on key components.
The right portion provides a detailed view of a single Transformer block, internal components included. We omit the depiction of multi-head attention for brevity. Note that different LLMs may use slightly different designs.
Self-Attention Mechanism: The most important innovation in the Transformer block is the self-attention mechanism. It enables each token in a sentence to “pay attention” to all other words when computing its own representation. This is achieved using an attention mask that guides which tokens should influence each other, thereby allowing the model to capture dependencies between words regardless of their distance in the text. Specifically, for a given input sentence, the model first converts each token into an embedding, [^20] This embedding differs from the text embedding used in the previous section. Each embedding here simply corresponds to one token without paying attention to the whole input sequence. The text embeddings used in the previous section are essentially the contextualized embeddings, which are generated by passing the tokens through multiple layers of Transformers, capturing rich contextual information and summarizing the whole input sequence. forming a matrix $X$ where each row corresponds to the embedding of a token. If the sentence contains $n$ tokens, and each token is represented by a $d$-dimensional embedding, then $X$ is an $n\times d$ matrix. From these embeddings, the model generates three matrices: Query $Q$, Key $K$, and Value $V$:
$Q=XW_{Q},\quad K=XW_{K},\quad V=XW_{V}$
where $W_{Q}$, $W_{K}$, and $W_{V}$ are trainable weight matrices of dimensions $d\times d_{q}$, $d\times d_{k}$, and $d\times d_{v}$ respectively. $d_{q}$ and $d_{k}$ are typically set to be the same, allowing for the dot-product operation used in computing the attention scores to be valid.
The attention scores are computed by first taking the dot product of the query matrix $Q$ and key matrix $K$, then scaling by the square root of the key vector dimension $d_{k}$, and applying an attention mask $M$. A softmax function is applied subsequently:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}+M\right)V$
The attention mask $M$ typically includes negative infinity at positions where attention is not applicable, ensuring the attention weights are effectively zero after applying the softmax. At positions where attention is relevant, the mask values are zero. In models like GPT and Llama, a causal attention mask, a.k.a, autoregressive attention, is utilized to prevent future tokens from influencing the prediction of the current token during training. The causal mask is a lower triangular matrix, where positions above the diagonal are set to negative infinity, and positions on and below the diagonal are set to zero. While BERT employs a bidirectional attention mechanism that allows each token to depend on all other tokens in the sequence.
Multi-Head Attention: To capture diverse aspects of the relationships between tokens, the Transformer enhances the self-attention mechanism through the use of multi-head attention. This approach involves performing multiple self-attention operations in parallel, with each “head” having its own independent $Q$, $K$, and $V$ matrices. The outputs from these operations are concatenated and then linearly transformed to produce the final attention output:
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W_{O},$
where each head computes an independent self-attention process:
$\text{head}_{i}=\text{Attention}(XW_{Q}^{i},XW_{K}^{i},XW_{V}^{i}),$
with $W_{Q}^{i}$, $W_{K}^{i}$, and $W_{V}^{i}$ being the learnable parameter matrices specific to the $i$-th head. $W_{O}$ is the weight matrix with dimension $hd_{v}\times d$, and is applied to the concatenated outputs from all heads, integrating the distinct perspectives captured by each head into a unified output.
Feed-Forward Networks: Each layer of the Transformer also includes a position-wise fully connected feed-forward network. This network comprises two linear transformations with a ReLU activation function in between:
$\text{FFN}(x)=\max(0,xW_{1}+b_{1})W_{2}+b_{2}$
where $W_{1}$, $W_{2}$, $b_{1}$, and $b_{2}$ are learnable parameters. The final output of the FNN has the same dimension $d$ as the input embeddings. This network acts on each position separately and identically, providing additional non-linearity and enhancing the model’s ability to learn complex patterns in the data.
Llama Architecture: The Llama model enhances the basic Transformer architecture to handle vast amounts of data and leverage extensive computational resources effectively, establishing it as one of the most powerful open-source language models currently available. Llama-3 retains essential components such as multi-head self-attention and feed-forward layers from the original Transformer model but incorporates several techniques to stabilize training and boost efficiency and performance. These enhancements include the application of pre-normalization with RMSNorm (zhang2019root), the SwiGLU activation function (shazeer2020glu), rotary positional embeddings (su2024roformer), and grouped-query attention (ainslie2023gqa), among others. Like Llama-3, other LLMs share a similar Transformer architecture but vary in the number of Transformer layers, training corpora, size of the weight matrices, activation functions, etc.
In the multi-layer structure of LLMs, the output of one layer serves as the input to the subsequent layer. Each layer’s output is a matrix of the same dimensions as the input matrix $X$, with each row representing the transformed embedding of a token, enriched with context from other tokens through the self-attention mechanism. After processing through all the layers, the final output matrix contains the contextualized embeddings of all tokens in the input sentence. Unlike RNNs that process input sequences one token at a time, the Transformer architecture can process whole input sequences in parallel, significantly enhancing processing speed and efficiency.
It is worth noting that LLMs do not necessarily operate in a conversational manner, nor do they always use language as their output. As we will demonstrate in Section 3.3.3, we employ the Llama model as a classifier, where the output is the distribution of predicted classes. In applications like conversational AI, exemplified by GPT, the contextualized embeddings from the last token of the last layer of the Transformer are fed into a linear layer followed by a softmax function, the configuration of which transforms these embeddings into a probability distribution over the vocabulary, enabling the model to predict the next token in the sequence.
3.3.2 LoRA Fine-Tuning Technique
Traditional fine-tuning of LLMs involves updating a substantial number of parameters, which can be computationally expensive and memory-intensive. To efficiently fine-tune the Llama-3 model on hardware with limited GPU memory, we employ Low-Rank Adaptation (LoRA), a popular Parameter-Efficient Fine-Tuning (PEFT) method developed by (hu2021lora), enabling practical fine-tuning of large-scale LLMs on medium-scale hardware. [^21] There are many alternatives for PEFT, like Prefix Tuning (li2021prefix), Adapter Layers (houlsby2019parameter), and BitFit (zaken2021bitfit). We select LoRA due to its excellent balance between simplicity and efficiency, making it the top choice for fine-tuning. LoRA operates under the assumption that updates during model adaptation exhibit an intrinsic low-rank property and introduce a set of low-rank trainable matrices into each layer of the Transformer model. This approach leverages the low-rank decomposition, which significantly reduces both memory usage and computational time through a significantly reduced number of trainable parameters.
We illustrate LoRA in Figure 7. We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$. We use $d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices. We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
$W^{\prime}=W+\Delta W,\quad\mathrm{with}\quad\Delta W=AB,\quad\mathrm{where}\quad A\in\mathbb{R}^{d\times r},\quad B\in\mathbb{R}^{r\times d_{W}}.$
Figure 7: Illustration of LoRA.
LoRA essentially introduces the low-rank matrices $A$ and $B$ to approximate the weight matrix change $\Delta W$ after the fine-tuning. These low-rank matrices are updated during the fine-tuning process, while the original weight matrix $W$ remains frozen. This strategy significantly reduces the number of trainable parameters from $d\times d_{W}$ to $r\times(d+d_{W})$ because the rank $r$ is typically set to a small value, such as 4 or 8, to balance the trade-off between model capacity and computational efficiency.
3.3.3 Implementation Details
We use exactly the same training and test sets as in Section 3.2, but the inputs here are the original headlines with the text information. To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them. For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 8. Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence. Thus, the last token in the sequence, being informed by all prior tokens, is used for classification. This embedding vector of the last token, output by the last Transformer layer and having a dimension of 4096, is processed through a linear layer producing an output of two dimensions representing the logits. Then the logits are transformed into probabilities by the softmax function, indicating the likelihood of each headline being more impactful. The softmax function is expressed as:
$\text{Softmax}(Z)_{i}=\frac{e^{z_{i}}}{\sum_{j}e^{z_{j}}}$
where $Z$ is the vector of logits, and $i$ indexes the specific class. During fine-tuning, our objective is to minimize the cross-entropy loss on the training set.
Figure 8: The classification pipeline using the Llama model. Headlines are first tokenized and each token is converted into an embedding vector. For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration. The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens. Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity. The output from the last token is fed into a classification head, which computes the probability that one headline is more impactful than the other.
We implement LoRA Fine-tuning using the PEFT library developed by Hugging Face. Our LoRA configuration settings include setting the dimension of the low-rank matrices, denoted as $r$, to 4. The lora_alpha parameter, acting as a scaling factor for the low-rank matrices, is set to 4. This parameter controls how much the learned low rank matrix would be weighted. Adjusting lora_alpha allows us to control the adaptation strength, enabling the model to integrate new information effectively without losing its pre-trained knowledge.
Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora). Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$. We set the dimension of the low-rank matrices $r$ to 4. To mitigate the overfitting, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices. Furthermore, we enable 16-bit mixed precision training. Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly. This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which takes only 0.023% of the total 7,506,636,800 parameters. We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training. A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process. This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
Lastly, we want to briefly discuss the GPU requirement and monetary cost of implementing LoRA. The GPU requirements for fine-tuning depend on factors such as model size, batch size, optimization algorithm, and floating-point precision. In our experiment, without LoRA, full-parameter fine-tuning would require over 100 GB of GPU memory. However, LoRA fine-tuning only requires 30 GB of GPU memory. To provide a more straightforward demonstration of the benefit of the saved GPU requirement, we showcase two GPU models that are commonly used for LLM training. Without LoRA, the NVIDIA H100 GPU with nearly the largest GPU memory, 80 GB is insufficient for fine-tuning. In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient. The price of an NVIDIA H100 GPU is ten times higher than an NVIDIA V100 GPU in May 2024. Thus, using LoRA significantly reduces the cost, making fine-tuning more accessible to researchers and practitioners.
3.3.4 Performance Analysis of Fine-tuning-based Approaches
We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 9. We observe a slightly better performance of 82.79% accuracy compared to the best result (82.33% accuracy) using OpenAI embeddings in Section 3.2. It is important to note that, if compared under the same LLM, LoRA fine-tuned Llama-3-8b outperforms the logistic classifier using embeddings from Llama-3-8b (around 77% accuracy).
Figure 9: Loss curve and accuracy curve as the number of training epochs increases. We repeat the fine-tuning process three times and display the average loss and accuracy. For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits. The final accuracies of the three replicates on the test set are 83.04%, 82.73%, and 82.59% respectively. We add the best performance (dash line) obtained using OpenAI’s embedding in Section 3.2 for a comparison.
From Figure 9, we observe that the loss evaluated on the training set is consistently lower than the loss on the test set, and the accuracy on the training set is consistently higher than the accuracy on the test set, indicating overfitting. Overfitting refers to the phenomenon that occurs when a machine learning model learns the details and noise in the training data to the extent that it negatively impacts the model’s performance on new, unseen data. This often happens because the model becomes too complex or because the training set is too small, causing the model to memorize specific examples and noise rather than general patterns. As a result, the model performs exceptionally well on the training data but fails to generalize, leading to relatively poor performance on test data.
In our case, the Llama-3-8b model is inherently complex. Although using LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex. Common approaches to address overfitting include reducing model complexity, adding regularization, and increasing the dataset size. We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay, in an attempt to mitigate overfitting. However, these modifications generally reduce performance on the training set without yielding significant improvements on the test set. We suspect that overfitting could be further alleviated, and test performance improved, with a larger and more diverse dataset. The insight for practitioners is that continuously updating the model with new data, such as through A/B testing on a news website to collect more pairwise comparison data, may lead to better generalization and performance over time. Instead of fine-tuning a model version and continuously using it, we suggest that companies keep updating the model as they collect larger datasets.
After fine-tuning the model, several cost-efficient techniques are available for deploying it for inference, significantly reducing computational costs and hardware requirements. For instance, llama.cpp allows LLM models to be deployed on machines without a GPU, making implementation more affordable. Moreover, model compression techniques such as quantization, sparsification, and knowledge distillation can drastically cut inference computation costs while maintaining, or only slightly affecting, model performance. Our approach ensures that the fine-tuned model is not only cost-efficient but also robust and secure when hosted on our own servers. For a comprehensive survey on efficient LLM deployment and inference techniques, please refer to (zhou2024survey).
4 Our Approach: LLM-Assisted Online Learning Algorithm
So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 82% for the embedding-based and LoRA fine-tuning approaches). The main advantage of these pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it. Thus, there is no need to use any traffic for experimentation. The downside, of course, is that these models are not fully accurate—in around 18% of the cases, they are likely to mispredict which headline is better. This can of course, lead to lower clicks and revenues for the firm. [^22] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth. However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples. Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
On the other hand, we have the current status quo, which is the use of experimentation-based approaches to choose headlines/content. As mentioned in Section 1, there are two broad experimentation-based methods—(1) A/B tests and (2) Online learning algorithms or bandits. In particular, the standard practice at Upworthy during the time of the data collection was A/B tests, where the firm allocated a fixed amount to different headlines and then chose the best-performing one for the remaining traffic. This approach is also known as Explore and Commit (E&C) and is commonly used in the industry due to its simplicity and low engineering cost; see chapter six of (lattimore2020bandit). The main benefit of this approach is that, with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased; as such it is the gold-standard approach for identifying the best arm (or headline in this case). The downside, of course, is that by assigning traffic to all the headlines during the exploration phase or A/B test, the firm incurs high regret (i.e., low profits since some of headlines in the test are likely have low CTRs). Therefore, many modern technology and media firms such as NYTimes and Yahoo use online learning algorithms or bandits that leverage the explore-exploit paradigm to lower the experimentation cost while still learning (lattimore2020bandit). These algorithms move traffic away from poorly performing arms/headlines dynamically in each period, based on the observed performance of each arm till that period. Because of this greedy behavior, they tend to switch to the best-performing arms quickly, and incur lower regret. Nevertheless, one drawback of these methods is that they start with the assumption that all headlines are equally good, i.e., they suffer from a cold-start problem, which can result in wasting traffic on the weaker headlines early on.
In this section, we present a framework that combines the benefits of LLM-based models with experimentation-based methods. Our key idea is that firms can address the cold-start problem in content experiments by leveraging predictions from LLM-based approaches as bayesian priors in a first step and combine them with online learning algorithm in the second step. Thus, our approach, termed LLM-Assisted Online Learning Algorithm (LOLA), continuously optimizes traffic allocation based on LLM predictions together with the data collected from the online experiment in real-time.
The rest of this section is organized as follows. In Section 4.1, we discuss the general LOLA framework, where we formalize the media firm’s problem and then present our algorithm. Next, in Section 4.2, we discuss the implementation of our proposed algorithm, benchmarks, and the experimental setup. Finally, we present our numerical results in Section 4.3.
4.1 Bandit Framework and LOLA
In the following, we first present the Multi-Armed Bandits (MAB) framework to formalize our problem and then discuss our proposed approach. MAB is a class of sequential decision-making problems where a learner must choose between multiple arms over periods to maximize cumulative reward. In our setting, different arms refer to different headlines associated with the same content, and periods refer to impressions where different headlines with the same content are displayed. Let $[K]:=\{1,2,\dots,K\}$ represent the set of $K$ arms, and $[T]:=\{1,2,\dots,T\}$ represent the entire decision horizon for a test. At each time step $t$, the learner selects an arm $a_{t}\in[K]$ and receives a reward $r_{a_{t}}$ drawn from a probability distribution specific to the chosen arm $a_{t}$. In our setting, we assume each arm has a constant click-through rate (CTR), and the reward is the click feedback, which follows the Bernoulli distribution with the probability equal to the CTR [^23] This reflects a stochastic bandit setting rather than a Bayesian bandit setting. However, it is important to note that our proposed algorithms can also be applied in a Bayesian setting by changing the definition of regret and the process of generating reward feedback. For completeness, we also present a Thompson sampling version of our proposed LOLA algorithm in Web Appendix C..
We first introduce the most common framework, regret minimization. The goal in this framework is to minimize cumulative regret, which is defined as the difference between the reward obtained by always pulling the optimal arm and the reward actually accumulated by the learner. Mathematically, the regret $R_{T}$ after $T$ trials is given by:
$R_{T}=\max_{k\in[K]}\sum_{t=1}^{T}r_{a^{*}}-\sum_{t=1}^{T}r_{a_{t}},$
where $a^{*}=\arg\max_{k\in[K]}\mathbb{E}[r_{k}]$ is the optimal arm with the highest expected reward.
As a complement to regret minimization, we also tackle the problem using the Best Arm Identification (BAI) framework. The goal of BAI is to identify the arm with the highest reward within a given limited number of pulls. Formally, given a fixed budget $B$, we aim to find the arm $a^{*}$ that maximizes the expected reward after $B$ trials.
Before we introduce our LLM-assisted online learning algorithm, we introduce the widely applied pure learning algorithm without LLMs, the Upper Confidence Bound (UCB) algorithm. The UCB stands out because it is provably asymptotically optimal. The UCB policy is straightforward and selects the arm that maximizes upper confidence bound on the estimated rewards, specifically:
$a_{t}=\arg\max_{k\in[K]}\left(\bar{\mu}^{t}_{k}+\sqrt{\frac{2\log(1/\delta)}{n^{t}_{k}}}\right),$
where $\bar{\mu}^{t}_{k}$ is the estimated mean reward of arm $k$ up to time $t$, $n^{t}_{k}$ is the number of times arm $k$ has been played up to time $t$, and $\delta$ is called the confidence level, which quantifies the degree of certainty. The intuition behind UCB’s effectiveness in balancing exploration and exploitation lies in its construction. By adding a confidence term $\sqrt{{2\log(1/\delta)}/{n^{t}_{k}}}$ to the estimated mean reward $\bar{\mu}_{k}$, UCB ensures that arms with fewer selections (higher uncertainty) are given a chance to be explored. As time progresses and more data is collected, the confidence term diminishes, allowing the algorithm to exploit arms with higher observed rewards. This dynamic adjustment enables the UCB algorithm to systematically explore under-explored arms while exploiting arms with high rewards, thus effectively addressing the exploration-exploitation trade-off inherent in the MAB problem. However, we need to set the key hyperparameter $\delta$ such that it should be small enough to ensure optimality with high probability, but not too small that suboptimal arms are over-explored. Typically, one can set the confidence term as $\alpha\sqrt{\log t/n^{t}_{k}}$ and fine-tune the scale term $\alpha$ for specific problems.
Under the current framework, we may have more than two headlines in one test (i.e., $K>2$), rendering the previous LLM-based binary classification models inapplicable. To compare multiple arms, we adapt the previous LLM-based classification models to predict CTR by changing the loss function from log loss to mean squared error (MSE) while keeping the training and fine-tuning procedures unchanged. Using CTR as the outcome measure does not adversely affect decision-making performance. For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.56% accuracy obtained with logistic regression. This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
Given the above formalization of the decision process and the LLM-based CTR prediction model, the remaining key challenge is to design a decision policy that incorporates these LLM-based CTR predictions. Inspired by the 2-Upper Confidence Bounds (2-UCBs) policy designed by (gur2022adaptive), we propose the following LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) algorithm. Essentially, we treat the LLM’s CTR predictions as auxiliary information prior to the start of online experiments, equating this auxiliary CTR information to additional impressions and click outcomes for each arm.
Algorithm 1 LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs)
1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
2:Hyperparameter Fine-Tuning: Use another subset of data to select two hyperparameters for the algorithm, $\alpha\in\mathbb{R}^{+}$ for controlling the upper bound, $n^{\text{a}ux}$ as the LLM’s equivalent auxiliary sample size.
3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$, LLM-based CTR prediciton $\bar{\mu}^{\text{a}ux}_{k}\leftarrow\mathcal{M}(x_{k})$, regular CTR initialization $\bar{\mu}^{1}_{k}=0$, accumulated impressions $n^{1}_{k}\leftarrow 1$, and accumulated clicks $c^{1}_{k}\leftarrow 0$, for all arms $k\in[K]$.
4:for $t=1$ to $T$ do
5:     Calculate the first UCB for all arms $k\in[K]$: $U_{k}^{1}=\bar{\mu}^{t}_{k}+\alpha\sqrt{\frac{\log t}{n^{t}_{k}}}$.
6:     Calculate the second UCB for all arms $k\in[K]$: $U_{k}^{2}=\frac{c^{t}_{k}+\bar{\mu}^{\text{a}ux}_{k}n^{\text{a}ux}}{n^{t}_{k}+n^{\text{a}ux}}+\alpha\sqrt{\frac{\log t}{n^{t}_{k}+n^{\text{a}ux}}}.$
7:     Play the arm $a_{t}=\arg\max_{k\in[K]}\min\{U_{k}^{1},U_{k}^{2}\}$.
8:     Observe the payoff $r_{a_{t}}$
9:     Update $n^{t+1}_{a_{t}}\leftarrow n^{t}_{a_{t}}+1$, $c^{t+1}_{a_{t}}\leftarrow c^{t}_{a_{t}}+r_{a_{t}}$, and $\bar{\mu}^{t+1}_{a_{t}}\leftarrow\frac{c^{t+1}_{a_{t}}}{n_{a_{t}}^{t+1}}$.
10:end for
In the training phase of the LLM-2UCBs algorithm, we can utilize various LLM-based methods, including LoRA fine-tuning using the original text information $x$, and embedding-based CTR prediction models where $x$ represents the embedding vector. The feature vector $x$ can also include additional information, such as user-specific data. However, since Upworthy does not collect user information and ultimately recommends the same headline to all users, incorporating user information into the algorithm is unnecessary in this specific context. Nonetheless, we emphasize that our algorithm is flexible and can be adapted to incorporate user information when it is available.
There are two hyperparameters in the LLM-2UCBs. The upper bound control hyperparameter, $\alpha$, also shows up in the regular UCB algorithm. The hyperparameter $n^{\text{aux}}$ is specific to our proposed algorithm and indicates the equivalent auxiliary samples for initialization. At a high level, we can regard the LLM predictions as prior information or the information flow before the algorithm starts. This prediction is approximately equivalent to initially generating $n^{\text{aux}}$ impressions for each arm and observing their outcomes to obtain the sample average CTR, $\bar{\mu}^{\text{aux}}_{k}$. The hyperparameter fine-tuning process is simple. We can utilize another randomly sampled dataset to test combinations of two hyperparameters and select the combination that yields the highest accumulated rewards.
During the online learning phase, the key step is balancing the exploration and exploitation trade-off. We capture this trade-off using the 2-UCBs rule. Essentially, we have two valid UCBs: $U_{k}^{1}$ is the regular UCB, which uses observed impressions and clicks to construct the mean reward estimator and adds a regular upper bound. The second UCB, $U_{k}^{2}$, incorporates LLM predictions. The calculation of $U_{k}^{2}$ is straightforward. The mean reward component is simply the sample average, taking into account the “auxiliary” samples generated by the LLM before the learning phase begins. These auxiliary samples are imaginary and approximate the richness of information contained in the LLM prediction. The next step is to choose the smaller of the two UCBs, $\min\{U_{k}^{1},U_{k}^{2}\}$, as the final UCB for selecting the arm. Note that if $n^{\text{aux}}=0$, the 2-UCBs rule is the same as the standard UCB rule. Conversely, when $n^{\text{aux}}\rightarrow\infty$, the 2-UCBs rule is equivalent to a pure exploitation policy (a.k.a. greedy policy) using the LLM-based CTR prediction. After playing the arm and observing the outcome, we update the impressions, clicks, and mean reward accordingly, following standard bandit algorithm procedures.
The 2-UCB algorithm offers several advantages. First, its nonparametric nature means it does not assume a linear relationship, making it flexible for various applications. Second, it is easy to implement, requiring the tuning of only two hyperparameters. Third, it can be applied to any LLM-based model. Notice that we do not use the contextual information $x$ during the online learning phase. The primary reason for this is to avoid updating the LLM prediction model during the learning phase, which would require significant engineering effort due to the large-scale nature of training LLMs. However, if engineering costs are not a concern, and for the completeness of this paper, we provide a linear bandit algorithm based on (chu2011contextual), taking into account LLM-based information and user features for personalized recommendations. Despite this, the linear bandit has limitations due to its linear assumption. Therefore, we propose another potential solution based on the FALCON algorithm (simchi2022bypassing). We discuss the pros and cons of these bandit algorithms in Appendix B.
Note that this framework is general and can accommodate any number of headlines and does not rely on binary classification and/or assume that differences between headlines are significant. As such, in the rest of the paper, we utilize the full dataset from Upworthy to test the performance of our proposed algorithm and benchmark approaches (unlike the analysis in Section 3, where we focus only on pairs of headlines with significant differences).
4.2 Benchmarks and Implementation Details in Regret Minimization
In this section, we introduce each of the following algorithms and their implementations in detail, as we will compare them in the next subsection.
  • $\bullet$
    Explore-then-Commit (E&C)
  • $\bullet$
    Pure LLM
  • $\bullet$
    Standard UCB (UCB)
  • $\bullet$
    LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs)
At a high level, we randomly split our data into three exclusive folds based on the test ID, ensuring that all headlines in the same test are in the same fold and that the same headline does not appear in different folds. Specifically, we use 70% of the data for training, 10% for fine-tuning, and 20% for testing. All hyperparameters are selected using the fine-tuning data, and the performance of the algorithms is evaluated on the test data. Note that only our proposed LLM-2UCBs and the pure LLM method utilize the training data. In our numerical study, we train the linear regression model with 3,072 dimensions text embeddings from OpenAI as inputs to predict CTRs
Upworthy assigns traffic to headlines with equal probability during the experiment phase, [^24] Upworthy explicitly mentioned they didn’t conduct power analysis before the experiment, and based on the disclosed impression amount data in the same test, we make this claim with high confidence due to the almost same amount of impressions. and, at the end of the experiment phase, select the winner headline for all future impressions. This practice closely resembles the E&C algorithm in the bandit literature. We implement this method as a benchmark in our numerical experiments. The E&C method has a single hyperparameter to tune: the proportion of periods for exploration. We select the proportion from $\{0.1,0.2,0.3,0.4,0.5\}$ with the highest accumulated rewards. This ratio is fine-tuned for experiments with different total periods $T$, as a longer time horizon requires a smaller proportion of traffic for exploration, given a constant number of arms.
Interestingly, the other two methods can be viewed as special cases of our proposed LLM-2UCBs algorithm. The Pure LLM (a.k.a pure exploitation, greedy algorithm) is a special case of LLM-2UCBs by setting $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results. This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon. Similarly, the standard UCB algorithm is a special case of LLM-2UCBs with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions. In our UCB implementation, we fine-tune the confidence control hyperparameter $\alpha$ from $\{0.02,0.04,\dots,0.10\}$ for different $T$. After fine-tuning, we find $\alpha=0.08$ in UCB gives the highest accumulated rewards. For LLM-2UCBs, we fine-tune the combination of hyperparameters $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$. After fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ in LLM-2UCBs.
After the fine-tuning of algorithms, we evaluate all algorithms under the test dataset, which has never been seen for CTR training and algorithm tuning. Algorithms do not know this true testing CTR in advance. We use the average CTR in the test data as the true CTRs to generate the click feedback in the numerical simulations. We argue that using this CTR as the true CTR in the simulation is justified because each headline was tested with a large amount of traffic, resulting in a relatively narrow confidence interval at Upworthy.
UCB and LLM-2UCBs are both asymptotically optimal in minimizing regret, which means there is no difference between these two algorithms when the number of time horizons $T$ goes to infinity, and the LLM-based prior information is not helpful in the long run. However, in practice, the time horizon is never infinite. Therefore, we compare different algorithms under different time horizons. Since there can be a varying number of headlines/arms in different tests, we scale the time horizon using the metric “Traffic/Impression per headline” (we use the notion $\tau$ to represent) from $\{50,100,200,400,600,800,1000\}$. Specifically, if there are three headlines in one test and the multiplier is $100$, we set $T=100\times 3=300$ for this test. It is important to investigate the small number multipliers because they also partially address the question of what happens if we are concerned about inference and can only use a small amount of traffic for adaptive experiments.
4.3 Numerical Results
We visualize the numerical results in Figure 10 and document the pairwise comparison of different algorithms, reporting the relative percentage reward improvement and the p-value from the t-tests.
LLM-2UCBs generally performs the best among all algorithms. Notably, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm. For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends. This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance.
When comparing LLM-2UCBs with Pure LLM, we find that LLM-2UCBs outperforms Pure LLM in most cases. An exception occurs when $\tau=50$, where there is no significant difference between LLM-2UCBs and Pure LLM. This is due to the constrained exploration ability of LLM-2UCBs in short planning horizons, leading it to perform similarly to using only LLM prior information. However, as $\tau$ increases to 1000, LLM-2UCBs significantly outperforms Pure LLM, generating 5.46% more clicks. This observation indicates that we cannot fully rely on pure LLM predictions alone.
Comparing LLM-2UCBs with regular UCB, it is evident that leveraging LLM can always enhance UCB performance, although this improvement diminishes as the time horizon increases. This is due to the asymptotic optimality of the UCB algorithm, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period.
E&C consistently underperforms compared to LLM-2UCBs because it neither leverages LLM predictions nor achieves asymptotic optimality. Even when $\tau=1000$, LLM-2UCBs is still better than E&C by 4.41%, which is a large improvement in this industry. E&C is also significantly worse than Pure LLM when $\tau\leq 400$, and shows no significant difference from Pure LLM when $\tau\geq 600$. This suggests that the current Upworthy practice is suboptimal.
In summary, our findings underscore the superior performance of LLM-2UCBs, particularly as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.
Figure 10: Average clicks per experiment per period under different time horizon multipliers. Note that the Y-axis captures the average clicks per test per period. For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$. The Y value is simply the average of this number $0.03$ over all tests. This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
Parameter $\tau$ LLM-2UCBs vs Pure LLM LLM-2UCBs vs UCB LLM-2UCBs vs E&C Pure LLM vs UCB Pure LLM vs E&C UCB vs E&C
50 $0.53$ $4.06$ $6.63$∗∗∗∗ $3.52$ $6.07$∗∗∗ $2.46$
100 $1.04$ $5.01$∗∗∗∗ $6.20$∗∗∗∗ $3.93$∗∗ $5.11$∗∗∗∗ $1.13$
200 $1.92$∗∗∗∗ $4.14$∗∗∗∗ $4.89$∗∗∗∗ $2.18$ $2.92$∗∗ $0.72$
400 $3.57$∗∗∗∗ $2.51$∗∗∗ $6.36$∗∗∗∗ $-1.02$ $2.70$∗∗∗ $3.76$∗∗∗∗
600 $3.78$∗∗∗∗ $1.74$∗∗ $3.60$∗∗∗∗ $-1.96$∗∗ $-0.17$ $1.83$∗∗∗
800 $5.00$∗∗∗∗ $1.02$ $4.97$∗∗∗∗ $-3.79$∗∗∗∗ $-0.03$ $3.91$∗∗∗∗
1000 $5.46$∗∗∗∗ $1.10$ $4.41$∗∗∗∗ $-4.13$∗∗∗∗ $-1.00$ $3.27$∗∗∗∗
Table 7: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001. For example, under the impression per headline equal to 50, the average click per test of LLM-2UCBs and Pure LLM is 0.05149 and 0.05122, respectively; so the relative improvement is calculated as ${0.05149}/{0.05122}-1=0.53\%$. Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
5 Conclusion
We note that our approach is quite general and is agnostic to the exact nature of the LLM-based approach used. Thus, any of the LLM-based approaches discussed in Section 3 can be used in the first step of our approach. This is important since the LLM models and fine-tuning approaches continue to advance quickly; so the plug-and-play nature of our approach extends well to this setting. In addition, when considering personalization and accessing contextual information (e.g., readers’ features), we can incorporate this extra information as inputs to the LLM-based CTR prediction model and integrate it into contextual bandit algorithms. Specifically, we provide two variants of LOLA for this purpose in Web Appendix B.
In conclusion, our study introduces LOLA, a novel framework that combines the predictive power of LLMs with the adaptive efficiency of online learning algorithms to enhance content experimentation. By leveraging LLM-based prediction models as priors, our approach minimizes regret in real time, leading to significant improvements in total reward/clicks. Furthermore, through the numerical analysis based on a large-scale dataset from Upworthy, we demonstrate that LOLA outperforms traditional A/B testing and pure online learning algorithms, particularly in settings with limited experimental traffic or a large number of arms. These conditions are prevalent in the current content industry, where platform users can easily create and share their content, such as on platforms like X and TikTok. Our numerical results also indicate that while LLMs provide valuable predictions, one cannot fully trust the prediction results solely from LLMs since experimentation-based approaches still remain the gold standard and can be leveraged to correct the errors in LLMs.
From a managerial perspective, LOLA offers a versatile solution applicable to various scenarios where firms need to decide on content display, such as news articles, headlines, advertisements, or website content. Utilizing open-source LLM models, this method is cost-effective and can be easily deployed across different tasks once LLM models are fine-tuned on relevant datasets. The fine-tuning process is cost-efficient and can be performed on entry-level GPUs, making advanced LLM-based techniques accessible to firms with limited budgets. Additionally, we document that prompt-based approaches, which have been the standard method in recent marketing literature, perform poorly in this content attractiveness comparison task, underscoring the importance of the fine-tuning approach.
Finally, we highlight that LOLA can be readily adapted to other measures beyond clicks. The click-based metric, while widely used by Upworthy a decade ago, has raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity. The industry is rapidly evolving, with Upworthy now focusing more on content quality and humanity, although clicks still remain an important measure; see more detailed discussion in (Upworthy2023). Our LOLA framework can easily accommodate these changes by adjusting the reward notion in the algorithm to align with new business goals, demonstrating its flexibility and relevance in a dynamic industry.
Code and Data
The code used in this study is publicly available at LLM News Github Repository to enable the reproduction of our results and to support future research that applies LLMs to experimentation and business problems.
Funding and Competing Interests Declaration
Author(s) have no competing interests to declare.
Appendix A Supplementary Materials for Section 3
A.1 Pairwise Comparison Between GPT Prompts
To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on 1,000 correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other. The mean differences and p values are documented in Figure A1 and Figure A2. The random guess model simply predicts the first headline wins with a half probability. The nomenclature for GPT prompts in these two figures follows a specific format. The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are flipped. A value of 0 indicates the correct label, while a value of 1 indicates the wrong label. For example, GPT-4-0-0 means zero-shot prompting using GPT-4, and GPT-3.5-5-1 means in-context learning with 5 demonstrations and wrong labels using GPT-3.5.
Figure A1: Average accuracy differences between different GPT prompts. The number in the heatmap presents the accuracy of prompts indicated on the left minus the accuracy of prompts indicated at the bottom. Shaded blocks indicate insignificant differences.
Figure A2: The number in the heatmap presents the p-value from the paired t-test between different GPT prompts.
A.2 Implementation Details of MLP Classification Models
Hyperparameters of the network structure. We employ a fully connected neural network with ReLU activation functions for the hidden layers and a sigmoid activation function for the output layer. This configuration is widely used in practice. To determine the number of layers and the hidden nodes of each layer, we choose the number of layers within the range of $\{1,2,3,4,5\}$, and set the number of nodes in the first fully connected layer from the set $\{128,256,512,1024,2048\}$. Besides, we adjust the previous-post ratio, which is the proportion of nodes in each layer relative to those in the subsequent layer. This ratio here is in $\{2,4,8\}$. Note that the previous-post ratio becomes irrelevant to the results when there’s only one hidden layer. This is because the final output layer has only one dimension for ranking the most attractive headline. We incorporate the dropout after each hidden layer to mitigate the overfitting. Dropout layers randomly deactivate neurons based on a specified dropout rate, while scaling up the remaining neurons to maintain the same mean of the layer. In this work, we set the dropout rate in $\{0.3,0.5\}$. All hyperparameters are automatically selected through hyperparameter optimization, which is introduced below.
Hyperparameters of the training procedure. The MLP is implemented in PyTorch, with data split into 70% for training, 10% for validation, and 20% for testing. To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch. Determining the appropriate batch size is crucial, as large batch sizes may lead to convergence to local optima, while small batch sizes may hinder convergence. We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$. Additionally, to mitigate overfitting, we apply $L_{2}$ regularization, with the regularization weight chosen from $\{1e-6,1e-4,1e-2\}$. We use binary cross-entropy as the loss function and employ stochastic gradient descent for training. We implement early stopping by monitoring the validation set loss, halting the training process if the reduction in loss over 60 epochs falls below a tolerance threshold of $10^{-4}$.
Hyperparameter optimization. We optimize the parameters mentioned above through Bayesian Optimization (snoek2012practical). This Bayesian technique uses Bayesian inference to construct a probabilistic model of the objective function, where the parameters are updated iteratively with new observations. This iterative process balances exploration of the parameter space to discover promising regions and exploitation of known high-performing areas. In this work, we consider the accuracy of the validation set as the objective and optimize hyperparameters selected from the specified parameter space, as delineated by the ranges mentioned previously. We choose the Gaussian process as the estimator with the automatic optimizer, which is commonly adopted by the community. To ensure a judicious trade-off between efficacy and efficiency, we repeat the optimization process 50 times for input data of 256 dimensions and 30 times for data of 3,072 dimensions.
A.3 Results of DistilBERT LoRA Fine-Tuning
One difference between DistilBERT and Llama-3 is that DistilBERT is a bidirectional Transformer model, which means that its attention mechanism allows the model to learn the context from both directions, while Llama-3 is a unidirectional model. An illustration of the structure of the DistilBERT model combined with a classification head is shown in Figure A3. By default, DistilBERT can handle two headlines together as input. The format of the input is: [CLS] Headline 1 [EOS] Headline 2 [EOS]. The [CLS] token is a special token used to get the output of the model for classification, and the output of the last Transformer layer corresponding to this token is used as the input to the classification head. The [EOS] token is used to separate the two sentences.
Figure A3: The classification pipeline using the DistilBERT model. Headlines are first tokenized, and each token is converted into an embedding vector. For illustrative purposes, only two Transformer layers are shown. The figure highlights the use of bidirectional attention in DistilBERT, whereby each token can incorporate information from any token. Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity. The output from the first token (CLS token) is fed into a classification head, which computes the probability that one headline is more impactful than the other.
Another difference is that DistilBERT is a smaller model compared to Llama-3-8b. DistilBERT has around 66 million parameters, which is only 0.94% of the parameters of Llama-3-8b. Due to the smaller model size, we are able to conduct full parameter updates during the fine-tuning process without using LoRA. We use the same learning rate scheduling, number of training epochs, and train-test set split as in the experiment detailed in Section 3.3.3. The accuracy of DistilBERT LoRA fine-tuning on the same test set is around 75%. More details can be found in the code repository.
Appendix B Algorithms with Contextual Information
For the completeness of the paper, we introduce two additional LOLA algorithms in this appendix, which leverage contextual information.
For linear contextual bandits, we assume that at each round, we first observe the context vector $x_{t,k}\in\mathbb{R}^{d}$ for each arm $k\in[K]$. This context information can include LLM embedding vectors and user information, among other features. We denote by $r_{t,k}$ the random reward of arm $k$ at round $t$. After pulling one arm at each round, only the reward of that arm, $r_{t}$, is revealed. We make the linear realizability assumption that the expected reward is linear in the context vectors, which means $\mathbb{E}[r_{t,k}|x_{t,k}]=x_{t,k}^{\top}\theta^{*}$. The weight vector $\theta^{*}\in\mathbb{R}^{d}$ is unknown and needs to be learned. Additionally, we assume access to a historical dataset of size $M$, denoted by $\{(x_{t,k},r_{t})\}_{t=-M}^{-1}$. With these definitions, we are ready to present our proposed Algorithm 2, named LLM-Assisted Linear Contextual Bandit. The LLM-LinUCB algorithm extends LinUCB (chu2011contextual). Under the linear realizability assumption, this algorithm achieves asymptotically optimal regret, matching the lower bound up to logarithmic factors.
Algorithm 2 LLM-Assisted Linear Contextual Bandit (LLM-LinUCB)
1:LLM Training Phase: Calculate $A\leftarrow I_{d}+\sum_{t=-M}^{-1}x_{t,a_{t}}x_{t,a_{t}}^{\top}$ (The $d$-by-$d$ identity matrix), and $b\leftarrow\sum_{t=-M}^{-1}x_{t,a_{t}}r_{t}$
2:Hyperparameter Fine Tune: $\alpha\in\mathbb{R}^{+}$ for controlling the upper bound.
3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$.
4:for $t=1$ to $T$ do
5:     Get weight estimator: $\theta_{t}\leftarrow A^{-1}b$
6:     Observe $K$ context vectors, $x_{t,1},x_{t,2},\dots,x_{t,K}\in\mathbb{R}^{d}$, generated through an LLM-based embedding model concatenated with user context information.
7:     for $k=1,2,\dots,K$ do
8:         $\hat{r}_{t,k}\leftarrow\theta_{t}^{\top}x_{t,k}+\alpha\sqrt{x_{t,k}^{\top}A^{-1}x_{t,k}}$ (Computes the upper confidence bound)
9:     end for
10:     Choose action $a_{t}=\arg\max_{k\in[K]}\hat{r}_{t,k}$.
11:     Observe reward $r_{t}$.
12:     $A\leftarrow A+x_{t,a_{t}}x_{t,a_{t}}^{\top}$.
13:     $b\leftarrow b+x_{t,a_{t}}r_{t}$.
14:end for
The algorithm proceeds as follows: During the LLM training phase, we prepare the matrix $A$ and vector $b$ to estimate the linear weight vector. We may need an LLM-based embedding model to extract embedding vectors from the original text information. Similar to LLM-2UCBs, we fine-tune the upper bound control parameter $\alpha$ using another fold of data and a bisection search to minimize regret on this fine-tuning dataset. The online learning phase remains the same as the standard LinUCB.
The advantages of LLM-LinUCB are its transparency due to the linear assumption, ease of execution with only one hyperparameter, and theoretical soundness with optimal regret. However, it is restricted to linear forms. There are two issues with this assumption. First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over embedding-based classification. Second, once concatenated with user contexts, it is unclear whether the linear form is a good approximation, necessitating further numerical study.
To alleviate the limitations of the linear assumption, we introduce another LLM-assisted contextual bandit algorithm (Algorithm 3), which extends the FAst Least-squares-regression-oracle CONtextual bandits (FALCON) algorithm proposed by (simchi2022bypassing). This type of contextual bandit algorithm generalizes the linear realizability assumption to a more flexible framework, assuming the true prediction model $f^{*}\in\mathcal{F}$. The functional class $\mathcal{F}$ can be very general and is not limited to linear models. This assumption is reasonable and promising because LLMs have demonstrated their capability across a wide range of tasks. Essentially, $\mathcal{F}$ can be viewed as a class of LLM models, and the goal is to find the true LLM model $f^{*}$ that performs best for our task, i.e., $f^{*}(x_{t,k},k)=\mathbb{E}[r_{t,k}|x_{t,k}]$ for any $k\in[K]$ and $x_{t,k}\in\mathcal{X}$. Under this setting, the context information is not restricted to real numbers and can include more general data, such as text information. We now present the LLM-FALCON algorithm.
Algorithm 3 LLM-Assisted FAst Least-squares-regression-oracle CONtextual bandits (LLM-FALCON)
1:LLM Training Phase: Obtain the historical data $\{(x_{t,k},r_{t})\}_{t=-M}^{0}$, and construct an LLM $\mathcal{F}$ class for training.
2:Hyperparameter Fine Tuning: Confidence control parameter $\delta$.
3:Online Learning Phase: Initialize $T$, $K$, and the epoch schedule $0=\tau_{0}<\tau_{1}<\tau_{2}<\cdots$.
4:for epoch $m=1,2,\dots$ do
5:     Set $\gamma_{m}=(1/30)\sqrt{(K\tau_{m-1})/\log(|\mathcal{F}|\tau_{m-1}/\delta)}$ (for epoch 1, $\gamma_{1}=1$).
6:     Compute $\hat{f}_{m}=\arg\min_{f\in\mathcal{F}}\sum_{t=-M}^{\tau_{m-1}}(f(x_{t,a_{t}},a_{t})-r_{t})^{2}$ via the offline least squares oracle.
7:     for round $t=\tau_{m-1}+1,\dots,\tau_{m}$ do
8:         Observe contexts $x_{t,k}\in\mathcal{X}$,$\forall k\in[K]$.
9:         Compute $\hat{f}_{m}(x_{t,k},k)$ for each arm $k\in[K]$. Let $\hat{a}_{t}=\arg\max_{k\in[K]}\hat{f}_{m}(x_{t,k},k)$.
10:         Define
$p_{t}(k)=\begin{cases}\frac{1}{K+\gamma_{m}(\hat{f}_{m}(x_{t,\hat{a}_{t}},\hat{a}_{t})-\hat{f}_{m}(x_{t,k},k))},&\forall k\neq\hat{a}_{t},\\ 1-\sum_{k\neq\hat{a}_{t}}p_{t}(k),&k=\hat{a}_{t}.\end{cases}$
11:         Sample $a_{t}\sim p_{t}(\cdot)$ and observe reward $r_{t}$.
12:     end for
13:end for
At the beginning of each epoch, we retrain the LLM using all available data up to that point, as detailed in Line 6. For simplicity, this algorithm assumes no optimization error during training, utilizing the offline least squares oracle instead. However, in practice, achieving zero optimization error is impossible, and one can incorporate this error into the final regret term using the regret triangle decomposition. Although this algorithm can handle original text information and fits into a general LLM model (not restricted to linear models), the primary challenge is the cost of training the LLM at each epoch because training a full LLM is typically time-consuming and costly. One potential solution, as investigated in our paper, is to fine-tune the LLM using LoRA instead of fully retraining it with new data. However, the empirical performance of this method remains unclear.
After getting an LLM trained on new data, then the key step in this algorithm is using the “inverse proportional to the gap” rule to balance exploration and exploitation. At each round, we first identify the empirically best arm $\hat{a}_{t}$ and then compute the gap in expected reward between this arm and the others. The sampling probability $p_{t}(k)$ is inversely proportional to this gap, meaning that a larger estimated reward $\hat{f}_{m}(x_{t,k},k)$ results in a smaller gap and a higher probability of being pulled. For a more detailed discussion of this algorithm, please refer to (simchi2022bypassing). Our purpose here is to adapt this algorithm in the LLM setting and discuss its pros and cons for completeness.
Appendix C LLM-Assisted Thompson Sampling
For completeness, we also present another version of LOLA that combines LLM and Thompson Sampling. We will update its numerical results in the next version of the manuscript.
Algorithm 4 LLM-Assisted Thompson Sampling (LLM-TS)
1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
2:Hyperparameter Fine-Tuning: Use another subset of data to select the LLM’s equivalent auxiliary sample size, $n^{\text{a}ux}$ and the standard TS prior $(\alpha^{0},\beta^{0})$.
3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$, LLM-based CTR prediciton $\bar{\mu}^{\text{a}ux}_{k}\leftarrow\mathcal{M}(x_{k})$, the parameters in Beta priors $(\alpha^{1}_{k},\beta^{1}_{k})\leftarrow(\alpha^{0},\beta^{0})+(n^{\text{a}ux}\bar{\mu}^{\text{a}ux}_{k},n^{\text{a}ux}(1-\bar{\mu}^{\text{a}ux}_{k}))$ for all arms $k\in[K]$.
4:for $t=1$ to $T$ do
5:     Sample $\theta_{k}^{t}\sim\text{beta}(\alpha^{t}_{k},\beta^{t}_{k})$ for all arms $k\in[K]$.
6:     Play the arm $a_{t}=\arg\max_{k\in[K]}\theta_{k}^{t}$.
7:     Observe the payoff $r_{a_{t}}$.
8:     Update $(\alpha^{t+1}_{k},\beta^{t+1}_{k})\leftarrow(\alpha^{t}_{k},\beta^{t}_{k})+(r_{a_{t}},1-r_{a_{t}})$ if $a_{t}=k$; update $(\alpha^{t+1}_{k},\beta^{t+1}_{k})\leftarrow(\alpha^{t}_{k},\beta^{t}_{k})$ if $a_{t}\neq k$.
9:end for
Appendix D Survey Results
To illustrate the difficulty of the headline selection task and compare it with LLM-based methods, we conducted a survey among students, asking respondents which headline was catchier.
Specifically, we conducted the survey study among undergraduate students, acknowledging that these participants were less experienced in editing than Upworthy editors and writers. Despite this limitation, we aim to highlight the challenges of this task for humans. It is important to note that all candidate headlines were chosen by Upworthy editors, who were uncertain which one was catchier, prompting the need for further A/B testing.
We designed two similar surveys using Qualtrics. The first survey randomly sampled 3,000 pairs from 39,158 headline pairs, differentiated by CTRs at a 0.05 significance level, along with 500 pairs showing no significant differences in CTRs. The second survey utilized the same 3,000 significant samples as the first. In both surveys, students were asked to select the catchier headline from each pair. The introductory text for the surveys was:
“Please help us understand what makes a news headline appealing to readers. In this survey, you will be presented with 20 pairs of headlines. Please choose the one that captures your attention more effectively or feels more engaging in each pair, assuming both headlines relate to the same news content. There are no right or wrong answers; we are simply interested in your opinions.”
After providing identification information, including the last four digits of their phone and student ID, respondents proceeded to a new page featuring 20 headline comparison questions with the prompt, “Please select the headline that you find more attractive,” with only one permissible response. Notably, the first survey included an additional response option: “Both headlines seem equally good.” At the survey’s conclusion, we asked, “How often do you read the news?” with options “Never”, “1-2 days per week”, “3-4 days per week”, “5-6 days per week”, “Everyday”.
Detailed information and survey results are recorded in Table A1. Because some students did not finish the survey or skipped some comparisons, one can see there are 3,809 responses out of $202\times 20=4,040$. To calculate the accuracy metric, we only focused on the significant pairs, and in the first survey, we ruled out the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys. “Accuracy” measure here is consistent with the LLM tests. We viewed the headline with a higher CTR as the correct answer, and calculated the proportion of correct answers as accuracy. By doing so, one should expect 50% accuracy by random guess since we randomly swapped the order of the headlines to make sure the displayed order does not correlate with the correct answer.
Detailed information and survey results are documented in Table A1. Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$). To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys. Accuracy here aligns with the LLM tests. We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy. Accordingly, one might expect a 50% accuracy rate from random guessing, as we randomized the order of the headlines to ensure that the display order did not correlate with the correct answer.
The survey results indicate that respondents’ ability to identify catchier headlines is comparable to random guessing, with no accuracies significantly deviating from 50% at the 0.05 significance level. Furthermore, accuracies did not vary significantly across different groups based on news reading frequency. We also incorporated two attention checks in the second survey. These findings suggest that the human ability to discern catchier headlines is limited, highlighting the intrinsic difficulty of the task.
Survey 1 Survey 2
conducted time 4/14 - 4/21 4/29 - 5/17
# of respondents 202 339
headline pair pool 3000 significant + 500 insignificant pairs 3000 significant pairs
# of sampled pairs in survey 20 20
# total received responses 3,809 4,571
overall accuracy % 45.22 51.58
Results based on different answer to “How often do you read the news?”
Never proportion % 21.78 16.22
accuracy % 45.76 51.62
1-2 days per week proportion % 48.51 49.56
accuracy % 45.36 52.54
3-4 days per week proportion % 18.81 18.29
accuracy % 46.83 51.39
5-6 days per week proportion % 4.95 5.31
accuracy % 39.05 50.15
Everyday proportion % 5.94 8.26
accuracy % 42.35 47.48
Table A1: Survey Response Analysis. No accuracies in this table are significantly lower or higher than 50% at 0.05 significant level. 8 respondents did not finish Survey 2 in the limited given time, so the summation of proportion from different subgroups is smaller than 100% in Survey 2.

### v2

@@
 This paper introduces the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery.
-Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests aimed at evaluating the performance of various headlines associated with the same article content, we first investigate three broad pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs.
-Our findings indicate that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline among two options.
-In contrast, OpenAI-embedding-based classification models and fine-tuned Llama-3-8b models achieve comparable accuracy, around 82-84%, though still falling short of the performance of experimentation with sufficient traffic.
-We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to adaptively allocate traffic and maximize clicks.
-Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B testing method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic or numerous arms.
-Our approach is both scalable and broadly applicable to content experiments across a variety of digital settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
-Keywords: LLMs, Content Experiments, Bandits, News
+Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests, we first investigate three pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs.
+We find that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline.
+In contrast, both OpenAI-embedding-based classification models and fine-tuned Llama-3 with 8 billion parameters achieve an accuracy of around 82-84%.
+We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to allocate traffic and maximize clicks adaptively.
+Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B test method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic.
+Our approach is scalable and applicable to content experiments across various settings where firms seek to optimize user engagement, including digital advertising and social media recommendations.
+Keywords: LLMs, Content Experiments, Bandits, News, Digital marketing.
 1 Introduction
@@
 Traditionally, media firms and publishers have relied on experimentation-based approaches to address this problem.
-Broadly speaking, there are two general types of experimentation styles—(1) Standard A/B tests and (2) Online learning algorithms or bandits.
+Broadly speaking, there are two types of experimentation styles—(1) Standard A/B tests and (2) Online learning algorithms or bandits.
 [^1] We use the terms adaptive experimentation, online learning algorithm, and bandits interchangeably throughout the paper.
-We now discuss each briefly.
-A/B testing is the most straightforward experimentation method, where a firm allocates a fixed portion of traffic to different treatment arms, assesses the results, and then goes with the best-performing arm.
+A/B test is the most straightforward experimentation method, where a firm allocates a fixed portion of traffic to different treatment arms, assesses the results, and then goes with the best-performing arm.
 This is also known as the Explore and Commit strategy, or E&C, where the firm explores for a fixed period using A/B tests and then commits to one treatment based on the results of those tests.
@@
 For instance, The New York Times recently built a centralized internal A/B test platform, ABRA (A/B Reporting and Allocation architecture), allowing different teams to experiment with their content; see (NYTimes2024) for more details.
-Another example is Upworthy, which has implemented extensive A/B testing to refine its headlines (matias2021upworthy).
-Upworthy editors create several headlines for the same article and randomly assign readers to see different headlines with equal probability for a fixed amount of traffic, and then show the best-performing headline to all the users.
+Another example is Upworthy, which has implemented extensive A/B tests to choose headlines for news articles (matias2021upworthy).
 The main advantage of this approach is that it is trustworthy—with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased and accurate.
@@
 This method provides a more efficient way to optimize content delivery by continually learning and adapting.
-For instance, both Yahoo News and The New York Times have successfully implemented contextual bandit algorithms to identify content that maximizes user engagement.
-Yahoo was one of the earliest platforms to use the linear contextual bandit for news article recommendation (li2010contextual).
-More recently, The New York Times has also adopted a similar algorithm to power its article recommendation system (NYTimes2019).
+For instance, both Yahoo News and The New York Times have successfully implemented contextual bandit algorithms to identify content that maximizes user engagement (li2010contextual; NYTimes2019).
 Online learning algorithms typically outperform A/B tests in terms of overall regret or reward.
@@
 As a result, they tend to switch to the best-performing arms quickly and incur lower regret compared to standard E&C strategies.
-This has been established both theoretically (lattimore2020bandit) as well empirically.
-For example, (li2010contextual) demonstrate that the linear contextual bandit algorithm improves the total clicks on news articles over the random policy (i.e., A/B testing with equal probability for arms) at Yahoo News using off-policy evaluation on field data.
+This has been established both theoretically (lattimore2020bandit) as well empirically (li2010contextual).
 [^2] Similarly, in the context of digital advertising, (schwartz2017customer) use field experiments to demonstrate that adaptive algorithms outperform the standard balanced A/B test.
-Nevertheless, these online algorithms have notable drawbacks.
-In particular, in standard content experiments, these algorithms typically suffer from a cold-start problem.
-That is, firms typically start these adaptive experiments under the assumption that all have an equal probability of being the best-performing one.
-This can lead to a situation where more-than-necessary traffic is allocated to sub-optimal arms and higher regret as less appealing content is shown to users during the exploration phase.
-Additionally, due to their probabilistic nature, these algorithms can inadvertently converge on a suboptimal arm, especially when the experimental traffic is small, increasing the probability of identifying a suboptimal arm as the best one.
-These issues arise because online algorithms are designed for large traffic regimes, where the experimenter does not care too much about initialization since it has little impact on the asymptotic regret.
-However, as discussed earlier, in the news industry, there is limited traffic available for experimentation, given that news content becomes stale quickly.
+Nevertheless, online algorithms also have notable drawbacks.
+In standard content experiments, these algorithms suffer from a cold-start problem.
+That is, firms typically start these adaptive experiments assuming that all arms are equally likely to be the best-performing ones.
+This can lead to a situation where more-than-necessary traffic is allocated to sub-optimal arms, leading to higher regret.
+Additionally, due to their probabilistic nature, these algorithms can inadvertently converge on a sub-optimal arm, especially when the experimental traffic is small.
+These issues arise because online algorithms are designed for large traffic regimes, where the initialization of arms has little impact on the asymptotic regret.
+However, as discussed earlier, there is limited traffic available for experimentation in the media industry since content becomes stale quickly.
 Therefore, the negative effects of the “equally effective” initialization do not diminish sufficiently during the time horizon of interest, leading to higher regret.
 Given these issues with content experimentation, the main question that this paper asks and answers is whether we can leverage the recently developed Large Language Models (LLMs) to enhance current content experimentation practices and simplify the process of identifying appealing content.
-LLMs, trained on extensive human-generated data, have demonstrated the ability to mimic human preferences and behavior in various contexts.
-For example, (li2024frontiers) investigate whether LLM-generated survey responses produce perceptual maps similar to those generated by human survey responses in car branding analysis.
-Their study shows a high agreement rate of over 75%.
-More specifically, in the context of news articles, (yoganarasimhan2024feeds) show that LLMs are effective at understanding the polarization of news content.
-Hence, a relevant question is whether news publishers can directly use LLMs to predict the appeal of different content, potentially replacing traditional experimentation methods.
+LLMs, trained on extensive human-generated data, have demonstrated the ability to mimic human preferences and behavior in a variety of consumer research tasks (li2024frontiers; brand2023using).
+Specifically, in the context of news articles, (yoganarasimhan2024feeds) show that LLMs can correctly predict the polarization of news content.
+Hence, a relevant question is whether firms can directly use LLMs to predict the appeal of different content, potentially replacing traditional experimentation.
 This approach could significantly reduce the costs associated with experimentation while simplifying the content delivery processes.
-To address this question, we leverage a large-scale dataset on content experiments by the publishing firm Upworthy conducted in the 2013–2015 time period.
-The dataset consists of 17,618 headline tests spanning 77,245 headlines, over 277 million impressions, and over 3.7 million clicks.
+To address this question, we leverage a large dataset of content experiments by the publishing firm Upworthy conducted in the 2013–2015 timeframe.
+The dataset consists of 17,618 headline tests spanning 77,245 headlines, 277 million impressions, and 3.7 million clicks.
 The goal of each A/B test was to identify the best headline among a set of candidate headlines (for a given article).
-This dataset is ideal for our study since it was collected from A/B tests where each headline received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance.
-Such a high-quality dataset provides a trustworthy ground for comparing different approaches.
-Additionally, since the dataset was sourced from the real media industry, our methods can reflect real-world scenarios and offer practical guidance to both researchers and practitioners.
-In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can predict headline attractiveness accurately.
+This dataset is an excellent test-bed for our study for a few reasons – (1) it is based on an extremely large number of A/B tests over a high volume of impressions, (2) each A/B test received a relatively large number of impressions, allowing for accurate estimates of each headline’s performance, and (3) it is sourced from a real media firm and represents the actions of a very large number of readers over a sufficiently long period to offer meaningful insights.
+In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can accurately predict headline attractiveness.
 For this exercise, we focus on pairs of headlines (within a given A/B test) whose CTRs are significantly different (based on the observed data).
@@
 Thus, this serves as a preliminary testbed to evaluate whether pure-LLM-based approaches can predict which of the two headlines is catchier (or clickable).
-We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM-based text-embedding approaches, and (3) Fine-tuning approaches.
-We discuss each approach and its performance next.
-First, for prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning.
+We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM text-embedding approaches, and (3) Fine-tuning approaches.
+For prompt-based methods, we consider two approaches: (a) Zero-shot prompting and (b) In-context learning.
 Zero-shot prompting is a technique in which an LLM generates responses without being explicitly trained on specific examples.
 This is the most common way in which most users interact with the LLM.
-In-context learning involves providing the LLM with several demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation).
+In-context learning involves providing the LLM with a few demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation).
 We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches, with accuracy hovering around 50%.
@@
 We then concatenate two embeddings as inputs, along with the winner with higher CTR as the output, to train binary classification models, including logistic regression and multilayer perceptron (a type of neural network).
-We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82% accuracy, while more complex neural networks do not improve performance over the simple logistic model.
+We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82.50% accuracy, while more complex neural networks do not improve performance over the simple logistic model.
 Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora).
-This approach performs the best, with 83% accuracy, which is a small improvement over the embedding-based approach.
-[^3] On the other hand, it also marginally increases training costs and engineering complexity.
-Overall, we find that prompt-based approaches, which have been the standard method in recent marketing literature for using LLMs so far, perform poorly.
-In contrast, models based on LLM embeddings or fine-tuned open-source LLMs offer significantly more accurate predictions.
-The primary advantage of LLM-based methods is that predictions can be made without wasting any traffic.
-Nonetheless, even the best LLM-based models do not match the accuracy of experimentation-based approaches (with sufficient traffic).
+This approach performs the best, with 83.40% accuracy, which is a small improvement over the embedding-based approach.
+However, even the best-fine-tuned LLMs are unable to perfectly generate/predict which types of content are most appealing to users.
+As such, none of the LLM-based approaches can match the accuracy of experimentation-based approaches.
 Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm).
-Our key insight is to use predictions from fine-tuned LLM models as priors in online algorithms to optimize experimentation.
+Our key insight is to use predictions from an LLM model as priors in online algorithms to optimize experimentation.
 The approach involves two steps.
-First, we train an LLM-based CTR prediction model, which can be an LLM-embedding-based model or a fine-tuned open-source LLM model, or indeed any other LLM-based approach.
+First, we train an LLM-based CTR prediction model using LoRA fine-tuning.
 The second step integrates these predictions into online learning algorithms.
 LOLA is a general framework, and depending on the specific goals of the experimentation, it can be adapted for different prediction models in the first step and different online algorithms in the second step.
-For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs), detailed in Section 4.1.
+For example, if the goal is to minimize regret (i.e., maximize total reward/clicks), we propose a modified Upper Confidence Bound algorithm called LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) that is detailed in §$\S$4.2.
 The idea of LLM-2UCBs is intuitive: we can view LLM-based CTR predictions as auxiliary samples before the start of the online algorithm (e.g., a 1% CTR can be viewed as 100 impressions with 1 click or 1,000 impressions with 10 clicks).
-After fine-tuning this auxiliary impression size hyperparameter, we incorporate these imaginary samples into the standard UCB algorithm to get the reward estimator and shrink the upper confidence bound.
+After fine-tuning this auxiliary impression size hyperparameter, we incorporate these auxiliary samples into the standard UCB algorithm (auer2002finite) to obtain the reward estimator and shrink the upper confidence bound.
 LOLA combines the advantages of both experimentation and LLMs, making it accurate and efficient.
-By using LLMs, we reduce the waste of impressions on poor-performing headlines, especially in the initial stage when we face a cold-start problem and lack good prior information about the headlines’ performance.
-On the other hand, through experimentation, we can correct the errors in LLMs.
-Even though LLMs achieve around 82% accuracy, there are still 18% of incorrect cases.
-This hybrid approach ensures that while LLMs provide reasonable initial predictions, ongoing experimentation refines and improves overall performance, making LOLA an effective solution for content optimization.
-We present a comprehensive evaluation of LOLA and compare its performance with three natural benchmarks using the Upworthy dataset: (1) Explore and Commit, (2) pure online learning algorithm, and (3) pure LLM-based model.
-The first benchmark, E&C, was the status quo at Upworthy during the time of data collection, wherein the editors first ran an A/B test on a set of headlines for a fixed set of impressions, and then displayed the winner for the rest of the traffic.
+By integrating an LLM-based model to predict CTRs before the online deployment phase, we avoid wasting impressions on poorly performing headlines, especially in the initial stage when we face a cold-start problem.
+On the other hand, through experimentation, we can correct the errors in LLM predictions.
+This hybrid approach ensures that while LLMs provide reasonable initial predictions, ongoing experimentation refines and improves overall performance, making LOLA an effective solution for optimizing content delivery.
+We present a comprehensive evaluation of LOLA and compare its performance with three natural benchmarks using the Upworthy dataset: (1) Explore and Commit, (2) a pure online learning algorithm, and (3) a pure LLM-based model.
+The first benchmark, E&C, was the status quo at Upworthy during the time of data collection, wherein the editors first ran an A/B test on a set of headlines for a fixed set of impressions and then displayed the winner for the rest of the traffic.
 For the pure online learning algorithm, we use the standard Upper Confidence Bound (UCB) algorithm, which is initialized with uniform CTRs.
-For the pure LLM-based approach we use the linear CTR prediction model in conjunction with OpenAI embeddings.
-We compute CTRs for all headlines and select the one with the highest CTR for all impressions, bypassing the experimentation phase.
-[^4] We chose the embedding-based approach for the evaluation instead of the fine-tuned model, given its ease of use and almost comparable performance to the fine-tuned open-source LLM.
-In LOLA, we use the same CTR prediction model as in the pure LLM approach and employ LLM-2UCBs to maximize accumulated clicks.
-Our numerical comparison primarily considers the regret minimization framework, as minimizing regret in Upworthy’s context equates to maximizing total clicks over all impressions, including those for experimentation.
-This metric is pertinent to business profitability since Upworthy, as a free online publisher without subscription fees, relies heavily on advertising revenue.
-Increased user clicks enable Upworthy to display more advertisements, thereby increasing revenue.
-In our empirical evaluation, we find that LOLA outperforms all the baseline approaches.
-We find that for the earlier impressions, both LOLA and the pure LLM-based approach beat the pure experimentation-based approaches (i.e., the standard A/B test and UCB algorithm) due to relatively accurate LLM-based CTR predictions.
-However, as the number of impressions grows, the performance of the pure LLM-based approach declines rapidly, as it does not learn anything from past impressions and, as such, cannot improve its performance.
-Thus, its accuracy falls short compared to other experimentation-based methods.
-Thus, as discussed earlier, our experiments demonstrate that LOLA is able to leverage the power of LLMs early in the horizon and then build on it to take advantage of experimentation over time, thereby bringing together the strengths of both methods.
-Finally, we observe that the performance gap between LOLA and the standard UCB narrows, given that both methods are asymptotically optimal and the benefit provided by LLM diminishes over time.
-Thus, we expect that LOLA will be especially valuable in settings where the traffic available for experimentation is limited and/or the number of arms is large.
-Nevertheless, LOLA continues to outperform standard UCB, even after very long horizons.
+For the pure LLM-based approach, we use the LoRA fine-tuned Llama-3 to predict the CTRs of headlines and select the one with the highest CTR for all impressions, bypassing the experimentation phase.
+In LOLA, we use the same CTR prediction model as that used in the pure LLM approach and then employ LLM-2UCBs to maximize accumulated clicks.
+We find that LOLA outperforms all the baseline approaches, though the next-best algorithm varies based on the length of the time horizon (keeping the number of headlines fixed).
+Specifically, when time horizons are small, both LOLA and the pure LLM-based approach beat the pure experimentation-based approaches (i.e., E&C and the UCB algorithm) due to relatively accurate LLM-based CTR predictions, and lack of sufficient time for the experimentation to provide meaningful updates.
+Specifically, both LOLA and the pure LLM outperform E&C by 8-9% and UCB by 6-7% in this scenario.
+However, as the number of impressions/time horizon grows, the performance of the pure LLM-based approach falls behind other experimentation-based methods because pure LLM-based approaches are static and do not take advantage of experimentation.
+In this case, LOLA beats pure LLM methods and E&C by 4–5%, and UCB by 2-3%.
+In sum, our experiments demonstrate that LOLA is able to leverage the power of LLMs early in the horizon and then build on it to take advantage of experimentation over time, thereby bringing together the strengths of both methods.
+Finally, while we focus on regret minimization in a stochastic bandit setting given the Upworthy context, LOLA is a general-purpose framework and can easily accommodate different goals and/or settings.
+We provide variants of the standard LOLA (and results on the empirical performance of these variants when applicable) for a few natural extensions, e.g., for Bayesian bandit settings where the second stage algorithm can be Thompson Sampling instead of UCB, for the best arm identification problem where the goal is to identify the best content/headline rather than regret minimization, and for settings where the firm has access to user/contextual features that can be used to personalize content delivery.
 In summary, our paper makes two key contributions to the literature.
@@
 To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques.
-We demonstrate the value of our approach using a large-scale dataset from Upworthy, showcasing significant improvements in total clicks and profitability.
-Second, from a managerial perspective, LOLA can be used in various settings where firms need to decide which content to show, such as news articles, headlines, or even advertisements and website content.
-This method is based on open-source LLM models, making it cost-effective and easily deployable across different domains once fine-tuned on relevant datasets.
-Compared to the substantial benefits gained from improved content optimization, the cost of fine-tuning these models is negligible.
-Because this fine-tuning process on the open-sourced LLMs can be efficiently executed even on entry-level GPUs, making it accessible for a wide range of applications.
-This accessibility ensures that even firms with limited budgets can implement advanced LLM-based techniques to enhance their content strategy.
-More broadly, while we focused mainly on the news/publishing industry, our framework is general and can easily extend to other settings where firms need to evaluate which type of content maximizes user engagement, including online/digital advertising and social media platforms.
+We demonstrate the value of our approach using a large-scale dataset from Upworthy when compared to baseline methods.
+Second, from a managerial perspective, LOLA can be used in a broad range of settings where firms need to decide which content to show.
+While we focus on the news/publishing industry given our empirical context, the LOLA framework is general and can easily extend to other settings, including digital advertising, email marketing, and website design.
+Further, the method is cost-effective, can be built on open-source LLMs (alleviating data privacy concerns), and is easily deployable across different domains once fine-tuned on relevant datasets.
 2 Upworthy Data and Experiments
-We now discuss data sourced from Upworthy, a U.S.
-media publisher known for its innovative use of A/B test in online media publishing.
-In a bid to increase clicks to articles, Upworthy conducted randomized experiments with each article published, exploring different combinations of headlines and images to determine which elements most effectively engaged viewers.
-The archival record from January 24, 2013, to April 30, 2015, detailed in (matias2021upworthy), demonstrates how strategic packaging of headlines, subheadings, and images played a crucial role in Upworthy’s growth strategy.
-Such A/B testing practices are also widely used in the media and publishing industry, technology industry, and political campaigns.
-A/B testing allows firms to assess and refine the impact of various editorial and content choices (e.g., headlines or text of an article/email) systematically and automatically.
 Test ID Headline Impressions Clicks
@@
 We display the key columns for our analysis, including the ID for the A/B tests, the text of headlines, and the number of impressions and clicks in the test.
-During the experiment period, Upworthy’s editorial team created several versions of headlines and/or images for each article (internally called “package” by Upworthy).
+We now discuss the data sourced from Upworthy, a U.S.
+media publisher known for its extensive use of A/B tests in digital publishing.
+Upworthy conducted randomized experiments with each article published, exploring different combinations of headlines and images to determine which elements most effectively led to higher clicks.
+The archival record from January 24, 2013, to April 30, 2015, detailed in (matias2021upworthy), demonstrates how the packaging of headlines and images played a crucial role in Upworthy’s growth strategy.
+Upworthy’s editorial team created several versions of headlines and/or images for each article (internally called “package” by Upworthy).
 A package is defined as one treatment or arm for an article and consists of a headline, image, or a combination of both.
@@
 Then, they A/B tested these packages or treatments to identify which one resonated the best with their audience.
-Specifically, Upworthy’s A/B testing system recorded how many impressions and clicks each package received.
-Key metrics from these tests were shown via a dashboard, which guided decisions on whether to adjust the content further or finalize a particular version.
-Table 1 shows two examples of A/B tests, with the relevant columns.
+During the A/B test, users only saw the headline (and an accompanying image in some cases), but not the article itself.
+Upon clicking on the headline, they were taken to the article.
+As such, the content of the article itself did not have any effect on users’ clicking behavior.
+Upworthy’s A/B test system recorded how many impressions and clicks each package received.
+Table 1 shows two examples of A/B tests with the relevant columns.
 The full details of all the A/B tests and their results are available from the Upworthy Research Archive.
@@
 Therefore, we focus on the headline tests and filter out all the image tests.
-After this filtering, there are 17,681 tests comparing headlines, totaling 77,245 packages.
+After this filtering, there are 17,681 tests, totaling 77,245 packages.
 On average, each article was tested with 4.37 headlines.
@@
 We also report the total number of impressions for the entire Upworthy website, as recorded in their data from Google Analytics.
-It is noteworthy that the impressions attributable to the tests constitute only 22.9% of the total traffic.
-This relatively small proportion of traffic can be attributed to Upworthy’s policy of conducting no more than one test per webpage to avoid interference effects.
+We see that the impressions attributable to the tests constitute only 22.9% of the total traffic.
 # of headlines in one test # of tests % of samples
@@
 To examine whether these A/B tests were informative, or not, we generate pairs of headlines and conduct pairwise t-tests on the CTRs of the headlines.
-For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^5] It is important to note that the issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives.
+For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^3] The issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives.
 In our dataset, tests involving 10 pairs or less (equivalent to no more than 5 packages) account for over 80% of all tests.
 The risk of multiple testing is relatively low when conducting fewer than 10 tests, and conventional significance levels, e.g., $\alpha=0.05$ are generally adequate to account for this risk.
-To ensure the robustness of our findings, we conducted an additional analysis by excluding tests with more than 5 packages.
-The comparisons among pure-LLM-based methods remained consistent because all methods were conducted on the same dataset.
-It is important to note that a higher risk of multiple testing results in a larger number of actually insignificant pairs within our significant subset, which may only decrease the prediction accuracy.
-In other words, the results we present using the significant pairs may underestimate the prediction accuracy.
-Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in Section 4, we utilize all available data without concern for multiple testing.
+Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in §$\S$4, we utilize all available data without concern for multiple testing.
 see Table 4 for examples of such tests for the two A/B tests highlighted earlier in Table 1.
-After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level, and 23,682 at a 0.01 level.
-We do not observe a fundamental difference when conducting our experiments with 0.05 or 0.01 significant level pairs; therefore, for simplicity, we primarily use 0.05 level data in the main analysis in Section 3.
+After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level and 23,682 at a 0.01 level.
 In summary, we find that only 28% of the tests were significant at the 0.05 level, and a mere 17% are significant at the 0.01 level.
 Overall, these results suggest that most pairs of headlines tend to perform similarly, and the task of learning which headline is catchier is non-trivial.
+To further illustrate the difficulty of this task, we conduct a survey, where we give human users pairs of headlines from the set of headline pairs that are significantly different from each other and ask them to identify the catchier headline.
+Even within this set of headline pairs, we find that the respondents’ accuracy in the survey was not significantly better than random guessing.
+This further highlights the fact that the task of identifying engaging content is inherently challenging, even for humans.
+See Web Appendix §$\S$§A for more details of the survey and its results.
 Test ID Headline 1 Headline 2 CTR 1 CTR 2 p value Winner
@@
 3 Pure-LLM-Based Methods
-In this section, we analyze 39,158 headline pairs, representing 28.05% of the total, at a 0.05 significance level.
+In this section, we consider the 39,158 headline pairs that significantly differ from each other at the 0.05 significance level.
 We focus on these significant pairs for three reasons: First, insignificant pairs suggest that decisions in final headline selection do not substantially affect users’ click behavior, implying that accuracy in predicting outcomes for “equally good” samples is not critically important from a business perspective.
-Second, to avoid information overload in our survey study, we limit our focus to significant pairs for a fair comparison between LLMs and humans.
-Third, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in Section 4.
+Second, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in §$\S$4.
 As such, relative comparison across methods is more important as long as we use the same dataset for all methods.
-Furthermore, we note that the methods discussed in this section can be readily adapted from binary to multi-class classification tasks.
-Indeed, in Section 4, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
+Third, as demonstrated in Web Appendix §$\S$§D.2.1, our robustness checks confirm that including insignificant pairs in the training phase does not affect the overall performance of the methods considered.
+Furthermore, we note that all methods discussed in this section can be readily adapted from binary to multi-class classification tasks.
+Indeed, in §$\S$4 where we introduce our LOLA approach, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
 In the rest of this section, we investigate the performance of three widely used LLM-based approaches—(1) GPT prompt-based approaches, including zero-shot prompting and in-context learning, (2) embedding-based classification models, and (3) fine-tuning of LLMs.
 3.1 Prompt-based Approaches
-In this section, we consider prompt-based approaches to evaluate which headline among a pair of headlines is catchier.
-These approaches simply use GPT prompts to learn which headline is catchier, and there is no further modeling or explicit fine-tuning involved.
-Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far .
-We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning. We discuss both in detail below.
+Prompt-based approaches use GPT prompts to learn which headline is catchier, and there is no modeling or explicit fine-tuning involved.
+Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far and have been used to examine if and when LLMs can simulate consumer behavior in market research studies (gui2023challenge; brand2023using; li2024frontiers).
+We consider two prompt-based approaches: (1) Zero-shot prompting and (2) In-context learning.
 3.1.1 Zero-shot Prompting
 Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly trained on specific examples.
-In this study, we evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models.
+We evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models.
 We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses.
@@
 The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation.
-Additionally, we set the default temperature parameter of 1.666The temperature parameter scales and controls the randomness of the model’s responses.
-GPT model generates responses token by token.
-Tokens are common sequences of characters found in a set of text.
-A helpful rule of thumb is that one token generally corresponds to about four characters of text in common English, translating to roughly three-quarters of a word (so 100 tokens $\approx$ 75 words), with each token sampled from a token distribution provided by the model.
-The temperature parameter determines whether this distribution is flattened or sharpened.
+Additionally, we set the default temperature parameter of 1.444The temperature parameter scales and controls the randomness of the model’s responses.
 A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused.
 For simplicity, we use the default setting with temperature as 1.0.
-We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy/performance.
+We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy.
 Figure 1: Zero-Shot Prompting for Headline Selection.
-In our setting, we design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1.
+We design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1.
 We test this prompt using three OpenAI large language models, [^5] As recommended by OpenAI, see this page, accessed on May 24, 2024, one can default to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o.
-OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence., each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters, enhancing its capabilities.
+OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence.
+each differing in parameter size: (1) GPT-3.5 (GPT-3.5-turbo-0125), which uses around 175 billion parameters, (2) GPT-4 (GPT-4-turbo-2024-04-09) which uses approximately 1.7 trillion parameters, and (3) GPT-4o (GPT-4o-2024-05-13), which uses an even larger number of parameters.
 The goal here is to investigate whether improvements in model architecture and parameter size can improve the accuracy of our predictions via standard prompts.
 3.1.2 In-context Learning
+Figure 2: In-Context Learning Prompt for Headline Selection.
 In-context learning is a technique where a model is provided with demonstrations within the input prompt to help it perform a task.
@@
 They provide empirical evidence and theoretical proofs showing that both Transformers and LSTMs can exhibit in-context learning on synthetic datasets.
-[^8] More generally, beyond LLMs, (garg2022can) investigate the capacity of Transformers to in-context learn various function classes, demonstrating that these models can be trained from scratch to learn linear, sparse linear, two-layer neural networks, and decision trees, even under distribution shifts between training data and inference-time prompts.
-We design the in-context learning prompt shown in Figure 2.
+We design in-context learning prompts as shown in Figure 2.
 Our prompts include demonstrations of catchier headlines and ask the model to determine which headline is catchier for a subsequent pair of headlines.
@@
 A few demonstrations are typically sufficient to guide the model.
-Improvement tends to have diminishing returns quickly; adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking).
+Improvement tends to have diminishing returns quickly, i.e., adding more demonstrations does not significantly enhance performance after a certain point (min2022rethinking).
 In our numerical experiments, we use two or five demonstrations to investigate the impact of the number of examples on performance.
 [^6] We did run experiments with a larger number of demonstrations, but did not find any significant improvements from doing so; and hence do not report them here.
-Figure 2: In-Context Learning Prompt for Headline Selection.
 An interesting aspect of in-context learning is the model’s robustness to incorrect labels in the demonstrations.
@@
 Second, the GPT models can produce different answers even when given the exact same prompt.
-To mitigate this variability, we randomly select 1,000 significant pairs and conduct all our experiments on this same test set.
+To mitigate this variability, we randomly select 1,000 significant pairs of headlines and conduct all our experiments on this same test set.
 Similarly, all the demonstrations are also randomly sampled and fixed for all experiments to reduce the variance of prompt outputs.
 The results of all the prompt-based experiments are shown in Table 5.
-A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix A.1.
-The second column of 5, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting.
+A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix §$\S$§B.1.
+The second column of Table 5, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting.
 Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with flipped outcome labels.
@@
 Neither the number of demonstrations used nor the label type have any significant impact on the accuracy of the GPT-3.5 predictions.
-Interestingly, we find zero-shot prompting using GPT-3.5 is significantly better than the random prediction, although by only 4.20%, and marginally outperforms the in-context learning.
+Interestingly, zero-shot prompting using GPT-3.5 is slightly better than random prediction and also marginally outperforms in-context learning.
 In comparison, experiments with GPT-4 have accuracy rates around 60-65%, which is a significant improvement in performance compared to GPT-3.5.
@@
 This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
-Finally, we also report the monetary cost and total run-time of the prompt-based experiments.
-The cost is \$1.13 for GPT-3.5 with 1,000 pairs and \$20.11 for GPT-4, which is approximately 20 times the cost of GPT-3.5.
-The total run-time is 121.5 minutes for the entire experiment, which involved testing a total of 10,000 prompts.
-[^10] Note that the run-time may depend on the API usage tiers, which provide different rate limits.
-In summary, we find that prompt-based methods are relatively slow (compared to other methods introduced later in this section), have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), and high monetary cost.
+In summary, we find that prompt-based methods have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), slow running time (compared to other methods introduced later in this section), and high monetary cost (see Web Appendix §$\S$§E for a detailed discussion of costs and running time).
 All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
 3.2 Classification Models with LLM-based Text Embeddings
-In this section, we consider another LLM-based method for this prediction task.
+We now consider another LLM-based method for the prediction task.
 Specifically, we leverage the text embeddings from OpenAI and utilize classification models, including Logistic Regression (Logit) and Multilayer Perceptron (MLP), to predict which of the headlines in a pair is likely to generate a higher CTR.
-The Logit model is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables.
-On the other hand, MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data.
-We refer interested readers to Chapter 5 in (zhang2021dive) for more details.
-3.2.1 Text Embedding Technique
-Text embedding is an advanced technique in natural language processing (NLP) that transforms larger chunks of text, such as sentences, paragraphs, or documents, into numerical vectors.
-Such vector forms a representation of the semantic information of the input text, and can be further utilized by downstream machine learning models.
-Models like BERT (devlin2018bert), Sentence-BERT (reimers2019sentence), and OpenAI’s embedding models (openai-embedding) (text-embedding-3-large for example) are viewed as state-of-the-art embedding models at the date of their release.
-Essentially, text embedding is a numerical representation of text that captures the rich information in the text.
-These embedding vectors are used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which determines the emotional tone of a piece of text.
-It is important to distinguish between embedding models and standard LLMs used for prompts, e.g., GPT.
-While they are based on similar underlying Transformer architectures and extensive pretraining on large text corpora, they are optimized for different tasks.
-Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts, while embedding models focus on producing high-quality embeddings for downstream tasks.
-For a more detailed survey of the most recent text embedding models (patil2023survey).
-In the remainder of this section, we leverage the embedding model from OpenAI to convert headline texts into numerical embedding vectors, which are then used to train the downstream binary classification models to identify which headline is better.
+Text embedding techniques transform large chunks of text, such as sentences, paragraphs, or documents, into numerical vectors.
+These embedding vectors can then be used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which is used to determine the emotional tone of a piece of text.
+[^7] Embedding models and standard prompt-based LLMs used share some common features, e.g., the underlying Transformer architectures and extensive pretraining on large text corpora.
+However, they are optimized for different tasks.
+Models like GPT are optimized for generating coherent and contextually relevant text based on input prompts while embedding models focus on producing high-quality embeddings for downstream tasks.
+See patil2023survey for a more detailed survey of the most recent text embedding models and their applications.
+An interesting question is why text embeddings obtained from LLM models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for downstream classification tasks.
+(saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text.
+Thus, the embeddings, which can be roughly viewed as the next token probability generated from a next-word prediction task, inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
 3.2.1 Implementation Details
-In our numerical study, we employ the embedding model from OpenAI to facilitate a fair comparison with GPT-based prompts.
-At the time of this study, the latest and best-performing embedding model available was text-embedding-3-large.
-This model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021.
-A notable feature of this model is its flexibility, allowing users to tailor the embedding dimension to their specific needs.
-[^11] Other models capable of text embedding, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained.
+Figure 3 illustrates the conceptual framework of our approach here.
+We concatenate the embeddings of two headline embeddings888 Because embeddings approximate the next token probability, one can also use the difference (subtraction) of two embedding vectors instead of the concatenation as the input to the downstream classification model.
+In our task, these two input types have no significantly different impact on the final classification accuracy.
+and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR.
+We consider three embedding models:
+  • $\bullet$
+    OpenAI embeddings: At the time of this study, the latest and best embedding model available from OpenAI was text-embedding-3-large.
+These embeddings are obtained using the OpenAI embedding API in Python.
+This embedding model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021.
+A notable feature of this model is its flexibility – it allows users to tailor the embedding dimension to their specific needs.
+[^9] Other models capable of producing text embeddings, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained.
 These dimensions cannot be changed or resized for different applications.
@@
 This sparsity can increase the possibility of over-fitting.
-Figure 3 illustrates the conceptual framework of our approach here.
-First, we call the OpenAI embedding API in Python to obtain text embeddings for the headlines based on text-embedding-3-large.
-We then concatenate the embeddings of two headline embeddings and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR.
-To assess the impact of model complexity on performance, we test two classification models: Logit and MLP.
-The embedding model parameters remain fixed during training, while only the classification model parameters are updated.
-Figure 3: The pipeline of the classification with the LLM-based text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
-As discussed earlier, we chose all headline pairs that are significant at 0.05 level for our analysis.
-It is worth noting that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment.
+  • $\bullet$
+    Llama-3-8b: Meta’s Llama-3 (metallama3) is regarded as one of the most advanced open-source LLMs to date (llama3modelcard).
+We obtain embeddings from Llama-3 with 8 billion parameters (Llama-3-8b).
+We fine-tune this exact model using LoRA in §$\S$3.3; therefore, using the embeddings from this model allows for a direct comparison of fine-tuning vs.
+embedding-based classification methods (see §$\S$3.4 for comparisons).
+Llama-3-8b does not produce embeddings explicitly designed to represent the semantic meaning of entire input sequences.
+To address this issue, we apply the technique described in (behnamghader2024llm2vec), which transforms the Llama-3-8b model into a more suitable text embedding model, outputting embeddings with a dimension of 4,096.
+  • $\bullet$
+    Word2Vec embeddings: Word2Vec, proposed by mikolov_etal_2013a, was the leading embedding model before the widespread adoption of Transformers in the Natural Language Processing (NLP) literature.
+Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information.
+Embeddings from Word2Vec have been used in the recent marketing literature in a variety of use cases, including product attribute modeling (timoshenko_hauser_2019; wang_etal_2022), the effect of creativity on product popularity (sozuer_etal_2024).
+We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec.
+To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
+Next, to assess the impact of model complexity on performance, we consider two classification models:
+  • $\bullet$
+    Logit model with linear features: This is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables.
+  • $\bullet$
+    Multilayer perceptron: MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data.
+It employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting.
+We refer interested readers to Chapter 5 in (zhang2021dive) for more details.
+We train the MLP using the stochastic gradient descent algorithm to minimize binary cross-entropy and perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration.
+Please see Web Appendix §$\S$§C.1 for additional details of the fine-tuning process and the architectures used in the final implementation.
+Figure 3: The pipeline of the classification with the LLM text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
+Finally, in terms of the data, as discussed earlier, we choose all headline pairs that are significant at the 0.05 level.
+Note that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment.
 To avoid information leakage, we first randomly split the data into 80% training and 20% test sets based on the test ID (the experiment ID for Upworthy) to ensure all headline pairs under the same experiment are either in the training set or the test set.
 Notice that Upworthy allowed additional experiments for the same headline under a new experiment ID because the editor may require an extra test to make an informed decision.
-It means that we may have the same headlines in both the training and test sets, even if they were in totally different experiments.
-In such cases, we discard headlines that appeared in the training set from the test set to avoid information leakage.
+It means we may have the same headlines in both the training and test sets, even if they were in different experiments.
+In such cases, we discard headlines that have appeared in the training set from the test set to avoid information leakage.
 [^11] There are alternative ways to do the train-test split without discarding any samples.
-One approach is assigning one headline pair at one time to training or test groups and then adding all pairs with the same headline in the same group until all samples are allocated.
-However, in our study, after conducting several experiments, we found that the splitting procedure has minimal impact on the final results.
-Hence, to maintain clarity and simplicity in our exposition, we simply discard duplicated testing samples, which constitute around 25% of the original test set.
-One may still be concerned if this Upworthy data has been used as part of OpenAI’s training corpus, which could potentially invalidate our analysis.
-However, we argue this is unlikely for several reasons.
-First, it is improbable that OpenAI would have specifically downloaded this Upworthy csv format dataset from a subfolder in Upworthy Research Archive.
+One approach is assigning one headline pair at one time to a training or test set and then adding all pairs with the same headline in the same group until all samples are allocated.
+However, after conducting several experiments in our study, we found that the splitting procedure has minimal impact on the final results.
+Hence, to maintain clarity and simplicity in our exposition, we simply discard around 25% of test headlines that have appeared in the training set.
+3.2.2 Performance Analysis of Embedding-based Approaches
+In Table 6, we present the performance of various embedding-based classification models.
+Embedding Model Classification Model Accuracy
+OpenAI-256E Logit 77.31%
+MLP 77.62%
+OpenAI-3072E Logit 82.50%
+MLP 82.26%
+Llama-3-8b-4096E Logit 74.03%
+MLP 79.40%
+Word2Vec-256E Logit 67.89%
+MLP 67.32%
+Word2Vec-3072E Logit 67.55%
+MLP 67.02%
+Table 6: The performance of different embedding and classification models on test data.
+All models use the same training data, and accuracy is evaluated on the test data.
+Note that the Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents.
+Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average accuracy from three reruns.
+At a high level, we see that all the embedding-based classification models perform better than prompt-based approaches.
+This performance can be attributed to the fact that these models are trained on a customized dataset, allowing them to learn decision rules tailored to the headline selection task.
+Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust.
+In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations.
+Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus.
+Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
+Next, we compare performance under different embedding models.
+We see that OpenAI embeddings outperform both Llama-3-8b and Word2Vec-based models.
+This is understandable since OpenAI’s LLMs are state-of-the-art models that naturally perform better than the classic Word2Vec embeddings.
+Also, the performance of embeddings from open-source Llama-3-8b falls between that of the proprietary OpenAI model and the older Word2Vec that does not use the transformer architecture.
+We also see that for OpenAI, higher-dimensional embeddings give higher-quality encoding and, thus, higher prediction accuracy.
+However, there is some limit beyond which there are diminishing returns.
+In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 5%.
+A more striking observation is that classification model complexity does not improve the prediction for OpenAI embeddings.
+Note that there is no significant difference in accuracy between MLP and Logit for the following embedding models – OpenAI-256E, OpenAI-3072E, Word2Vec-256E, and Word2Vec-3072E.
+But for Llama-3-8b embeddings, the MLP classification model outperforms Logit classification model by approximately 5%.
+We hypothesize that with more dimensions in text embeddings, the separation between different classes becomes more complex (e.g., higher sparsity and higher risk of over-fitting), requiring a more sophisticated model for classification.
+Overall, our findings suggest that a Logit classifier with linear features on top of fixed text embeddings from OpenAI can effectively capture the attractiveness of content.
+We now investigate a few other aspects of these models and summarize our findings below.
+For brevity, we simply discuss the findings here and refer interested readers to the Web Appendix for details.
+Note that in all these analyses, we focus on OpenAI embeddings (since it offers the best performance) and the logistic regression with linear features as our classification model (since the MLP does not offer any significant improvement over the logit model for OpenAI embeddings and is easier to train).
+  • $\bullet$
+    First, we examine the conditions under which these approaches fail.
+We use cosine similarity measures to quantify the similarity between the two headlines in a pair and find that these approaches tend to perform worse when the headlines are similar.
+This is understandable since text embeddings may not be able to capture the differences between two headlines when they are similar.
+As a result, it is hard for a classification model to differentiate between them.
+See Web Appendix §$\S$§C.2 for details.
+  • $\bullet$
+    Second, we quantify the role of the number of training samples on predictive accuracy.
+For the 256E-dimensional embedding, we find that approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance.
+However, when we use the 3072E-dimensional embedding, there is evidence of overfitting even when we use 95% of the data for training.
+This suggests that for larger embedding vectors, we need much larger training datasets.
+See Web Appendix §$\S$§C.3 for details.
+  • $\bullet$
+    Third, we examine whether we can improve the performance and alleviate any overfitting issues by adopting dimensionality reduction techniques (PCA) and/or regularization (adding $L^{1}$/$L^{2}$ norms).
+However, we don’t see any significant improvement with these methods.
+See Web Appendix §$\S$§C.4 for details.
+  • $\bullet$
+    Finally, there are two potential threats to the inference from our analysis.
+First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate the results.
+A second concern is that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created.
+We examine both these issues and provide a detailed discussion and robustness checks to alleviate concerns around them in Web Appendix §$\S$§C.5.
+3.3 Fine-Tuning Open-Source LLMs with LoRA
+In §$\S$3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself.
+The best-performing embedding-based model reaches an accuracy of 82.50% without modifying any parameters in the LLM.
+This raises a natural question: if we can train LLMs directly, can we achieve better performance?
+However, training LLMs from scratch is costly and resource-intensive.
+Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLM to improve performance while keeping the computational cost manageable.
+Note that to fine-tune LLMs, we need to be able to access the underlying model parameters.
+This is only feasible for open-source LLMs, like BERT and Llama.
+[^12] Some proprietary LLMs offer black-box fine-tuning.
+For instance, OpenAI allows fine-tuning GPT-3.5, GPT-4/GPT-4o (with limited experimental access) through its API in a black-box manner.
+Users provide the training and test datasets and select the base GPT model, but are not provided any details of the fine-tuning process or the underlying model parameter.
+This approach also has a number of other drawbacks – the fine-tuning API is quite expensive compared to the open-source fine-tuning approaches, the fine-tuned model cannot be hosted on a local machine, data privacy concerns since the media firm’s proprietary data will need to be submitted to OpenAI.
+Given these drawbacks, it is preferable for firms to fine-tune open-source LLMs, as the process and data are fully under the control of the firm.
+However, for the completeness of comparison, we fine-tuned the ChatGPT-3.5 model on the training set using this API, with the default configuration provided by OpenAI, and the same training-test dataset split.
+We obtained an accuracy of only 79.80% on the test set.
+The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
+For the fine-tuning task, we use Meta’s Llama-3-8b as our primary base model.
+zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance.
+Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b).
+We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see §$\S$3.3.3 for details).
+Hence, we avoid fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated.
+The rest of this section is organized as follows.
+We present the LoRA method for efficient fine-tuning in §$\S$3.3.1.
+The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning.
+Then we detail our implementation in §$\S$3.3.2 and report our results in §$\S$3.3.3.
+Background information related to Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 are given in Web Appendix §$\S$§D.1.
+3.3.1 LoRA Fine-Tuning Technique
+Traditional fine-tuning of LLMs involves updating a substantial number of parameters, which can be computationally expensive and memory-intensive.
+To efficiently fine-tune the Llama-3 model on hardware with limited GPU memory, we employ Low-Rank Adaptation (LoRA), a popular Parameter-Efficient Fine-Tuning (PEFT) method developed by (hu2021lora), enabling practical fine-tuning of large-scale LLMs on medium-scale hardware.
+[^13] There are many alternatives for PEFT, like Prefix Tuning (li2021prefix), Adapter Layers (houlsby2019parameter), and BitFit (zaken2021bitfit).
+We select LoRA due to its excellent balance between simplicity and efficiency, making it the top choice for fine-tuning.
+LoRA operates under the assumption that updates during model adaptation exhibit an intrinsic low-rank property and introduce a set of low-rank trainable matrices into each layer of the Transformer model.
+This approach leverages the low-rank decomposition, which significantly reduces both memory usage and computational time through a significantly reduced number of trainable parameters.
+We illustrate LoRA in Figure 4.
+We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$.
+Specifically, $W_{Q}$ is the weight matrix for the query vectors, $W_{K}$ is for the key vectors, and $W_{V}$ is for the value vectors in the attention mechanism of the Transformer architecture.
+These matrices store most of the information and knowledge learned by LLM, and are crucial for calculating the attention scores and subsequently determining the relevance of different tokens in the input sequence.
+For more technical background, we refer to Web Appendix §$\S$§D.1.
+We use $d\times d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices.
+We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
+$W^{\prime}=W+\Delta W,\quad\mathrm{with}\quad\Delta W=AB,\quad\mathrm{where}\quad A\in\mathbb{R}^{d\times r},\quad B\in\mathbb{R}^{r\times d_{W}}.$
+Figure 4: Illustration of LoRA Fine-Tuning.
+LoRA essentially introduces the low-rank matrices $A$ and $B$ to approximate the weight matrix change $\Delta W$ after the fine-tuning.
+These low-rank matrices are updated during the fine-tuning process, while the original weight matrix $W$ remains frozen.
+This strategy significantly reduces the number of trainable parameters from $d\times d_{W}$ to $r\times(d+d_{W})$ because the rank $r$ is typically set to a small value, such as 4 or 8, to balance the trade-off between model capacity and computational efficiency.
+3.3.2 Implementation Details
+We use exactly the same training and test sets as in §$\S$3.2, but the inputs here are the original headlines with the text information.
+To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them.
+For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
+We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 5.
+Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence.
+Thus, the last token in the sequence, being informed by all prior tokens, is used for classification.
+This embedding vector of the last token (which is 4096-dimensional), output by the last Transformer layer, is processed through a linear layer producing a two-dimensional logit vector, denoted by $Z=(Z_{1},Z_{2})$.
+The $Z$ vector is then transformed into probabilities by the softmax function, indicating the likelihood of each headline being catchier, i.e.,
+$\mathbb{P}\{\text{i-th headline is catchier}\}=\frac{e^{Z_{i}}}{e^{Z_{1}}+e^{Z_{2}}}\mathrm{~{}for~{}}i=1,2.$
+Figure 5: The classification pipeline using the Llama model.
+Headlines are first tokenized and each token is converted into an embedding vector.
+For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration.
+The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens.
+Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity.
+The output from the last token is fed into a classification head, which computes the probability that one headline is more impactful than the other.
+During fine-tuning, our objective is to minimize the cross-entropy loss on the training set.
+We implement LoRA fine-tuning using the PEFT library developed by Hugging Face.
+We set the dimension of the low-rank matrices as $r=4$.
+Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora).
+Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$.
+To mitigate over-fitting issues, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices.
+[^14] Furthermore, we enable 16-bit mixed precision training.
+Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly.
+This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
+With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which translates to only 0.023% of the total 7,506,636,800 parameters.
+We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training.
+A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process.
+This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
+3.3.3 Performance Analysis of Fine-tuning-based Approaches
+Figure 6: Loss curve and accuracy curve as the number of training epochs increases.
+We repeat the fine-tuning process three times and display the average loss and accuracy.
+For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits.
+The final accuracies of the three replicates on the test set are 83.59%, 82.70%, and 83.93% respectively, and 83.40% on average.
+We add the best performance (red dash line) obtained using OpenAI’s embedding in §$\S$3.2 for comparison.
+We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 6.
+We observe an accuracy of 83.40% accuracy on the test data, which is marginally better than the best result (82.50% accuracy) using OpenAI embeddings in §$\S$3.2.
+However, such marginal improvements should not undermine the utility of fine-tuning.
+Comparing the fine-tuning of Llama-3-8b with the use of OpenAI embeddings is not straightforward due to inherent differences in models’ capability.
+To highlight the benefits of fine-tuning, we observe that the MLP classification model using embeddings derived from Llama-3-8b (79.40% accuracy in Table 6) performs worse than LoRA fine-tuned Llama-3-8b (83.40% accuracy).
+This demonstrates that fine-tuning Llama-3-8b can enhance performance compared to Llama-3-8b embeddings.
+Similar to the case of the embedding-based classification model, we notice that the fine-tuning approach also suffers from some overfitting issues; see Figure 6.
+This is because Llama-3-8b is inherently complex; even though LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex.
+Common approaches to address over-fitting include reducing model complexity, adding regularization, and increasing the dataset size.
+We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting.
+However, these modifications generally reduced performance on the training set without yielding significant improvements on the test set.
+We also consider two other versions of the LoRA fine-tuned model to document their performance:
+  • $\bullet$
+    Larger dataset: One way to improve performance and alleviate overfitting is to fine-tune the model on a larger dataset that also includes data on headline pairs that are not significantly different from each other.
+In our setting, including insignificant pairs of headlines expands the dataset from 39,158 pairs to 134,941 pairs.
+However, even with this expanded training data, accuracy on the test set continues to hover around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset.
+We refer interested readers to Web Appendix §$\S$§D.2.1 for details of this analysis and numerical results.
+  • $\bullet$
+    DistilBERT Fine-tuning: Before the widespread adoption of the term “LLM”, BERT was considered one of the best language models available, and it is still widely used as a benchmark in many NLP tasks.
+To provide a comparison with these earlier models, we therefore fine-tune DistilBERT – a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (sanh2019distilbert; devlin2018bert).
+Fine-tuned DistilBERT provides an accuracy of about 75% on the test dataset, which is good, but still lower than that of the fine-tuned Llama-3-8b model.
+Please see Web Appendix §$\S$§D.2.2 for details on the implementation and performance of the fine-tuned DistilBERT.
+In sum, we find that using more data and/or different base models does not lead to significant improvement in performance in our setting.
+3.4 Summary of Performance of Pure-LLM Based Methods
+We summarize the performance of various pure-LLM-based methods in Table 7.
+We also compare the accuracy of human respondents on the same task.
+Details of the survey questions and the accuracy analysis of human respondents are shown in Web Appendix §$\S$§A.
+Method Accuracy Notes
+Human respondents 51.58% On 4,571 human labeled pairs
+Prompt-based approach $<$65% GPT-4 with in-context learning, evaluated on 1,000 random samples
+Classification model with LLM text embeddings 82.50% OpenAI-3072E embedding model with logistic regression
+Fine-tuned LLM model 83.40% LoRA fine-tuned Llama-3-8b
+Table 7: Summarization of performance of various pure-LLM-based methods and comparison with survey results.
+All results in this table have been reported before.
+We only report the best performance of each method.
+There are four main takeaways from these comparisons.
+First, human respondents cannot easily predict which headline would be more engaging and likely to lead to higher CTRs.
+This is consistent with the fact that media firms (and even expert editors) rely on experimentation to decide which headlines and articles to display to users.
+Second, while simple prompt-based approaches like in-context learning with examples are marginally better than purely random guesses or human predictions, they nevertheless do not provide sufficiently high accuracy.
+Third, both embedding-based models and fine-tuning perform equally well (at least in our setting), and provide equally good performance with around 83% accuracy.
+They each have their pros and cons.
+The main advantage of the logit classification model with OpenAI embeddings is that it is extremely easy to implement and use in business settings since it does not require any significant model training/deployment.
+However, embeddings from the OpenAI models are proprietary, and as such, we lack visibility into the exact nature of the embeddings.
+In contrast, fine-tuning approaches use open-source models such as Llama-3-8b, and we have access to the underlying parameters of the LLM.
+However, this transparency comes with significant costs – fine-tuning LLMs can be computationally expensive and requires non-trivial training and deployment skills that not all businesses may have or want to invest in.
+The fourth and most important takeaway is that while pure-LLM-based approaches are informative, they are unable to reach the accuracy of experimental approaches.
+That is, firms cannot replace content experimentation with LLM-based predictions.
+Finally, we discuss the monetary costs associated with three LLM-based approaches.
+The cost of prompt-based, embedding-based, and LoRA fine-tuning in our analysis are around \$20, \$0.18, \$3, respectively.
+The cost of prompt-based and embedding-based comes from the access to OpenAI’s API in order to get responses and text embeddings.
+The cost of fine-tuning comes from renting a cloud computing server equipped with one NVIDIA A100 GPU.
+Overall, we find that the prompt-based approaches are the most expensive, while embedding and fine-tuning are the cheapest in terms of cost.
+However, in terms of run time and effort, fine-tuning is more costly.
+The detailed comparison of the cost, run-time, and engineering effort required for each pure-LLM method are discussed in detail in Web Appendix §$\S$§E.
+4 Our Approach: LLM-Assisted Online Learning Algorithm
+So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 83% for the embedding-based and LoRA fine-tuning approaches).
+The main advantage of pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it.
+Thus, there is no need to use any traffic for experimentation.
+The downside, of course, is that these models are not fully accurate – in a pairwise comparison, in around 17% of the cases, they are likely to mispredict which headline is better.
+This can, of course, lead to lower clicks and revenues for the firm.
+[^15] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth.
+However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples.
+Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
+On the other hand, we have the current status quo, which is the use of experimentation-based approaches to choose headlines/content.
+As mentioned in §$\S$1, there are two broad experimentation-based methods—(1) A/B tests and (2) Online learning algorithms or bandits.
+In particular, the standard practice at Upworthy during the time of the data collection was A/B tests, where the firm allocated a fixed amount to different headlines and then chose the best-performing one for the remaining traffic.
+This approach is also known as Explore and Commit (E&C) and is commonly used in the industry due to its simplicity and low engineering cost; see Chapter 6 of (lattimore2020bandit).
+The main benefit of this approach is that, with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased; as such, it is the gold-standard approach for identifying the best arm (or headline in this case).
+The downside, of course, is that by assigning traffic to all the headlines during the exploration phase or A/B test, the firm incurs high regret (i.e., low profits since some headlines in the test are likely to have low CTRs).
+Therefore, many modern technology and media firms such as NYTimes and Yahoo use online learning algorithms or bandits that leverage the explore-exploit paradigm to lower the experimentation cost while still learning (lattimore2020bandit).
+These algorithms move traffic away from poorly performing arms/headlines dynamically in each period, based on the observed performance of each arm till that period.
+Because of this greedy behavior, they tend to switch to the best-performing arms quickly and incur lower regret.
+Nevertheless, one drawback of these methods is that they start with the assumption that all headlines are equally good, i.e., they suffer from a cold-start problem, which can result in wasting traffic on the weaker headlines early on.
+In this section, we present a framework that combines the benefits of LLM-based models with experimentation based methods.
+Our key idea is that firms can address the cold-start problem in content experiments by leveraging predictions from LLM-based approaches as CTR priors in the first step and combining them with an online learning algorithm in the second step.
+Thus, our approach, termed LLM-Assisted Online Learning Algorithm (LOLA), continuously optimizes traffic allocation based on LLM predictions together with the data collected from the online experiment in real time.
+The rest of this section is organized as follows.
+In §$\S$4.1, we present the media firm’s problem and the standard bandit framework, and then in §$\S$4.2, we present our LOLA framework.
+Next, in §$\S$4.3, we discuss the implementation of our proposed algorithm, benchmarks, and the experimental setup.
+We present our numerical results in §$\S$4.4, and finally, in §$\S$4.5, we show how our LOLA approach can be easily generalized to a wide variety of settings and present a set of natural variants and extensions of the basic LOLA framework.
+4.1 Bandit Framework
+Multi-Armed-Bandits (MAB) is a class of sequential decision-making problems where a learner must choose between multiple arms over periods to maximize cumulative reward.
+While there is a large literature on it in the machine learning and statistics community, the literature in marketing is small, but growing; researchers have used it for optimizing website design, pricing, advertising, and to learn consumer preferences (hauser_etal_2009; schwartz2017customer; misra_etal_2019; liberali_ferecatu_2022; aramayo_etal_2023; jain_etal_2024).
+In our setting, different arms refer to different headlines associated with the same article, and periods refer to impressions where different headlines with the same article are displayed.
+Let $[K]:=\{1,2,\dots,K\}$ represent the set of $K$ arms, and $[T]:=\{1,2,\dots,T\}$ represent the entire decision horizon for a test.
+At each time step $t$, the learner selects an arm $a_{t}\in[K]$ and receives a reward $r_{a_{t}}$ drawn from a probability distribution specific to the chosen arm $a_{t}$.
+We assume each arm has a constant click-through rate (CTR), and the reward is the click feedback, which follows the Bernoulli distribution with a probability equal to the CTR.
+The learner’s goal is to minimize cumulative regret, which is defined as the difference between the reward obtained by always choosing the optimal arm and the reward actually accumulated by the learner under the policy played by the learner.
+Mathematically, the regret $R_{T}$ after $T$ trials is given by:
+$R_{T}=\max_{k\in[K]}\sum_{t=1}^{T}r_{a^{*}}-\sum_{t=1}^{T}r_{a_{t}},$ (1)
+where $a^{*}=\arg\max_{k\in[K]}\mathbb{E}[r_{k}]$ is the optimal arm with the highest expected reward.
+In MAB settings with finite arms, the standard algorithm that is typically used is the Upper Confidence Bound (UCB) algorithm (auer2002finite).
+UCB stands out among all online learning algorithms because it is provably asymptotically optimal.
+The UCB policy is straightforward – it selects the arm that maximizes the upper confidence bound on the estimated rewards.
+Specifically:
+$a_{t}=\arg\max_{k\in[K]}\left(\bar{\mu}^{t}_{k}+\sqrt{\frac{2\log(1/\delta)}{n^{t}_{k}}}\right),$ (2)
+where $\bar{\mu}^{t}_{k}$ is the estimated mean reward of arm $k$ up to time $t$, $n^{t}_{k}$ is the number of times arm $k$ has been played up to time $t$, and $\delta$ is the confidence level, which quantifies the degree of certainty.
+The intuition behind UCB’s effectiveness in balancing exploration and exploitation lies in its construction.
+By adding a confidence term $\sqrt{{2\log(1/\delta)}/{n^{t}_{k}}}$ to the estimated mean reward $\bar{\mu}_{k}$, UCB ensures that arms with fewer selections (higher uncertainty) are given a chance to be explored.
+As time progresses and more data is collected, the confidence term diminishes, allowing the algorithm to exploit arms with higher observed rewards.
+This dynamic adjustment enables the UCB algorithm to explore under-explored arms while exploiting arms with high rewards systematically, thus effectively addressing the exploration-exploitation trade-off inherent in the MAB problem.
+However, we need to set the key hyperparameter $\delta$ such that it is small to ensure optimality with high probability, but not so small that suboptimal arms are over-explored; see Chapter 7.1 (lattimore2020bandit) for a detailed discussion.
+Typically, we set the confidence term as $\alpha\sqrt{\log t/n^{t}_{k}}$ and fine-tune the scale term $\alpha$ for specific problems.
+4.2 LOLA
+We now discuss our approach LOLA, and how the standard UCB approach discussed above can be extended to include information from LLMs in the first step preceding active experimentation.
+LOLA has two key steps, and we describe each in detail.
+In the first step, we use LLMs to get an initial prediction of each headline’s CTR.
+Note the prediction task here is somewhat different from that in §$\S$3.
+In the previous section, we used LLM-based approaches to perform pairwise comparisons of headlines and predict which headline amongst a pair is more engaging.
+However, as discussed earlier, that approach does not work when we have more than two headlines (which is the case for over 90% of the A/B tests in our data).
+Further, we need actual CTR predictions for each headline to feed into the second step of our algorithm.
+Therefore, we adapt the LLM-based classification models from §$\S$3 for predicting CTRs by making the following modifications: changing the loss function from log loss to mean squared error (MSE), using single headline inputs instead of headline pairs, and switching from binary output to continuous CTR prediction.
+The training and fine-tuning procedures remain unchanged.
+[^16] Using CTR as the outcome measure does not adversely affect the performance.
+For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.31% accuracy obtained with logistic regression.
+This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
+We compare two approaches for CTR prediction: an OpenAI-embedding-based method and a LoRA fine-tuned Llama-3-8b model.
+Our numerical results indicate that the R-squared values for the embedding-based approach and the LoRA fine-tuned method are 0.241 and 0.315, respectively.
+When selecting the headlines with the highest predicted CTRs, the accuracy—defined as the percentage of times the chosen headline indeed has the highest actual CTR—is 42.66% for the embedding-based method and 46.55% for the LoRA fine-tuned method.
+Note that this accuracy is significantly lower than the binary classification accuracy discussed in §$\S$3 because we can have more than two headlines in a test and many headlines have no significant difference in CTRs.
+Based on these results, we select the LoRA fine-tuned Llama-3-8b model as the best CTR prediction model in the first step.
+Additional details on the CTR prediction methods and their results can be found in Web Appendix §$\S$§F.
+In the second step, we build on the bandit framework from §$\S$4.1 to design an online algorithm that incorporates the LLM-based CTR predictions from the first step.
+Inspired by the 2-Upper Confidence Bounds (2-UCBs) policy designed by (gur2022adaptive), we propose Algorithm 1, titled LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs).
+Essentially, in this algorithm, we treat the LLM’s CTR predictions as auxiliary information prior to the start of online experiments, equating this auxiliary CTR information to additional impressions and click outcomes for each arm.
+Algorithm 1 LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs)
+1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
+2:Hyperparameter Fine-Tuning: Use another subset of data to select two hyperparameters for the algorithm, $\alpha\in\mathbb{R}^{+}$ for controlling the upper bound, $n^{\text{aux}}$ as the LLM’s equivalent auxiliary sample size.
+3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$, LLM-based CTR prediciton $\bar{\mu}^{\text{aux}}_{k}\leftarrow\mathcal{M}(x_{k})$, regular CTR initialization $\bar{\mu}^{1}_{k}=0$, accumulated impressions $n^{1}_{k}\leftarrow 1$, and accumulated clicks $c^{1}_{k}\leftarrow 0$, for all arms $k\in[K]$.
+4:for $t=1$ to $T$ do
+5:     Calculate the first UCB for all arms $k\in[K]$: $U_{k}^{1}=\bar{\mu}^{t}_{k}+\alpha\sqrt{\frac{\log t}{n^{t}_{k}}}$.
+6:     Calculate the second UCB for all arms $k\in[K]$: $U_{k}^{2}=\frac{c^{t}_{k}+\bar{\mu}^{\text{aux}}_{k}n^{\text{aux}}}{n^{t}_{k}+n^{\text{aux}}}+\alpha\sqrt{\frac{\log t}{n^{t}_{k}+n^{\text{aux}}}}.$
+7:     Play the arm $a_{t}=\arg\max_{k\in[K]}\min\{U_{k}^{1},U_{k}^{2}\}$.
+8:     Observe the payoff $r_{a_{t}}$
+9:     Update $n^{t+1}_{a_{t}}\leftarrow n^{t}_{a_{t}}+1$, $c^{t+1}_{a_{t}}\leftarrow c^{t}_{a_{t}}+r_{a_{t}}$, and $\bar{\mu}^{t+1}_{a_{t}}\leftarrow\frac{c^{t+1}_{a_{t}}}{n_{a_{t}}^{t+1}}$.
+10:end for
+There are two hyperparameters in LLM-2UCBs – (1) The upper bound control hyperparameter, $\alpha$, that also shows up in the regular UCB algorithm, and (2) The hyperparameter $n^{\text{aux}}$, which is specific to our proposed algorithm and indicates the equivalent auxiliary samples for initialization.
+At a high level, we can regard the LLM predictions as prior information or the information flow before the algorithm starts.
+This prediction is approximately equivalent to initially generating $n^{\text{aux}}$ impressions for each arm and observing their outcomes to obtain the sample average CTR, $\bar{\mu}^{\text{aux}}_{k}$.
+The hyperparameter fine-tuning process is simple.
+We can utilize another randomly sampled dataset to test combinations of two hyperparameters and select the combination that yields the highest accumulated rewards.
+During the online learning phase, the key step is balancing the exploration and exploitation trade-off.
+We capture this trade-off using the 2-UCBs rule.
+Essentially, we have two valid UCBs: $U_{k}^{1}$ is the regular UCB, which uses observed impressions and clicks to construct the mean reward estimator and adds a regular upper bound.
+The second UCB, $U_{k}^{2}$, incorporates LLM predictions.
+The calculation of $U_{k}^{2}$ is straightforward.
+The mean reward component is simply the sample average, taking into account the “auxiliary” samples generated by the LLM before the learning phase begins.
+These auxiliary samples are imaginary and approximate the richness of information contained in the LLM prediction.
+The next step is to choose the smaller of the two UCBs, $\min\{U_{k}^{1},U_{k}^{2}\}$, as the final UCB for selecting the arm.
+Note that if $n^{\text{aux}}=0$, the 2-UCBs rule is the same as the standard UCB rule.
+Conversely, when $n^{\text{aux}}\rightarrow\infty$, the 2-UCBs rule is equivalent to a pure exploitation policy (a.k.a.
+greedy policy) using the LLM-based CTR prediction.
+Thus, when $n^{\text{aux}}$ is a positive integer, the 2-UCBs algorithm incorporates information from both LLMs and the actual click outcomes.
+After playing the arm and observing the outcome, we update the impressions, clicks, and mean reward accordingly, following standard bandit algorithm procedures.
+The 2-UCBs algorithm offers three advantages.
+First, it is easy to implement, requiring the tuning of only two hyperparameters.
+Second, it can be used in conjunction with any LLM-based CTR prediction model in the first step and any outcome of interest in the second step.
+We use clicks as the outcome of interest given our application setting and data; however, other outcomes, such as time spent on a page or other measures of engagement, can be easily used instead of clicks.
+Finally, it is intuitive and easy to implement.
+Because it can easily accommodate the edge cases, e.g., one where the manager has no auxiliary information or one where the manager wants to fully rely on the LLM predictions, there is no need for extra engineering effort to set up separate systems for different edge cases.
+This adaptability is extremely valuable given the rapid advances in LLM technology.
+When the capability of the pure-LLM method continues to advance, the manager can easily fine-tune and increase $n^{\text{aux}}$ to take more advantage of LLM.
+4.3 Benchmarks and Implementation Details
+We now discuss the implementation details and present a set of benchmark algorithms.
+First, we randomly split our data into three exclusive folds based on the test ID, ensuring that all headlines in the same test are in the same fold and that the same headline does not appear in different folds.
+We use 70% of the data for training, 10% for hyperparameter fine-tuning, and 20% for testing.
+The training data is used to train the LLM models and is only used by our proposed LOLA approach (the LLM-2UCBs specifically) and pure LLM-based benchmarks.
+All hyperparameters in all the algorithms (both LOLA and the benchmark algorithms) are selected using the fine-tuning data, and the performance of the algorithms is evaluated on the test data.
+We now describe a set of commonly used algorithms that serve as benchmarks and our LOLA approach.
+  • $\bullet$
+    Explore-then-Commit (E&C): In this approach, the firm runs the A/B test for a fixed number of periods (i.e., explores) and then commits to the best-performing arm at that stage; see Chapter 6 (lattimore2020bandit).
+This is the standard A/B test approach used by many digital firms in the industry.
+In particular, Upworthy’s approach during the data collection period closely resembles the E&C algorithm – recall that they assign traffic to candidate headlines with equal probability during the experiment phase and then show the winning headline to all future impressions.
+The E&C method has a single hyper-parameter that needs tuning: the proportion of periods for exploration.
+We consider the following proportions – $\{0.1,0.2,0.3,0.4,0.5\}$, and choose the one with the highest accumulated rewards based on the 10% of data used for hyper-parameter tuning.
+We finally choose 0.2 for exploration.
+  • $\bullet$
+    Pure LLM-based approach: In this approach, the firm uses LLMs to predict the CTR for all the arms (headlines) and then allocates all the traffic to the arm with the highest predicted CTR.
+This is equivalent to a fully greedy policy based on LLM predictions.
+We use the training data to learn a model that predicts the CTR of a headline based on an LLM (either through text embeddings or LoRA fine-tuning).
+In our numerical study, we use the best-performing pure-LLM-based approach based on our experiments (see Web Appendix §$\S$§F), i.e., the LoRA fine-tuned Llama-3-8b with MSE loss, which gives a 46.55% accuracy.
+[^17] Note that since this is a pure prediction-based greedy approach, there are no hyper-parameters associated with the algorithm, and hence we do not use the fine-tuning data.
+  • $\bullet$
+    Standard UCB (UCB): This is the standard UCB algorithm discussed in §$\S$4.1.
+Here, the firm does not use any first-stage LLM predictions as an input, and there is only one hyperparameter to fine-tune, the confidence control, $\alpha$.
+We consider values from the set $\{0.02,0.04,\dots,0.10\}$, and find that $\alpha=0.08$ gives the highest accumulated rewards.
+  • $\bullet$
+    LOLA: This is our proposed approach.
+In the first step, we use CTR predictions from the best-performing pure-LLM-based CTR prediction model, i.e., the LoRA fine-tuning Llama-3-8b with MSE loss.
+Note that this is the same model that we use for the pure LLM-based approach, which allows for a fair comparison of the two approaches.
+For the second step, we use the LLM-2UCBs algorithm as outlined in §$\S$4.2.
+We fine-tune the combination of the two hyperparameters on the fine-tuning data, and we consider $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$.
+Based on fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ for our setting.
+    Note that the former two methods can be viewed as special cases of our proposed LOLA algorithm.
+Both our algorithm and the pure LLM-based approach use the exact same CTR prediction algorithm.
+Hence, the pure LLM-based approach (a.k.a pure exploitation, greedy algorithm) is a special case of LOLA, wherein we set $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results.
+This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon.
+Similarly, the standard UCB algorithm is a special case of LOLA with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions.
+Again, note that we use the same hyper-parameter $\alpha=0.08$ across both algorithms, which ensures that our algorithm defaults to the standard UCB with $\alpha=0.08$ when we ignore the CTR predictions from the LLM.
+After the fine-tuning of algorithms, we evaluate all algorithms on the test dataset, which has never been used for either CTR training or algorithm tuning.
+Therefore, none of the algorithms know the true CTRs of the headlines in the test data in advance.
+We use the average CTR for each headline in the test data as the true CTRs to generate the click feedback in our numerical simulations.
+This approach of generating click outcomes is reasonable in our setting since each headline was tested on a large amount of traffic by Upworthy, resulting in relatively narrow confidence intervals.
+UCB and LOLA are both asymptotically optimal in minimizing regret, which means there is no difference between these two algorithms when the number of time horizons $T$ goes to infinity, and the LLM-based prior information is not helpful in an infinite horizon case.
+However, in practice, the time horizon is never infinite.
+Therefore, we compare all the algorithms for different values of time horizons.
+One challenge in our setting is that the number of headlines varies across tests.
+Intuitively, for any given time horizon, it is always easier to minimize regret (or identify the best headline) when the A/B test has only two headlines compared to a case when there are three headlines; and, of course, the problem only gets more challenging as the number of headlines increases.
+As a result, comparisons across time horizons are meaningful only when we keep the number of headlines constant.
+Since the number of headlines varies in our dataset (see Table 3), we scale the time horizon for each A/B test using the metric “Traffic/Impressions per headline” (which we represent using the notion $\tau$), where we choose $\tau$ from $\{50,100,200,400,600,800,1000\}$.
+For example, if there are three headlines in an A/B test and the multiplier is $\tau=100$, then we set $T=100\times 3=300$ for this test.
+Thus, the effective horizon across all tests for a given $\tau$ is the same.
+In our experiments, it will be particularly important to investigate the performance of different algorithms at small-medium multipliers because these represent situations where the media firm only has a small amount of traffic to work with and needs to adaptively experiment and effectively allocate impressions to the right headline with limited experimentation ability.
+4.4 Numerical Results
+Figure 7: Average clicks per experiment per period under different time horizon multipliers.
+Note that the Y-axis captures the average clicks per test per period.
+For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
+The Y value is simply the average of this number $0.03$ over all tests.
+This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
+We visualize the results from our numerical simulations in Figure 7 and document the pairwise comparison of different algorithms, along with the relative percentage reward improvement and the p-values from the t-tests in Table 8.
+We find that, on average, LOLA performs the best among all algorithms across all time horizons.
+We document a few additional patterns of interest.
+Parameter $\tau$ LOLA vs UCB LOLA vs Pure LLM LOLA vs E&C UCB vs Pure LLM UCB vs E&C Pure LLM vs E&C
+50 $6.95$∗∗∗∗ $0.09$ $8.69$∗∗∗∗ $-6.41$∗∗∗∗ $1.62$ $8.59$∗∗∗∗
+100 $6.38$∗∗∗∗ $1.03$∗∗ $7.72$∗∗∗∗ $-5.02$∗∗∗∗ $1.26$ $6.61$∗∗∗∗
+200 $4.98$∗∗∗∗ $1.78$∗∗∗∗ $8.16$∗∗∗∗ $-3.05$∗∗∗∗ $3.03$∗∗ $6.27$∗∗∗∗
+400 $3.96$∗∗∗∗ $3.02$∗∗∗∗ $5.76$∗∗∗∗ $-0.90$ $1.74$ $2.66$∗∗∗
+600 $3.04$∗∗∗∗ $3.49$∗∗∗∗ $4.83$∗∗∗∗ $0.43$ $1.73$∗∗ $1.30$
+800 $2.35$∗∗∗∗ $4.20$∗∗∗∗ $5.36$∗∗∗∗ $1.81$∗∗∗ $2.94$∗∗∗∗ $1.12$
+1000 $2.23$∗∗∗∗ $5.02$∗∗∗∗ $5.08$∗∗∗∗ $2.73$∗∗∗∗ $2.79$∗∗∗∗ $0.05$
+Table 8: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001.
+For example, under the impression per headline equal to 50, the average click per test of LOLA and UCB is 0.053760 and 0.050267, respectively; so the relative improvement is calculated as ${0.053760}/{0.050267}-1=6.95\%$.
+Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
+First, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm.
+The small fluctuation at $\tau\in\{50,100,200\}$ is due to the randomness of average clicks per period, which tends to be larger under the smaller time horizon.
+For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends.
+This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance.
+Comparing the performance of LOLA with pure LLM methods, we find that LOLA generally outperforms Pure LLM, except when $\tau=50$, where there is no significant difference between LOLA and Pure LLM.
+This is because the short planning horizon does not allow LOLA to learn new information over and above the information already available from the LLM.
+As $\tau$ increases to 1000, we see that LOLA significantly outperforms Pure LLM, generating 5.02% more clicks.
+This suggests that if the horizon is very short, the planner/firm will not see significant improvement over simply relying on LLM predictions.
+However, as the horizon increases, LOLA will start showing significant gains.
+Next, comparing LOLA with regular UCB, we see that it always outperforms UCB.
+Interestingly, here we see that, as the time horizon increases, the relative improvement from LOLA diminishes.
+Intuitively, this is because both UCB and LOLA are active learning algorithms, and as the time horizon increases, they are both able to learn the CTRs better.
+While LOLA starts with a relative information advantage, as the time horizon increases, this advantage diminishes.
+Formally, this pattern is due to the asymptotic optimality of the UCB and LOLA algorithms, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period.
+We also see that E&C consistently underperforms compared to LOLA because it neither leverages LLM predictions nor achieves asymptotic optimality.
+While E&C shows better performance as the horizon increases, even when $\tau=1000$, LOLA is still better than E&C by 5.08%, which is a large improvement in the media and publishing industry.
+Finally, comparing the benchmark algorithms with each other, we see that when the horizon is short, the Pure LLM does better, while UCB is better when the horizon is longer.
+In general, this reflects the intuition that active learning is not very useful when the time horizons are very short since the algorithm has no time to explore and learn the CTRs effectively (and then exploit this learning).
+So simply going with the pure LLM predictions is better.
+In contrast, when the horizons are longer, UCB (or active learning) outperforms using static predictions from the LLM.
+Note that LOLA combines the strengths of both UCB and Pure LLM, and hence it is able to outperform both and provide superior performance in both short and long horizons.
+We also note that E&C is generally worse than both Pure LLM and UCB, which suggests that the current Upworthy practice is suboptimal.
+This is because E&C neither exploits LLM information and nor does it actively learn over time.
+In summary, we find that LOLA outperforms existing approaches for experimentation by leveraging the strengths of LLMs and combining them with the benefits of active/online learning.
+In particular, our results demonstrate the value of LOLA as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.
+4.5 LOLA Variants and Extensions
+So far, we have considered a version of LOLA that uses a fine-tuned CTR prediction model for LLM predictions in the first stage and an LLM-2UCBs algorithm that minimizes regret for online learning in the second step.
+However, LOLA is a general framework and can be easily adapted to a wide variety of settings.
+We highlight a few natural extensions and variants below.
+Details of the models and numerical results (when applicable) are shown in Web Appendix §$\S$§G.
+Best Arm Identification: So far, we focused on the goal of regret minimization, where the goal is to maximize the clicks/reward (per Equation (1)).
+While regret minimization is the most commonly studied goal in active learning and the natural goal for businesses in most settings, another commonly studied goal is Best Arm Identification (BAI).
+The goal of BAI problems is to identify the arm with the highest reward as fast and accurately as possible, regardless of accumulated regret.
+Formally, given a low failure rate $\delta$, we aim to find the arm $a^{*}$ with the highest expected reward, at least $1-\delta$ probability, while minimizing the number of total pulls.
+[^18] This is also called fixed-confidence setting (garivier2016optimal).
+Another common approach is the fixed-budget setting, where the goal is to maximize the probability of finding the best arm within a fixed number of pulls.
+We can easily modify the LOLA framework for a BAI goal.
+In Web Appendix §$\S$§G.1, we present a LLM-BAI algorithm that builds on the recently proposed BAI algorithm by mason2020finding.
+This algorithm is particularly effective in our setting since it can accommodate scenarios where the difference between the rewards of arms is small or insignificant.
+We show that our LLM-BAI algorithm significantly outperforms both the plain BAI algorithm (proposed by mason2020finding) as well as the standard A/B test under the same failure rate restriction.
+Please see Web Appendix §$\S$§G.1 for a detailed description of the LLM-BAI algorithm and its performance against the standard benchmarks.
+Thompson Sampling: Our bandit specification in §$\S$4.1 reflects a stochastic bandit setting rather than a Bayesian bandit setting.
+However, our proposed algorithms can easily be applied in a Bayesian setting.
+We present a Thompson sampling version of our proposed LOLA algorithm (LLM-TS) and its numerical performance in Web Appendix §$\S$§G.2.
+We see that UCB-based algorithms perform better than Thompson Sampling-based algorithms in both the LLM-assisted version and the standard version.
+That is, LLM-TS performs worse than LLM-2UCBs, and standard TS performs worse than UCB.
+This pattern has been observed in the literature, and it is generally known that TS suffers from over-exploration (min2020policy); in other words, UCB’s upper confidence shrinks faster, leading it to exploit more effectively.
+Nevertheless, in business settings where the firm already has a TS-based adaptive experimentation framework, it is easy to adapt the LOLA approach within this approach.
+User Information: So far, we have not considered user-level features or how to personalize the content shown to different users based on their behavioral/context features (mostly because the UpWorthy data does not have user-level features).
+Prior research has shown that using such features can significantly improve the match between users and content (li2010contextual; yoganarasimhan_2020; rafieian_yoganarasimhan_2021).
+It is, however, easy to expand our LOLA approach to include user features in both the LLM training phase as well as the online learning phase.
+We present two versions of LOLA that extend existing contextual bandit algorithms to our setting.
+First, we can extend the standard contextual linear bandits (chu2011contextual) to our LOLA framework by incorporating both user and text features in the LLM training and online learning phases of LOLA.
+However, this linear bandit approach has limitations since it assumes that the rewards model is linear in text and user features.
+Therefore, we also propose another solution based on the FALCON algorithm that does not make any linearity assumption (simchi2022bypassing).
+We present the details of both these contextual LOLA algorithms and their pros and cons in Web Appendix §$\S$§G.3.
+Alternative rewards: In the Upworthy setting, the measure of reward is clicks.
+However, click-based metrics have raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity.
+As a result, the news and media industry is now evolving to focus more on content quality and humanity, although clicks still remain an important measure; see (Upworthy2023) for a more detailed discussion.
+The LOLA framework can be easily adapted to alternative reward measures that align with new business goals, e.g., time spent on the platform during a session, or time spent on the article (and not just clicks on the headline).
+Apart from these extensions, the LOLA approach is quite general and adaptable on other dimensions as well.
+For instance, it is agnostic to the exact nature of the LLM-based approach used in the first step.
+Any of the LLM-based approaches discussed in §$\S$3 can be used in the first step.
+This is important since the LLM models and fine-tuning approaches continue to advance quickly.
+Further, LOLA can also be used in conjunction with content/creatives created by LLMs themselves.
+For instance, one simple way to leverage fine-tuned models is to use them to generate content (instead of using human-generated headlines/content).
+Indeed, recent studies have shown that promotional content (ads/emails) generated by fine-tuned LLMs tend to outperform traditional personalized ads and human-generated content in effectiveness (kumar2023generative; angelopoulos2024value).
+LLM-generated headlines/content can be used as competing arms/treatments within our LOLA framework.
+Additionally, the LOLA framework can also be easily extended to content recommendation systems used by social media platforms such as X and Facebook.
+In summary, the plug-and-play nature of our approach can easily accommodate newer advances and developments at all stages of the implementation process.
+5 Conclusion
+In conclusion, in this paper, we examine if and how firms can leverage the LLMs to enhance content experimentation in digital platforms.
+First, we examine how well LLMs can predict which content would be more appealing to users.
+We find that while LLMs provide informative predictions, even the best performing fine-tuned LLMs are unable to perfectly predict which content will be more appealing to users, and match the accuracy/regret of experimentation-based approaches.
+As such, firms cannot fully trust the prediction results solely from LLMs and replace experimentation.
+We introduce LOLA, a novel experimentation framework that combines the predictive power of LLMs with the adaptive efficiency of online learning algorithms to enhance content experimentation.
+By leveraging LLM-based prediction models as priors, our approach minimizes regret in real time, leading to significant improvements in total reward.
+Comprehensive numerical experiments based on a large-scale A/B testing dataset demonstrate that LOLA outperforms the traditional A/B tests and pure online learning algorithms.
+From a managerial perspective, LOLA offers a versatile solution to a broad set of scenarios where the firm needs to decide which content to display to users.
+LOLA is particularly valuable under limited traffic conditions that are common in many digital media platforms.
+This includes cases where content becomes stale quickly (e.g., news articles) or social media platforms like X and TikTok, where users generate a lot of new content regularly, and as a result, the amount of content is high relative to the amount of traffic.
+While our focal application is in the context of headline selection in the news and media industry, the framework can easily be applied to other problems, such as digital advertising, email marketing, and promotions.
+Indeed, a large literature in marketing focuses on how to use experiments in conjunction with machine learning methods to identify optimal treatments and/or personalize treatments; see rafieian_yoganarasimhan_2023 for a detailed review.
+The LOLA approach can be easily used in many of those settings and help with maximizing outcomes of interest while simultaneously reducing the cost of experimentation.
+Finally, from an implementation perspective, LOLA offers many advantages.
+Since it can utilize open-source LLM models, it is cost-effective and can be easily deployed across different tasks once LLM models are fine-tuned on relevant datasets.
+The fine-tuning process itself is cost-efficient and can be performed on entry-level GPUs, making advanced LLM-based techniques accessible to firms with limited budgets.
+Relying on open-source LLMs also ensures that the firm’s data stays within the firm and is not accessible to the owners of proprietary LLM models (such as Google and OpenAI).
+Further, our approach is compatible with many bandit-based experimental systems used by large digital firms such as Amazon (fiez_etal_2024), and as such, is easy to integrate and deploy.
+Code and Data
+The code used in this study is publicly available at LLM News Github Repository to enable the reproduction of our results and to support future research that applies LLMs to experimentation and business problems.
+Funding and Competing Interests Declaration
+Author(s) have no competing interests to declare.
+Appendix A User Survey
+We conduct a survey of undergraduate students from a large public university.
+The goal of the survey was to examine whether lay human users can accurately identify which headlines would lead to more clicks.
+Recall that the headlines themselves were created by Upworthy editors, but those editors need an experiment to identify the most appealing headline.
+We design two similar surveys using Qualtrics.
+The first survey randomly samples 3,000 pairs of headlines from 39,158 headline-pairs that are significantly different from each other (at a 0.05 significance level), along with 500 pairs that have no significant differences in CTRs.
+The second survey utilizes the same 3,000 significant samples as the first but does not include the insignificant samples.
+In both surveys, students were asked to select the catchier headline from a pair of headlines. The introductory text for the surveys was:
+“Please help us understand what makes a news headline appealing to readers.
+In this survey, you will be presented with 20 pairs of headlines.
+Please choose the one that captures your attention more effectively or feels more engaging in each pair, assuming both headlines relate to the same news content.
+There are no right or wrong answers; we are simply interested in your opinions.”
+Respondents then proceed to a new page featuring 20 headline comparison questions with the prompt, “Please select the headline that you find more attractive,” with only one permissible response.
+In all the comparisons, we randomly swap the order of the headlines to ensure that the display order does not correlate with the correct answer.
+We also incorporated two attention checks in the second survey.
+The first survey included an additional response option: “Both headlines seem equally good.” At the survey’s conclusion, we also collected data on users’ news consumption habits.
+Specifically, we ask, “How often do you read the news?” with options “Never”, “1-2 days per week”, “3-4 days per week”, “5-6 days per week”, “Everyday”.
+Detailed information on the surveys and the results from both surveys are shown in Table A1.
+Note that the total responses received is less than the number of respondents ×$\times 20$ because some students did not finish the survey or skipped some comparisons.
+For example, for the first survey, the total number of responses 3,809 is less than $202\times 20=4,040$.
+To calculate the accuracy metric, we only focus on the significant pairs, and in the first survey, we discard the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys.
+“Accuracy” measure here is consistent with the LLM tests.
+We treat the headline with a higher CTR as the correct answer and calculate the proportion of correct answers as accuracy.
+By doing so, one should expect 50% accuracy by random guess since
+Detailed information and survey results are documented in Table A1.
+Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$).
+To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys.
+Accuracy here aligns with the LLM tests.
+We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy.
+As such, we should expect a 50% accuracy rate from random guessing.
+The survey results indicate that respondents’ ability to identify catchier headlines is comparable to random guessing – the accuracy rate in both surveys is not significantly different from 50% at the 0.05 significance level.
+Furthermore, accuracies did not vary significantly across different groups based on news reading frequency.
+These findings suggest that the human ability to discern catchier headlines is limited, highlighting the intrinsic difficulty of the task.
+Survey 1 Survey 2
+conducted time 4/14 - 4/21 4/29 - 5/17
+# of respondents 202 339
+headline pair pool 3000 significant + 500 insignificant pairs 3000 significant pairs
+# of sampled pairs in survey 20 20
+# total received responses 3,809 4,571
+overall accuracy % 45.22 51.58
+Results based on different answer to “How often do you read the news?”
+Never proportion % 21.78 16.22
+accuracy % 45.76 51.62
+1-2 days per week proportion % 48.51 49.56
+accuracy % 45.36 52.54
+3-4 days per week proportion % 18.81 18.29
+accuracy % 46.83 51.39
+5-6 days per week proportion % 4.95 5.31
+accuracy % 39.05 50.15
+Everyday proportion % 5.94 8.26
+accuracy % 42.35 47.48
+Table A1: Survey Response Analysis.
+No accuracies in this table are significantly lower or higher than 50% at 0.05 significant level.
+8 respondents did not finish Survey 2 in the limited given time, so the summation of proportion from different subgroups is smaller than 100% in Survey 2.
+Appendix B Appendix for Prompt-based Approaches
+B.1 Pairwise Comparison Between GPT Prompts
+To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on 1,000 correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other.
+The mean differences and p values are documented in Figure A1 and Figure A2.
+The random guess model simply predicts the first headline wins with a half probability.
+The nomenclature for GPT prompts in these two figures follows a specific format.
+The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are flipped.
+A value of 0 indicates the correct label, while a value of 1 indicates the wrong label.
+For example, GPT-4-0-0 means zero-shot prompting using GPT-4, and GPT-3.5-5-1 means in-context learning with 5 demonstrations and wrong labels using GPT-3.5.
+Figure A1: Average accuracy differences between different GPT prompts.
+The number in the heatmap presents the accuracy of prompts indicated on the left minus the accuracy of prompts indicated at the bottom.
+Shaded blocks indicate insignificant differences.
+Figure A2: The number in the heatmap presents the p-value from the paired t-test between different GPT prompts.
+Appendix C Appendix to Embedding-based Classification Approaches
+C.1 Implementation Details of MLP Classification Models
+Hyperparameters of the network structure.
+We employ a fully connected neural network with ReLU activation functions for the hidden layers and a sigmoid activation function for the output layer.
+This configuration is widely used in practice.
+To determine the number of layers and the hidden nodes of each layer, we choose the number of layers within the range of $\{1,2,3,4,5\}$, and set the number of nodes in the first fully connected layer from the set $\{128,256,512,1024,2048\}$.
+Besides, we adjust the previous-post ratio, which is the proportion of nodes in each layer relative to those in the subsequent layer.
+This ratio here is in $\{2,4,8\}$.
+Note that the previous-post ratio becomes irrelevant to the results when there’s only one hidden layer.
+This is because the final output layer has only one dimension for ranking the most attractive headline.
+We incorporate the dropout after each hidden layer to mitigate the over-fitting.
+Dropout layers randomly deactivate neurons based on a specified dropout rate, while scaling up the remaining neurons to maintain the same mean of the layer.
+In this work, we set the dropout rate in $\{0.3,0.5\}$.
+All hyperparameters are automatically selected through hyperparameter optimization, which is introduced below.
+Hyperparameters of the training procedure.
+The MLP is implemented in PyTorch, with data split into 70% for training, 10% for validation, and 20% for testing.
+To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch.
+Determining the appropriate batch size is crucial, as large batch sizes may lead to convergence to local optima, while small batch sizes may hinder convergence.
+We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$.
+Additionally, to mitigate over-fitting, we apply $L^{2}$ regularization, with the regularization weight chosen from $\{1e-6,1e-4,1e-2\}$.
+We use binary cross-entropy as the loss function and employ stochastic gradient descent for training.
+We implement early stopping by monitoring the validation set loss, halting the training process if the reduction in loss over 60 epochs falls below a tolerance threshold of $10^{-4}$.
+Hyperparameter optimization.
+We optimize the parameters mentioned above through Bayesian Optimization (snoek2012practical).
+This Bayesian technique uses Bayesian inference to construct a probabilistic model of the objective function, where the parameters are updated iteratively with new observations.
+This iterative process balances exploration of the parameter space to discover promising regions and exploitation of known high-performing areas.
+In this work, we consider the accuracy of the validation set as the objective and optimize hyperparameters selected from the specified parameter space, as delineated by the ranges mentioned previously.
+We choose the Gaussian process as the estimator with the automatic optimizer, which is commonly adopted by the community.
+To ensure a judicious trade-off between efficacy and efficiency, we repeat the optimization process 50 times for input data of 256 dimensions and 30 times for data of 3,072 dimensions.
+C.2 Effect of Headline Similarity on Performance
+We now examine the conditions under which these approaches fail.
+, Specifically, we investigate whether headline similarity affects the performance of embedding-based classification models.
+We first calculate the cosine similarity based on the text embeddings for each pair of headlines.
+This metric captures the similarity between two vectors in inner product space and is defined as the dot product of the vectors divided by the product of their lengths, i.e., ${e_{1}\cdot e_{2}}/(\|e_{1}\|\|e_{2}\|)$, where $e_{1}\cdot e_{2}$ is the dot product of the embedding vectors, and $\|e_{1}\|$ and $\|e_{2}\|$ are the their Euclidean norms.
+Figure 3(a) shows the distribution of cosine similarity across all headline pairs, whereas Figure 3(b) shows the distributions of cosine similarity separately for two cases: (1) cases where our embedding-based classification correctly predicts which headline is catchier, and (2) cases where our embedding-based classification incorrectly predicts which headline is catchier.
+We observe a rightward shift in the distribution of incorrect predictions (compared to the distribution of correct predictions).
+This suggests that incorrect predictions are more common when the headlines are very similar to each other (i.e., higher cosine similarity).
+There are two possible reasons for this.
+First, in cases where the headline pairs are very similar, there is a higher chance of Type 1 error (in the t-tests based on our A/B test data), i.e., the headlines may not be significantly different in fact.
+Second, text embeddings may not be able to capture the differences between two headlines when they are similar, i.e., the embeddings may also be very similar.
+As a result, it is hard for the classification model to differentiate between them.
+[^19] Note that it is unlikely that this phenomenon is due to problems/issues with the classification model because both Logit and MLP achieve nearly identical performance.
+(a) Cosine similarity distribution over all pairs in training and test sets
+(b) Cosine similarity over incorrect/correct predicted test pairs
+Figure A3: The left panel shows the distribution of cosine similarity for all pairs.
+The right panel shows the distributions of cosine similarity for both incorrect and correct predictions using the logit classification model with OpenAI-3072E embeddings.
+C.3 Effect of Training and Test Data Size on Performance
+(a) Performance of OpenAI-256E
+(b) Performance of OpenAI-3072E
+Figure A4: Performance of Logit classification models with OpenAI embeddings under varying training dataset sizes and different embedding sizes.
+We now conduct experiments varying the proportion of data used for training vs.
+test.
+We use the Logit model for these experiments since it offers the best performance and is easy to train.
+The results from this exercise are shown in Figure A4.
+As we can see, when we use the 256-dimensional embeddings, approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance.
+However, when we use the 3072E-dimensional embedding, there is a clear over-fitting [^20] Over-fitting refers to the phenomenon that the accuracy on the training set is consistently higher than the accuracy on the test set.
+It occurs when a machine learning model learns the details and noise in the training data to the extent that it negatively impacts the model’s performance on new, unseen data.
+This often happens because the model becomes too complex or because the training set is too small, causing the model to memorize specific examples and noise rather than general patterns.
+issue even if we use the simplest Logit model and 95% of the data for training.
+This suggests that for larger embedding vectors, we need much larger training datasets.
+[^21] In principle, overfitting can also be due to model complexity.
+However, this is unlikely to be the case in our setting since we are using a simple Logit model with linear features.
+C.4 Dimensionality Reduction and Regularization
+In this section, we examine whether employing dimensionality reduction techniques and/or regularization can improve performance and reduce overfitting.
+C.4.1 PCA for Embeddings
+One can observe that OpenAI-256E in Figure A4(a) does not suffer from the over-fitting in contrast to OpenAI-3072E as shown in Figure A4(b).
+Thus, one straightforward method one can try to attack the over-fitting issue is reducing the dimension of embeddings using PCA or directly using lower-dimensional OpenAI embeddings.
+However, as we have discussed that OpenAI-256E has a worse performance than OpenAI-3072E regarding the accuracy in the testing dataset, we turn to PCA to see if PCA-reduced embeddings can improve the performance by solving the over-fitting issue.
+Specifically, we first use PCA to reduce 3072-dimensional OpenAI embeddings.
+Then, we concatenate two reduced embeddings from healing pairs and use Logit for the downstream classification task.
+The result is displayed in Figure A5, which shows that as the dimension is reduced, although there is a smaller accuracy gap between the training and test set, the accuracy on both datasets decreases, indicating that PCA does not work for our task, which prefers higher testing accuracy.
+Figure A5: Accuracy of the Logit classification model with varying PCA-reduced OpenAI embeddings as inputs.
+C.4.2 $L^{1}$, $L^{2}$ Regularization
+[…]
+Figure A6: Accuracy of the Logit classification model with varying $L^{1}$ (left) and $L^{2}$ (right) regularization weights. X-axis is the regularization weight defined as $1/\texttt{C}$.
+In sum, we find that neither PCA nor regularization improves the performance accuracy on the test data, which suggests that over-fitting is unlikely to be due to learning noise in the training data.
+Our analysis suggests that we need more A/B tests to learn the relationship between a headline’s text embeddings and its attractiveness effectively.
+In summary, these comparisons suggest that data samples from Upworthy are insufficient to unlock the potential of LLMs fully.
+Nevertheless, diverting more and more traffic to A/B tests is costly, and this observation is one of the motivations for our approach later in §$\S$4.
+C.5 Potential Threats and Robustness Checks
+Next, we discuss two potential threats to the inference from our analysis here and our robustness checks to alleviate concerns around these threats.
+First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus – which could potentially invalidate our analysis.
+However, this is unlikely for several reasons.
+First, it is improbable that OpenAI would have specifically downloaded these Upworthy CSV files (Comma-Separated Values files) dataset from a subfolder in Upworthy Research Archive.
 Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than paired headlines with winner information.
@@
 However, the observed accuracy was approximately 50%, indicating that the embeddings do not contain such information, confirming our assumption.
-Finally, as discussed earlier, we consider both a Logit model (with linear features) and a MLP for the classification task (i.e., to predict which headline is catchier).
-The MLP model employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting.
-The MLP network is trained using the stochastic gradient descent algorithm to minimize binary cross-entropy.
-We perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration.
-For the OpenAI-256E model, we used the following MLP architecture: Inputs(512)-fc(256)-ReLU-dropout(0.5)-fc(1)-sigmoid, and the training hyperparameters were set with an $L_{2}$ penalty of $10^{-4}$ and 74 training epochs.
-For the OpenAI-3072E model, the MLP architecture was Inputs(6144)-fc(2048)-ReLU-dropout(0.5)-fc(1)-sigmoid, with an $L_{2}$ penalty of $10^{-4}$ and 32 training epochs.
-Please see Web Appendix A.2 for additional details of this model.
-3.2.3 Performance Analysis of Embedding-based Approaches
-We now present the performance of the different embedding-based classification models in Table 6.
-We also include a benchmark using logistic regression with Word2Vec word embeddings for a comprehensive comparison.
-We use Word2Vec as it was one of the best embedding models before the wide adoption of Transformer in NLP.
-Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information.
-We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec.
-To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
-Embedding Model Classification Model Accuracy
-OpenAI-256E Logit 77.56%
-MLP 78.29%
-OpenAI-3072E Logit 82.33%
-MLP 82.06%
-Word2Vec-256E Logit 68.58%
-Word2Vec-3072E Logit 68.84%
-Table 6: The performance of different embedding and classification models on test data.
-All models use the same training data, and accuracy is evaluated on the test data.
-Note that Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents.
-Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average of the accuracy from three reruns: 78.27%, 78.29% and 78.32% for OpenAI-256E + MLP and 82.43%, 81.67% and 82.09% for OpenAI-3072E + MLP.
-We find that classification models using the 3072-dimensional text embedding from OpenAI achieve more than 80% accuracy, demonstrating superior performance compared to prompt-based approaches.
-This performance can be attributed to the fact that embedding-based classification models are trained on a customized dataset, allowing it to learn decision rules tailored precisely fine-tuned to the headline selection task.
-Because these models can leverage a larger labeled dataset during training, they are more generalizable and robust.
-In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations, resulting in sub-optimal performance compared to a task-specific trained classifier.
-Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus.
-Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
-Next, we discuss the detailed comparison between models using the embedding technique.
-We first compare the results under different embedding dimensions.
-Typically, more dimensions give greater quality encoding and, thus, higher prediction accuracy.
-However, there is some limit beyond which there are diminishing returns.
-In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 4.3%.
-It is notable that OpenAI embeddings with Logit outperform Word2Vec significantly by around 9.0% and 13.5% at the same dimensions of 256 and 3072, respectively.
-These observations are unsurprising since OpenAI’s LLMs are state-of-the-art language models that naturally perform better than the classic Word2Vec embeddings.
-A more striking observation is that classification model complexity does not improve the prediction (conditional on the same inputs).
-Specifically, there is no significant accuracy difference between MLP and Logit when using OpenAI-256E or OpenAI-3072E.
-This observation suggests that a linear classifier on top of fixed LLM embeddings could effectively capture the attractiveness of content.
-An interesting question is why text embeddings obtained from models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for our downstream classification task.
-(saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text.
-Thus, the embeddings generated from a next-word prediction task inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
-We further investigate two specific aspects of embedding-based approaches.
-First, we examine the conditions under which these approaches fail, and second, we quantify the role of the number of training samples on predictive accuracy.
-For both these analyses, we use logistic regression with linear features as our classification model (since this is the model with the highest accuracy and there is no significant improvement in the accuracy of MLP).
-To address the first question, we first calculate the cosine similarity based on the text embeddings for each pair of headlines.
-This metric captures the similarity between two vectors in inner product space and is defined as the dot product of the vectors divided by the product of their lengths, i.e., ${e_{1}\cdot e_{2}}/(\|e_{1}\|\|e_{2}\|)$, where $e_{1}\cdot e_{2}$ is the dot product of the embedding vectors, and $\|e_{1}\|$ and $\|e_{2}\|$ are the their Euclidean norms.
-Figure 5(a) shows the distribution of cosine similarity across all headline pairs, whereas Figure 5(b) shows the distributions of cosine similarity separately for two cases: (1) cases where our embedding-based classification correctly predicts which headline is catchier, and (2) cases where our embedding-based classification incorrectly predicts which headline is catchier.
-We observe a rightward shift in the distribution of incorrect predictions (compared to the distribution of correct predictions).
-This suggests that incorrect predictions are more common when the headlines are very similar to each other (i.e., higher cosine similarity).
-There are two possible reasons for this.
-First, in cases where the headline pairs are very similar, there is a higher chance of Type 1 error (in the t-tests based on our A/B test data), i.e., the headlines may not be significantly different in fact.
-[^14] Conducting multiple tests in our data can inflate the overall Type I error rate, known as the multiple testing problem.
-This means there is a higher likelihood of encountering at least one Type I error.
-Similar headlines are more likely to lead to these errors because the actual difference in their CTRs is likely to be small.
-This increases the probability that variation in randomly sampled impressions and clicks in A/B tests might lead to significant results by chance, even when there is no true effect.
-Second, text embeddings may not be able to capture the differences between two headlines when they are similar, i.e., the embeddings may also be very similar.
-As a result, it is hard for the classification model to differentiate between them.
-[^15] Note that it is unlikely that this phenomenon is due to problems/issues with the classification model because both Logit and MLP achieve nearly identical performance.
-(a) Performance of OpenAI-256E
-(b) Performance of OpenAI-3072E
-Figure 4: Performance of Logit classification models with OpenAI embeddings under varying training dataset sizes and different embedding sizes.
-Next, to address the second question, we conduct experiments varying the proportion of data used for training vs.
-test.
-We use the Logit model for these experiments since it offers the best performance and is easy to train.
-The results from this exercise are shown in Figure 4.
-As we can see, when we use the 256-dimensional embeddings, approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high performance.
-However, when we use the 3072E-dimensional embedding, there is a clear over-fitting issue even if we use the simplest Logit model and 100% of the data for training.
-This suggests that for larger embedding vectors, we need much larger training datasets.
-[^16] It is unlikely that this is due to issues with the classification model since the Logit model is inherently parsimonious.
-To further investigate this issue, we experimented with both $L_{1}$ and $L_{2}$ regularization to control over-fitting.
-[^17] This was done using the penalty and C control parameters in the LogisticRegression function from sklearn package.
-Unfortunately, regularization did not improve the accuracy on the test data, suggesting that over-fitting is unlikely to be due to learning noise in the training data.
-Recall that the average headline in the Upworthy tests received 3,569 impressions and 54.26 clicks.
-Our analysis suggests that we need much larger A/B tests to learn the relationship between a headline’s text embeddings and its attractiveness effectively.
-In summary, these comparisons suggest that data samples from Upworthy are insufficient to unlock the potential of LLMs fully.
-Nevertheless, diverting more and more traffic to A/B tests is costly, and this observation is one of the motivations for our approach later in Section 4.
-(a) Cosine similarity distribution over all pairs in training and test sets
-(b) Cosine similarity over incorrect/correct predicted test pairs
-Figure 5: The left panel shows the distribution of cosine similarity for all pairs.
-The right panel shows the distributions of cosine similarity for both incorrect and correct predictions using the logit classification model with OpenAI-3072E embeddings.
-Finally, we discuss the time and compute costs of embedding-based approaches, and contrast them with prompt-based approaches.
-We find that obtaining text embeddings through the OpenAI API is extremely efficient in terms of time usage and is also more economical compared to direct prompt-based interactions.
-When using the embedding API, data transmission involves smaller, fixed-size vectors instead of variable-length text, which reduces data overhead and accelerates processing speed.
-We spent less than \$1 to obtain the text embedding for all 77,245 headlines, while the prompt experiment with only 10,000 pairs cost us around \$20.181818At the time of this study (May 2024), GPT-4 (turbo version) prompt costs \$10 per 1M input tokens and \$30 per 1M output tokens.
-Further, the embeddings used in our study, text-embedding-3-large, only cost \$0.13 per 1M input tokens and there is no output token cost associated with embeddings.
-Our entire dataset comprises 77,245 headlines, with an average of 19 tokens per headline, resulting in a total embedding cost of approximately \$0.18.
-In addition, we found that it is much faster to process embeddings (more than 3,000 pairs per minute), while the prompt experiment is much slower (can only test around 83 pairs per minute).
-This efficiency can translate into lower operational costs and help with easier integration into real-time applications.
-Another key advantage of the embedding-based approach is its transparency and robustness.
-Using simple models like logistic regression with text embeddings from OpenAI can give researchers and managers insight into the training process and source of performance improvements.
-This transparency is crucial in business settings, where a clear understanding of a model’s behavior and performance is often necessary to mitigate potential risks and biases.
-In contrast, in-context learning is a black-box approach where the user/manager cannot obtain any insight into the source of the performance improvement.
-3.3 Fine-Tuning Open-Sourced LLMs with LoRA
-In Section 3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself.
-The performance of these embedding-based approaches reached 82% accuracy without modifying any parameters in the LLM.
-This raises a natural question: if we can train LLMs directly, can we achieve better performance?
-However, training LLMs from scratch is costly and resource-intensive.
-Therefore, in this section, we turn to fine-tuning techniques, which involve partially adjusting the parameters in the LLMs to improve performance while keeping the computational cost manageable.
-Note that to fine-tune LLMs, we need to be able to access the underlying model parameters.
-This is only feasible for open-source LLMs, like BERT and LLama, and not for proprietary LLMs like OpenAI’s GPT and Google’s Gemini.
-The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
-For the fine-tuning task, we use Meta’s Llama-3 (meta_llama_3) as our primary base model because it is regarded as one of the most advanced open-source LLMs to date (llama3_model_card).
-zhao_etal_2024 consider a large-scale experiment where they fine-tune different open-source LLMs and show that for NLP tasks, including news headline generation, Llama-3 offers one of the best performance.
-[^19] Before the widespread adoption of the term “LLM”, BERT was considered among the best language models available, and is still widely used as a benchmark.
-To provide a comparison with these earlier models, we also fine-tuned DistilBERT (sanh2019distilbert)—a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (devlin2018bert).
-Details on the implementation and performance of the fine-tuned DistilBERT are shown in Web Appendix A.3.
-Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b).
-We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see Section 3.3.4 for details).
-Hence, we do not consider fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated.
-Our numerical experiment shows LoRA fine-tuned Llama-3-8b achieves nearly 83% accuracy.
-The rest of this section is organized as follows.
-We first introduce the Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 in Section 3.3.1.
-Next, we present the LoRA method for efficient fine-tuning in Section 3.3.2.
-The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning.
-Finally, we detail our implementation in Section 3.3.3 and report our results in Section 3.3.4.
-3.3.1 Transformer and Llama Architecture
+Nevertheless, if a firm is concerned that its data is a part of public repositories or is concerned about privacy issues around submitting headlines to OpenAI to obtain embeddings, then instead of using embedding-based approaches, it can rely on fine-tuning based approaches; see §$\S$3.3 for details.
+Second, one may be concerned that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created, making our classification task invalid.
+To address this concern, we conducted a robustness check by splitting the dataset based on the creation time of the headlines (e.g., the initial 80% of the data, in chronological order, as the training set, and the subsequent 20% as the test set).
+We found its accuracy was 81.50% (close to the best embedding accuracy of 82.50%), indicating that time shift preference is not a concern in our data.
+Appendix D Appendix to LoRA Fine-Tuning
+D.1 Transformer and Llama Architecture
 The Transformer model, introduced by (vaswani2017attention), has revolutionized the field of NLP with its innovative attention mechanism and highly parallelizable architecture.
 Unlike traditional recurrent neural network (RNN) based models, the Transformer uses self-attention mechanisms to process input data.
-This shift allows for greater computational efficiency and significantly improved performance across a range of NLP tasks.
-As illustrated in the right portion of Figure 6, the Transformer block comprises multiple layers, each featuring two primary components: a multi-head self-attention mechanism and a position-wise fully connected feed-forward network.
+This shift allows for greater computational efficiency and significantly improved performance across various NLP tasks.
+As illustrated in the right portion of Figure A7, the Transformer block comprises multiple layers, each featuring two primary components: a multi-head self-attention mechanism and a position-wise fully connected feed-forward network.
 Key enhancements include layer normalization (ba2016layer) and the use of residual connections (he2016deep), which improve training stability and help prevent issues such as vanishing gradients.
 While some specifics are omitted to maintain brevity and readability, it’s important to note that such architectures are subject to rapid evolution and vary across different LLMs.
-Figure 6: Illustration of the Transformer block and architecture typical in most LLMs.
+Figure A7: Illustration of the Transformer block and architecture typical in most LLMs.
 The left portion of the figure displays the overall architecture of an LLM, highlighting the layers of Transformer blocks and the connections that represent the attention mask.
@@
 Additionally, only two layers are shown for simplicity; actual models may contain many more layers and connections, which are not depicted here to maintain clarity and focus on key components.
-The right portion provides a detailed view of a single Transformer block, internal components included.
+The right portion provides a detailed view of a single Transformer block, including internal components.
 We omit the depiction of multi-head attention for brevity.
@@
 Each embedding here simply corresponds to one token without paying attention to the whole input sequence.
-The text embeddings used in the previous section are essentially the contextualized embeddings, which are generated by passing the tokens through multiple layers of Transformers, capturing rich contextual information and summarizing the whole input sequence.
+The text embeddings used in the previous section are essentially the contextualized embeddings, which are generated by passing the tokens through multiple layers of Transformers, capturing rich contextual information, and summarizing the whole input sequence.
 forming a matrix $X$ where each row corresponds to the embedding of a token.
@@
 It is worth noting that LLMs do not necessarily operate in a conversational manner, nor do they always use language as their output.
-As we will demonstrate in Section 3.3.3, we employ the Llama model as a classifier, where the output is the distribution of predicted classes.
+As we will demonstrate in §$\S$3.3.2, we employ the Llama model as a classifier, where the output is the distribution of predicted classes.
 In applications like conversational AI, exemplified by GPT, the contextualized embeddings from the last token of the last layer of the Transformer are fed into a linear layer followed by a softmax function, the configuration of which transforms these embeddings into a probability distribution over the vocabulary, enabling the model to predict the next token in the sequence.
-3.3.2 LoRA Fine-Tuning Technique
-Traditional fine-tuning of LLMs involves updating a substantial number of parameters, which can be computationally expensive and memory-intensive.
-To efficiently fine-tune the Llama-3 model on hardware with limited GPU memory, we employ Low-Rank Adaptation (LoRA), a popular Parameter-Efficient Fine-Tuning (PEFT) method developed by (hu2021lora), enabling practical fine-tuning of large-scale LLMs on medium-scale hardware.
-[^21] There are many alternatives for PEFT, like Prefix Tuning (li2021prefix), Adapter Layers (houlsby2019parameter), and BitFit (zaken2021bitfit).
-We select LoRA due to its excellent balance between simplicity and efficiency, making it the top choice for fine-tuning.
-LoRA operates under the assumption that updates during model adaptation exhibit an intrinsic low-rank property and introduce a set of low-rank trainable matrices into each layer of the Transformer model.
-This approach leverages the low-rank decomposition, which significantly reduces both memory usage and computational time through a significantly reduced number of trainable parameters.
-We illustrate LoRA in Figure 7.
-We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$.
-We use $d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices.
-We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
-$W^{\prime}=W+\Delta W,\quad\mathrm{with}\quad\Delta W=AB,\quad\mathrm{where}\quad A\in\mathbb{R}^{d\times r},\quad B\in\mathbb{R}^{r\times d_{W}}.$
-Figure 7: Illustration of LoRA.
-LoRA essentially introduces the low-rank matrices $A$ and $B$ to approximate the weight matrix change $\Delta W$ after the fine-tuning.
-These low-rank matrices are updated during the fine-tuning process, while the original weight matrix $W$ remains frozen.
-This strategy significantly reduces the number of trainable parameters from $d\times d_{W}$ to $r\times(d+d_{W})$ because the rank $r$ is typically set to a small value, such as 4 or 8, to balance the trade-off between model capacity and computational efficiency.
-3.3.3 Implementation Details
-We use exactly the same training and test sets as in Section 3.2, but the inputs here are the original headlines with the text information.
-To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them.
-For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
-We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 8.
-Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence.
-Thus, the last token in the sequence, being informed by all prior tokens, is used for classification.
-This embedding vector of the last token, output by the last Transformer layer and having a dimension of 4096, is processed through a linear layer producing an output of two dimensions representing the logits.
-Then the logits are transformed into probabilities by the softmax function, indicating the likelihood of each headline being more impactful.
-The softmax function is expressed as:
-$\text{Softmax}(Z)_{i}=\frac{e^{z_{i}}}{\sum_{j}e^{z_{j}}}$
-where $Z$ is the vector of logits, and $i$ indexes the specific class. During fine-tuning, our objective is to minimize the cross-entropy loss on the training set.
-Figure 8: The classification pipeline using the Llama model.
-Headlines are first tokenized and each token is converted into an embedding vector.
-For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration.
-The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens.
-Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity.
-The output from the last token is fed into a classification head, which computes the probability that one headline is more impactful than the other.
-We implement LoRA Fine-tuning using the PEFT library developed by Hugging Face.
-Our LoRA configuration settings include setting the dimension of the low-rank matrices, denoted as $r$, to 4.
-The lora_alpha parameter, acting as a scaling factor for the low-rank matrices, is set to 4.
-This parameter controls how much the learned low rank matrix would be weighted.
-Adjusting lora_alpha allows us to control the adaptation strength, enabling the model to integrate new information effectively without losing its pre-trained knowledge.
-Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora).
-Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$.
-We set the dimension of the low-rank matrices $r$ to 4.
-To mitigate the overfitting, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices.
-Furthermore, we enable 16-bit mixed precision training.
-Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly.
-This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
-With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which takes only 0.023% of the total 7,506,636,800 parameters.
-We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training.
-A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process.
-This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
-Lastly, we want to briefly discuss the GPU requirement and monetary cost of implementing LoRA.
-The GPU requirements for fine-tuning depend on factors such as model size, batch size, optimization algorithm, and floating-point precision.
-In our experiment, without LoRA, full-parameter fine-tuning would require over 100 GB of GPU memory.
-However, LoRA fine-tuning only requires 30 GB of GPU memory.
-To provide a more straightforward demonstration of the benefit of the saved GPU requirement, we showcase two GPU models that are commonly used for LLM training.
-Without LoRA, the NVIDIA H100 GPU with nearly the largest GPU memory, 80 GB is insufficient for fine-tuning.
-In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient.
-The price of an NVIDIA H100 GPU is ten times higher than an NVIDIA V100 GPU in May 2024.
-Thus, using LoRA significantly reduces the cost, making fine-tuning more accessible to researchers and practitioners.
-3.3.4 Performance Analysis of Fine-tuning-based Approaches
-We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 9.
-We observe a slightly better performance of 82.79% accuracy compared to the best result (82.33% accuracy) using OpenAI embeddings in Section 3.2.
-It is important to note that, if compared under the same LLM, LoRA fine-tuned Llama-3-8b outperforms the logistic classifier using embeddings from Llama-3-8b (around 77% accuracy).
-Figure 9: Loss curve and accuracy curve as the number of training epochs increases.
-We repeat the fine-tuning process three times and display the average loss and accuracy.
-For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits.
-The final accuracies of the three replicates on the test set are 83.04%, 82.73%, and 82.59% respectively.
-We add the best performance (dash line) obtained using OpenAI’s embedding in Section 3.2 for a comparison.
-From Figure 9, we observe that the loss evaluated on the training set is consistently lower than the loss on the test set, and the accuracy on the training set is consistently higher than the accuracy on the test set, indicating overfitting.
-Overfitting refers to the phenomenon that occurs when a machine learning model learns the details and noise in the training data to the extent that it negatively impacts the model’s performance on new, unseen data.
-This often happens because the model becomes too complex or because the training set is too small, causing the model to memorize specific examples and noise rather than general patterns.
-As a result, the model performs exceptionally well on the training data but fails to generalize, leading to relatively poor performance on test data.
-In our case, the Llama-3-8b model is inherently complex.
-Although using LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex.
-Common approaches to address overfitting include reducing model complexity, adding regularization, and increasing the dataset size.
-We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay, in an attempt to mitigate overfitting.
-However, these modifications generally reduce performance on the training set without yielding significant improvements on the test set.
-We suspect that overfitting could be further alleviated, and test performance improved, with a larger and more diverse dataset.
-The insight for practitioners is that continuously updating the model with new data, such as through A/B testing on a news website to collect more pairwise comparison data, may lead to better generalization and performance over time.
-Instead of fine-tuning a model version and continuously using it, we suggest that companies keep updating the model as they collect larger datasets.
-After fine-tuning the model, several cost-efficient techniques are available for deploying it for inference, significantly reducing computational costs and hardware requirements.
-For instance, llama.cpp allows LLM models to be deployed on machines without a GPU, making implementation more affordable.
-Moreover, model compression techniques such as quantization, sparsification, and knowledge distillation can drastically cut inference computation costs while maintaining, or only slightly affecting, model performance.
-Our approach ensures that the fine-tuned model is not only cost-efficient but also robust and secure when hosted on our own servers.
-For a comprehensive survey on efficient LLM deployment and inference techniques, please refer to (zhou2024survey).
-4 Our Approach: LLM-Assisted Online Learning Algorithm
-So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 82% for the embedding-based and LoRA fine-tuning approaches).
-The main advantage of these pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it.
-Thus, there is no need to use any traffic for experimentation.
-The downside, of course, is that these models are not fully accurate—in around 18% of the cases, they are likely to mispredict which headline is better.
-This can of course, lead to lower clicks and revenues for the firm.
-[^22] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth.
-However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples.
-Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
-On the other hand, we have the current status quo, which is the use of experimentation-based approaches to choose headlines/content.
-As mentioned in Section 1, there are two broad experimentation-based methods—(1) A/B tests and (2) Online learning algorithms or bandits.
-In particular, the standard practice at Upworthy during the time of the data collection was A/B tests, where the firm allocated a fixed amount to different headlines and then chose the best-performing one for the remaining traffic.
-This approach is also known as Explore and Commit (E&C) and is commonly used in the industry due to its simplicity and low engineering cost; see chapter six of (lattimore2020bandit).
-The main benefit of this approach is that, with a sufficiently large amount of traffic allocated to the A/B test, the resulting inference is unbiased; as such it is the gold-standard approach for identifying the best arm (or headline in this case).
-The downside, of course, is that by assigning traffic to all the headlines during the exploration phase or A/B test, the firm incurs high regret (i.e., low profits since some of headlines in the test are likely have low CTRs).
-Therefore, many modern technology and media firms such as NYTimes and Yahoo use online learning algorithms or bandits that leverage the explore-exploit paradigm to lower the experimentation cost while still learning (lattimore2020bandit).
-These algorithms move traffic away from poorly performing arms/headlines dynamically in each period, based on the observed performance of each arm till that period.
-Because of this greedy behavior, they tend to switch to the best-performing arms quickly, and incur lower regret.
-Nevertheless, one drawback of these methods is that they start with the assumption that all headlines are equally good, i.e., they suffer from a cold-start problem, which can result in wasting traffic on the weaker headlines early on.
-In this section, we present a framework that combines the benefits of LLM-based models with experimentation-based methods.
-Our key idea is that firms can address the cold-start problem in content experiments by leveraging predictions from LLM-based approaches as bayesian priors in a first step and combine them with online learning algorithm in the second step.
-Thus, our approach, termed LLM-Assisted Online Learning Algorithm (LOLA), continuously optimizes traffic allocation based on LLM predictions together with the data collected from the online experiment in real-time.
-The rest of this section is organized as follows.
-In Section 4.1, we discuss the general LOLA framework, where we formalize the media firm’s problem and then present our algorithm.
-Next, in Section 4.2, we discuss the implementation of our proposed algorithm, benchmarks, and the experimental setup.
-Finally, we present our numerical results in Section 4.3.
-4.1 Bandit Framework and LOLA
-In the following, we first present the Multi-Armed Bandits (MAB) framework to formalize our problem and then discuss our proposed approach.
-MAB is a class of sequential decision-making problems where a learner must choose between multiple arms over periods to maximize cumulative reward.
-In our setting, different arms refer to different headlines associated with the same content, and periods refer to impressions where different headlines with the same content are displayed.
-Let $[K]:=\{1,2,\dots,K\}$ represent the set of $K$ arms, and $[T]:=\{1,2,\dots,T\}$ represent the entire decision horizon for a test.
-At each time step $t$, the learner selects an arm $a_{t}\in[K]$ and receives a reward $r_{a_{t}}$ drawn from a probability distribution specific to the chosen arm $a_{t}$.
-In our setting, we assume each arm has a constant click-through rate (CTR), and the reward is the click feedback, which follows the Bernoulli distribution with the probability equal to the CTR [^23] This reflects a stochastic bandit setting rather than a Bayesian bandit setting.
-However, it is important to note that our proposed algorithms can also be applied in a Bayesian setting by changing the definition of regret and the process of generating reward feedback.
-For completeness, we also present a Thompson sampling version of our proposed LOLA algorithm in Web Appendix C..
-We first introduce the most common framework, regret minimization.
-The goal in this framework is to minimize cumulative regret, which is defined as the difference between the reward obtained by always pulling the optimal arm and the reward actually accumulated by the learner.
-Mathematically, the regret $R_{T}$ after $T$ trials is given by:
-$R_{T}=\max_{k\in[K]}\sum_{t=1}^{T}r_{a^{*}}-\sum_{t=1}^{T}r_{a_{t}},$
-where $a^{*}=\arg\max_{k\in[K]}\mathbb{E}[r_{k}]$ is the optimal arm with the highest expected reward.
-As a complement to regret minimization, we also tackle the problem using the Best Arm Identification (BAI) framework.
-The goal of BAI is to identify the arm with the highest reward within a given limited number of pulls.
-Formally, given a fixed budget $B$, we aim to find the arm $a^{*}$ that maximizes the expected reward after $B$ trials.
-Before we introduce our LLM-assisted online learning algorithm, we introduce the widely applied pure learning algorithm without LLMs, the Upper Confidence Bound (UCB) algorithm.
-The UCB stands out because it is provably asymptotically optimal.
-The UCB policy is straightforward and selects the arm that maximizes upper confidence bound on the estimated rewards, specifically:
-$a_{t}=\arg\max_{k\in[K]}\left(\bar{\mu}^{t}_{k}+\sqrt{\frac{2\log(1/\delta)}{n^{t}_{k}}}\right),$
-where $\bar{\mu}^{t}_{k}$ is the estimated mean reward of arm $k$ up to time $t$, $n^{t}_{k}$ is the number of times arm $k$ has been played up to time $t$, and $\delta$ is called the confidence level, which quantifies the degree of certainty.
-The intuition behind UCB’s effectiveness in balancing exploration and exploitation lies in its construction.
-By adding a confidence term $\sqrt{{2\log(1/\delta)}/{n^{t}_{k}}}$ to the estimated mean reward $\bar{\mu}_{k}$, UCB ensures that arms with fewer selections (higher uncertainty) are given a chance to be explored.
-As time progresses and more data is collected, the confidence term diminishes, allowing the algorithm to exploit arms with higher observed rewards.
-This dynamic adjustment enables the UCB algorithm to systematically explore under-explored arms while exploiting arms with high rewards, thus effectively addressing the exploration-exploitation trade-off inherent in the MAB problem.
-However, we need to set the key hyperparameter $\delta$ such that it should be small enough to ensure optimality with high probability, but not too small that suboptimal arms are over-explored.
-Typically, one can set the confidence term as $\alpha\sqrt{\log t/n^{t}_{k}}$ and fine-tune the scale term $\alpha$ for specific problems.
-Under the current framework, we may have more than two headlines in one test (i.e., $K>2$), rendering the previous LLM-based binary classification models inapplicable.
-To compare multiple arms, we adapt the previous LLM-based classification models to predict CTR by changing the loss function from log loss to mean squared error (MSE) while keeping the training and fine-tuning procedures unchanged.
-Using CTR as the outcome measure does not adversely affect decision-making performance.
-For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.56% accuracy obtained with logistic regression.
-This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
-Given the above formalization of the decision process and the LLM-based CTR prediction model, the remaining key challenge is to design a decision policy that incorporates these LLM-based CTR predictions.
-Inspired by the 2-Upper Confidence Bounds (2-UCBs) policy designed by (gur2022adaptive), we propose the following LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs) algorithm.
-Essentially, we treat the LLM’s CTR predictions as auxiliary information prior to the start of online experiments, equating this auxiliary CTR information to additional impressions and click outcomes for each arm.
-Algorithm 1 LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs)
-1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
-2:Hyperparameter Fine-Tuning: Use another subset of data to select two hyperparameters for the algorithm, $\alpha\in\mathbb{R}^{+}$ for controlling the upper bound, $n^{\text{a}ux}$ as the LLM’s equivalent auxiliary sample size.
-3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$, LLM-based CTR prediciton $\bar{\mu}^{\text{a}ux}_{k}\leftarrow\mathcal{M}(x_{k})$, regular CTR initialization $\bar{\mu}^{1}_{k}=0$, accumulated impressions $n^{1}_{k}\leftarrow 1$, and accumulated clicks $c^{1}_{k}\leftarrow 0$, for all arms $k\in[K]$.
-4:for $t=1$ to $T$ do
-5:     Calculate the first UCB for all arms $k\in[K]$: $U_{k}^{1}=\bar{\mu}^{t}_{k}+\alpha\sqrt{\frac{\log t}{n^{t}_{k}}}$.
-6:     Calculate the second UCB for all arms $k\in[K]$: $U_{k}^{2}=\frac{c^{t}_{k}+\bar{\mu}^{\text{a}ux}_{k}n^{\text{a}ux}}{n^{t}_{k}+n^{\text{a}ux}}+\alpha\sqrt{\frac{\log t}{n^{t}_{k}+n^{\text{a}ux}}}.$
-7:     Play the arm $a_{t}=\arg\max_{k\in[K]}\min\{U_{k}^{1},U_{k}^{2}\}$.
-8:     Observe the payoff $r_{a_{t}}$
-9:     Update $n^{t+1}_{a_{t}}\leftarrow n^{t}_{a_{t}}+1$, $c^{t+1}_{a_{t}}\leftarrow c^{t}_{a_{t}}+r_{a_{t}}$, and $\bar{\mu}^{t+1}_{a_{t}}\leftarrow\frac{c^{t+1}_{a_{t}}}{n_{a_{t}}^{t+1}}$.
-10:end for
-In the training phase of the LLM-2UCBs algorithm, we can utilize various LLM-based methods, including LoRA fine-tuning using the original text information $x$, and embedding-based CTR prediction models where $x$ represents the embedding vector.
-The feature vector $x$ can also include additional information, such as user-specific data.
-However, since Upworthy does not collect user information and ultimately recommends the same headline to all users, incorporating user information into the algorithm is unnecessary in this specific context.
-Nonetheless, we emphasize that our algorithm is flexible and can be adapted to incorporate user information when it is available.
-There are two hyperparameters in the LLM-2UCBs.
-The upper bound control hyperparameter, $\alpha$, also shows up in the regular UCB algorithm.
-The hyperparameter $n^{\text{aux}}$ is specific to our proposed algorithm and indicates the equivalent auxiliary samples for initialization.
-At a high level, we can regard the LLM predictions as prior information or the information flow before the algorithm starts.
-This prediction is approximately equivalent to initially generating $n^{\text{aux}}$ impressions for each arm and observing their outcomes to obtain the sample average CTR, $\bar{\mu}^{\text{aux}}_{k}$.
-The hyperparameter fine-tuning process is simple.
-We can utilize another randomly sampled dataset to test combinations of two hyperparameters and select the combination that yields the highest accumulated rewards.
-During the online learning phase, the key step is balancing the exploration and exploitation trade-off.
-We capture this trade-off using the 2-UCBs rule.
-Essentially, we have two valid UCBs: $U_{k}^{1}$ is the regular UCB, which uses observed impressions and clicks to construct the mean reward estimator and adds a regular upper bound.
-The second UCB, $U_{k}^{2}$, incorporates LLM predictions.
-The calculation of $U_{k}^{2}$ is straightforward.
-The mean reward component is simply the sample average, taking into account the “auxiliary” samples generated by the LLM before the learning phase begins.
-These auxiliary samples are imaginary and approximate the richness of information contained in the LLM prediction.
-The next step is to choose the smaller of the two UCBs, $\min\{U_{k}^{1},U_{k}^{2}\}$, as the final UCB for selecting the arm.
-Note that if $n^{\text{aux}}=0$, the 2-UCBs rule is the same as the standard UCB rule.
-Conversely, when $n^{\text{aux}}\rightarrow\infty$, the 2-UCBs rule is equivalent to a pure exploitation policy (a.k.a.
-greedy policy) using the LLM-based CTR prediction.
-After playing the arm and observing the outcome, we update the impressions, clicks, and mean reward accordingly, following standard bandit algorithm procedures.
-The 2-UCB algorithm offers several advantages.
-First, its nonparametric nature means it does not assume a linear relationship, making it flexible for various applications.
-Second, it is easy to implement, requiring the tuning of only two hyperparameters.
-Third, it can be applied to any LLM-based model.
-Notice that we do not use the contextual information $x$ during the online learning phase.
-The primary reason for this is to avoid updating the LLM prediction model during the learning phase, which would require significant engineering effort due to the large-scale nature of training LLMs.
-However, if engineering costs are not a concern, and for the completeness of this paper, we provide a linear bandit algorithm based on (chu2011contextual), taking into account LLM-based information and user features for personalized recommendations.
-Despite this, the linear bandit has limitations due to its linear assumption.
-Therefore, we propose another potential solution based on the FALCON algorithm (simchi2022bypassing).
-We discuss the pros and cons of these bandit algorithms in Appendix B.
-Note that this framework is general and can accommodate any number of headlines and does not rely on binary classification and/or assume that differences between headlines are significant.
-As such, in the rest of the paper, we utilize the full dataset from Upworthy to test the performance of our proposed algorithm and benchmark approaches (unlike the analysis in Section 3, where we focus only on pairs of headlines with significant differences).
-4.2 Benchmarks and Implementation Details in Regret Minimization
-In this section, we introduce each of the following algorithms and their implementations in detail, as we will compare them in the next subsection.
-  • $\bullet$
-    Explore-then-Commit (E&C)
-  • $\bullet$
-    Pure LLM
-  • $\bullet$
-    Standard UCB (UCB)
-  • $\bullet$
-    LLM-Assisted 2-Upper Confidence Bounds (LLM-2UCBs)
-At a high level, we randomly split our data into three exclusive folds based on the test ID, ensuring that all headlines in the same test are in the same fold and that the same headline does not appear in different folds.
-Specifically, we use 70% of the data for training, 10% for fine-tuning, and 20% for testing.
-All hyperparameters are selected using the fine-tuning data, and the performance of the algorithms is evaluated on the test data.
-Note that only our proposed LLM-2UCBs and the pure LLM method utilize the training data.
-In our numerical study, we train the linear regression model with 3,072 dimensions text embeddings from OpenAI as inputs to predict CTRs
-Upworthy assigns traffic to headlines with equal probability during the experiment phase, [^24] Upworthy explicitly mentioned they didn’t conduct power analysis before the experiment, and based on the disclosed impression amount data in the same test, we make this claim with high confidence due to the almost same amount of impressions.
-and, at the end of the experiment phase, select the winner headline for all future impressions.
-This practice closely resembles the E&C algorithm in the bandit literature.
-We implement this method as a benchmark in our numerical experiments.
-The E&C method has a single hyperparameter to tune: the proportion of periods for exploration.
-We select the proportion from $\{0.1,0.2,0.3,0.4,0.5\}$ with the highest accumulated rewards.
-This ratio is fine-tuned for experiments with different total periods $T$, as a longer time horizon requires a smaller proportion of traffic for exploration, given a constant number of arms.
-Interestingly, the other two methods can be viewed as special cases of our proposed LLM-2UCBs algorithm.
-The Pure LLM (a.k.a pure exploitation, greedy algorithm) is a special case of LLM-2UCBs by setting $n^{\text{aux}}=\infty$, indicating complete trust in the LLM prediction results.
-This method fully relies on the LLM CTR prediction, directly selecting the headline with the highest predicted CTR throughout the entire horizon.
-Similarly, the standard UCB algorithm is a special case of LLM-2UCBs with $n^{\text{aux}}=0$, indicating no reliance on LLM CTR predictions.
-In our UCB implementation, we fine-tune the confidence control hyperparameter $\alpha$ from $\{0.02,0.04,\dots,0.10\}$ for different $T$.
-After fine-tuning, we find $\alpha=0.08$ in UCB gives the highest accumulated rewards.
-For LLM-2UCBs, we fine-tune the combination of hyperparameters $n^{\text{aux}}\in\{600,800,1000,1200,1400\}$ and $\alpha\in\{0.02,0.04,\dots,0.10\}$.
-After fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ in LLM-2UCBs.
-After the fine-tuning of algorithms, we evaluate all algorithms under the test dataset, which has never been seen for CTR training and algorithm tuning.
-Algorithms do not know this true testing CTR in advance.
-We use the average CTR in the test data as the true CTRs to generate the click feedback in the numerical simulations.
-We argue that using this CTR as the true CTR in the simulation is justified because each headline was tested with a large amount of traffic, resulting in a relatively narrow confidence interval at Upworthy.
-UCB and LLM-2UCBs are both asymptotically optimal in minimizing regret, which means there is no difference between these two algorithms when the number of time horizons $T$ goes to infinity, and the LLM-based prior information is not helpful in the long run.
-However, in practice, the time horizon is never infinite.
-Therefore, we compare different algorithms under different time horizons.
-Since there can be a varying number of headlines/arms in different tests, we scale the time horizon using the metric “Traffic/Impression per headline” (we use the notion $\tau$ to represent) from $\{50,100,200,400,600,800,1000\}$.
-Specifically, if there are three headlines in one test and the multiplier is $100$, we set $T=100\times 3=300$ for this test.
-It is important to investigate the small number multipliers because they also partially address the question of what happens if we are concerned about inference and can only use a small amount of traffic for adaptive experiments.
-4.3 Numerical Results
-We visualize the numerical results in Figure 10 and document the pairwise comparison of different algorithms, reporting the relative percentage reward improvement and the p-value from the t-tests.
-LLM-2UCBs generally performs the best among all algorithms.
-Notably, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm.
-For the other three approaches, which actively learn the CTR during the experiments, their average performance per period increases as the time horizon extends.
-This improvement is because the number of arms in the experiment remains constant, and with increasing periods, these algorithms have more opportunities to learn the CTRs, thus achieving better performance.
-When comparing LLM-2UCBs with Pure LLM, we find that LLM-2UCBs outperforms Pure LLM in most cases.
-An exception occurs when $\tau=50$, where there is no significant difference between LLM-2UCBs and Pure LLM.
-This is due to the constrained exploration ability of LLM-2UCBs in short planning horizons, leading it to perform similarly to using only LLM prior information.
-However, as $\tau$ increases to 1000, LLM-2UCBs significantly outperforms Pure LLM, generating 5.46% more clicks.
-This observation indicates that we cannot fully rely on pure LLM predictions alone.
-Comparing LLM-2UCBs with regular UCB, it is evident that leveraging LLM can always enhance UCB performance, although this improvement diminishes as the time horizon increases.
-This is due to the asymptotic optimality of the UCB algorithm, implying that as the time horizon approaches infinity, the initialization has a negligible effect on the average performance over an infinitely long period.
-E&C consistently underperforms compared to LLM-2UCBs because it neither leverages LLM predictions nor achieves asymptotic optimality.
-Even when $\tau=1000$, LLM-2UCBs is still better than E&C by 4.41%, which is a large improvement in this industry.
-E&C is also significantly worse than Pure LLM when $\tau\leq 400$, and shows no significant difference from Pure LLM when $\tau\geq 600$.
-This suggests that the current Upworthy practice is suboptimal.
-In summary, our findings underscore the superior performance of LLM-2UCBs, particularly as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.
-Figure 10: Average clicks per experiment per period under different time horizon multipliers.
-Note that the Y-axis captures the average clicks per test per period.
-For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
-The Y value is simply the average of this number $0.03$ over all tests.
-This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
-Parameter $\tau$ LLM-2UCBs vs Pure LLM LLM-2UCBs vs UCB LLM-2UCBs vs E&C Pure LLM vs UCB Pure LLM vs E&C UCB vs E&C
-50 $0.53$ $4.06$ $6.63$∗∗∗∗ $3.52$ $6.07$∗∗∗ $2.46$
-100 $1.04$ $5.01$∗∗∗∗ $6.20$∗∗∗∗ $3.93$∗∗ $5.11$∗∗∗∗ $1.13$
-200 $1.92$∗∗∗∗ $4.14$∗∗∗∗ $4.89$∗∗∗∗ $2.18$ $2.92$∗∗ $0.72$
-400 $3.57$∗∗∗∗ $2.51$∗∗∗ $6.36$∗∗∗∗ $-1.02$ $2.70$∗∗∗ $3.76$∗∗∗∗
-600 $3.78$∗∗∗∗ $1.74$∗∗ $3.60$∗∗∗∗ $-1.96$∗∗ $-0.17$ $1.83$∗∗∗
-800 $5.00$∗∗∗∗ $1.02$ $4.97$∗∗∗∗ $-3.79$∗∗∗∗ $-0.03$ $3.91$∗∗∗∗
-1000 $5.46$∗∗∗∗ $1.10$ $4.41$∗∗∗∗ $-4.13$∗∗∗∗ $-1.00$ $3.27$∗∗∗∗
-Table 7: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001.
-For example, under the impression per headline equal to 50, the average click per test of LLM-2UCBs and Pure LLM is 0.05149 and 0.05122, respectively; so the relative improvement is calculated as ${0.05149}/{0.05122}-1=0.53\%$.
-Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
-5 Conclusion
-We note that our approach is quite general and is agnostic to the exact nature of the LLM-based approach used.
-Thus, any of the LLM-based approaches discussed in Section 3 can be used in the first step of our approach.
-This is important since the LLM models and fine-tuning approaches continue to advance quickly; so the plug-and-play nature of our approach extends well to this setting.
-In addition, when considering personalization and accessing contextual information (e.g., readers’ features), we can incorporate this extra information as inputs to the LLM-based CTR prediction model and integrate it into contextual bandit algorithms.
-Specifically, we provide two variants of LOLA for this purpose in Web Appendix B.
-In conclusion, our study introduces LOLA, a novel framework that combines the predictive power of LLMs with the adaptive efficiency of online learning algorithms to enhance content experimentation.
-By leveraging LLM-based prediction models as priors, our approach minimizes regret in real time, leading to significant improvements in total reward/clicks.
-Furthermore, through the numerical analysis based on a large-scale dataset from Upworthy, we demonstrate that LOLA outperforms traditional A/B testing and pure online learning algorithms, particularly in settings with limited experimental traffic or a large number of arms.
-These conditions are prevalent in the current content industry, where platform users can easily create and share their content, such as on platforms like X and TikTok.
-Our numerical results also indicate that while LLMs provide valuable predictions, one cannot fully trust the prediction results solely from LLMs since experimentation-based approaches still remain the gold standard and can be leveraged to correct the errors in LLMs.
-From a managerial perspective, LOLA offers a versatile solution applicable to various scenarios where firms need to decide on content display, such as news articles, headlines, advertisements, or website content.
-Utilizing open-source LLM models, this method is cost-effective and can be easily deployed across different tasks once LLM models are fine-tuned on relevant datasets.
-The fine-tuning process is cost-efficient and can be performed on entry-level GPUs, making advanced LLM-based techniques accessible to firms with limited budgets.
-Additionally, we document that prompt-based approaches, which have been the standard method in recent marketing literature, perform poorly in this content attractiveness comparison task, underscoring the importance of the fine-tuning approach.
-Finally, we highlight that LOLA can be readily adapted to other measures beyond clicks.
-The click-based metric, while widely used by Upworthy a decade ago, has raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity.
-The industry is rapidly evolving, with Upworthy now focusing more on content quality and humanity, although clicks still remain an important measure; see more detailed discussion in (Upworthy2023).
-Our LOLA framework can easily accommodate these changes by adjusting the reward notion in the algorithm to align with new business goals, demonstrating its flexibility and relevance in a dynamic industry.
-Code and Data
-The code used in this study is publicly available at LLM News Github Repository to enable the reproduction of our results and to support future research that applies LLMs to experimentation and business problems.
-Funding and Competing Interests Declaration
-Author(s) have no competing interests to declare.
-Appendix A Supplementary Materials for Section 3
-A.1 Pairwise Comparison Between GPT Prompts
-To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on 1,000 correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other.
-The mean differences and p values are documented in Figure A1 and Figure A2.
-The random guess model simply predicts the first headline wins with a half probability.
-The nomenclature for GPT prompts in these two figures follows a specific format.
-The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are flipped.
-A value of 0 indicates the correct label, while a value of 1 indicates the wrong label.
-For example, GPT-4-0-0 means zero-shot prompting using GPT-4, and GPT-3.5-5-1 means in-context learning with 5 demonstrations and wrong labels using GPT-3.5.
-Figure A1: Average accuracy differences between different GPT prompts.
-The number in the heatmap presents the accuracy of prompts indicated on the left minus the accuracy of prompts indicated at the bottom.
-Shaded blocks indicate insignificant differences.
-Figure A2: The number in the heatmap presents the p-value from the paired t-test between different GPT prompts.
-A.2 Implementation Details of MLP Classification Models
-Hyperparameters of the network structure.
-We employ a fully connected neural network with ReLU activation functions for the hidden layers and a sigmoid activation function for the output layer.
-This configuration is widely used in practice.
-To determine the number of layers and the hidden nodes of each layer, we choose the number of layers within the range of $\{1,2,3,4,5\}$, and set the number of nodes in the first fully connected layer from the set $\{128,256,512,1024,2048\}$.
-Besides, we adjust the previous-post ratio, which is the proportion of nodes in each layer relative to those in the subsequent layer.
-This ratio here is in $\{2,4,8\}$.
-Note that the previous-post ratio becomes irrelevant to the results when there’s only one hidden layer.
-This is because the final output layer has only one dimension for ranking the most attractive headline.
-We incorporate the dropout after each hidden layer to mitigate the overfitting.
-Dropout layers randomly deactivate neurons based on a specified dropout rate, while scaling up the remaining neurons to maintain the same mean of the layer.
-In this work, we set the dropout rate in $\{0.3,0.5\}$.
-All hyperparameters are automatically selected through hyperparameter optimization, which is introduced below.
-Hyperparameters of the training procedure.
-The MLP is implemented in PyTorch, with data split into 70% for training, 10% for validation, and 20% for testing.
-To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch.
-Determining the appropriate batch size is crucial, as large batch sizes may lead to convergence to local optima, while small batch sizes may hinder convergence.
-We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$.
-Additionally, to mitigate overfitting, we apply $L_{2}$ regularization, with the regularization weight chosen from $\{1e-6,1e-4,1e-2\}$.
-We use binary cross-entropy as the loss function and employ stochastic gradient descent for training.
-We implement early stopping by monitoring the validation set loss, halting the training process if the reduction in loss over 60 epochs falls below a tolerance threshold of $10^{-4}$.
-Hyperparameter optimization.
-We optimize the parameters mentioned above through Bayesian Optimization (snoek2012practical).
-This Bayesian technique uses Bayesian inference to construct a probabilistic model of the objective function, where the parameters are updated iteratively with new observations.
-This iterative process balances exploration of the parameter space to discover promising regions and exploitation of known high-performing areas.
-In this work, we consider the accuracy of the validation set as the objective and optimize hyperparameters selected from the specified parameter space, as delineated by the ranges mentioned previously.
-We choose the Gaussian process as the estimator with the automatic optimizer, which is commonly adopted by the community.
-To ensure a judicious trade-off between efficacy and efficiency, we repeat the optimization process 50 times for input data of 256 dimensions and 30 times for data of 3,072 dimensions.
-A.3 Results of DistilBERT LoRA Fine-Tuning
-One difference between DistilBERT and Llama-3 is that DistilBERT is a bidirectional Transformer model, which means that its attention mechanism allows the model to learn the context from both directions, while Llama-3 is a unidirectional model.
+D.2 Additional LoRA Models
+D.2.1 Incorporating Insignificant Pairs for LoRA Fine-tuning
+Similar to the embedding-based logit model, we have experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting.
+However, as we observed, these modifications generally reduce performance on the training set without yielding significant improvements on the test set.
+In the following, we test whether incorporating more data, including insignificant pairs, can help improve the performance.
+Notice that by incorporating more significant pairs, we indeed observed improvement in test accuracy (see Figure A4), the observation of which holds for both embedding-based models and LoRA fine-tuned models.
+Here, we only present the results from fine-tuning Llama-3-8b using different sizes of training data, and the result from the embedding-based model remains the same.
+Recall that in §$\S$3.3.3, we use only significant pairs (with p-values smaller than or equal to 0.05, which includes 39,158 pairs) and split them into training and test sets.
+We continue to use the same significant test set for performance evaluation.
+However, for the training set, we obtain the data in the following manner: first, we remove the fixed test set from all pairs, leaving a total of 134,941 pairs.
+Then, given a ratio (e.g., 60%), we select these remaining pairs’ first 60% (ranked by ascending p-value) as the training set, see Figure A8 for a diagram of our dataset construction.
+Figure A8: Dataset Construction by Incorporating Both Significant and Insignificant Pairs in Llama-3-8b LoRA Fine-tuning
+We set the ratio to 20%, 40%, 60%, 80%, and 100%.
+For each ratio and size of the training set, we finetune the Llama-3-8b model and evaluate the performance on the fixed test set.
+Except for the training set, all other training configurations are the same as in §$\S$3.3.2.
+The performance is reported in Figure A9.
+The accuracy of the training set decreases as the ratio gets larger since we are incorporating more insignificant pairs.
+The accuracy on the test set maintains around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset.
+Notice that the decreasing trend of accuracy on the training set in Figure A4 is less steep because the pairs in that training set are all significant.
+However, the main goal here is to confirm that more insignificant pairs cannot improve the performance for either LoRA or the embedding method.
+Figure A9: Performance of fine-tuning Llama-3-8b under varying training dataset sizes.
+D.2.2 DistilBERT LoRA Fine-Tuning
+We now discuss One difference between DistilBERT and Llama-3 is that DistilBERT is a bidirectional Transformer model, which means that its attention mechanism allows the model to learn the context from both directions, while Llama-3 is a unidirectional model.
 An illustration of the structure of the DistilBERT model combined with a classification head is shown in Figure A10.
@@
 Due to the smaller model size, we are able to conduct full parameter updates during the fine-tuning process without using LoRA.
-We use the same learning rate scheduling, number of training epochs, and train-test set split as in the experiment detailed in Section 3.3.3.
-The accuracy of DistilBERT LoRA fine-tuning on the same test set is around 75%.
-More details can be found in the code repository.
-Appendix B Algorithms with Contextual Information
-For the completeness of the paper, we introduce two additional LOLA algorithms in this appendix, which leverage contextual information.
-For linear contextual bandits, we assume that at each round, we first observe the context vector $x_{t,k}\in\mathbb{R}^{d}$ for each arm $k\in[K]$.
-This context information can include LLM embedding vectors and user information, among other features.
-We denote by $r_{t,k}$ the random reward of arm $k$ at round $t$.
-After pulling one arm at each round, only the reward of that arm, $r_{t}$, is revealed.
+We use the same learning rate scheduling, number of training epochs, and train-test set split as in the experiment detailed in §$\S$3.3.2.
+The accuracy of DistilBERT LoRA fine-tuning on the same test dataset is around 75%.
+Thus, we find that the newer Llama-3-8b model offers better performance upon fine-tuning.
+We refer interested readers to our code repository for further details on the fine-tuning process.
+Appendix E Cost Comparison of Pure-LLM-Based Methods
+In Table A2, we summarize the monetary cost and effort required for each pure-LLM method used in this paper.
+Method Cost Runing time Effort
+Prompt \$21.24 (10,000 GPT prompts) 121.5 mins for 10,000 GPT prompts (82 prompt requests per min) Low
+Embedding \$0.18 (text-embedding-3-large) 26 mins for embedding all 77,245 headlines (3,000 embedding requests per min) Moderate
+LoRA \$3 (Llama-3-8b) 2 hours (Llama-3-8b) High
+Table A2: Summary of cost, run-time, and effort for different methods.
+Note that the reported run-time for Prompt-based (GPT-3.5 and GPT-4) and embedding-based methods refer to the time needed to obtain responses or embeddings from OpenAI.
+Note that the prompt and embedding response time can vary based on the API usage tiers, with our results specifically derived from Tier 1.
+For the prompt-based method utilizing OpenAI API, the cost was \$1.13 for 5,000 GPT-3.5 prompts and \$20.11 for 5,000 GPT-4 prompts, which is approximately 20 times the cost of GPT-3.5.
+The total run-time for the entire experiment was 121.5 minutes, involving the testing of 10,000 prompts.
+For the embedding-based methods, obtaining text embeddings via the OpenAI API is extremely time-efficient and more economical compared to prompt-based methods.
+When using the embedding API, data transmission involves smaller, fixed-size vectors instead of variable-length text, which reduces data overhead and accelerates processing speed.
+We spent approximately \$0.18 to obtain text embeddings for all 77,245 headlines, while the prompt experiment with only 10,000 prompts cost around \$20.232323At the time of this study (May 2024), GPT-4 (turbo version) prompt costs \$10 per 1M input tokens and \$30 per 1M output tokens.
+The embeddings used in our study, text-embedding-3-large, only cost \$0.13 per 1M input tokens and there is no output token cost associated with embeddings.
+Our entire dataset comprises 77,245 headlines, with an average of 19 tokens per headline, resulting in a total embedding cost of approximately \$0.18.
+Additionally, processing embeddings is much faster (the maximum Tier 1 rate limit—3,000 requests per minute reached) compared to prompt-based methods (around 82 pairs per minute).
+This efficiency can lead to lower operational costs and easier integration into real-time applications.
+For the LoRA fine-tuning method, the GPU requirements depend on factors such as model size, batch size, optimization algorithm, and floating-point precision.
+In our experiment, full-parameter fine-tuning without LoRA would require over 100 GB of GPU memory.
+However, LoRA fine-tuning only requires 30 GB of GPU memory.
+To illustrate the benefit of reduced GPU requirements, we compare two commonly used GPU models for LLM training.
+Without LoRA, the NVIDIA H100 GPU, which has nearly the largest GPU memory (80 GB), is insufficient for fine-tuning.
+In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient.
+The price of an NVIDIA H100 GPU is ten times higher than an NVIDIA V100 GPU as of May 2024.
+Thus, using LoRA significantly reduces costs, making fine-tuning more accessible to researchers and practitioners.
+In our experiment, each replicate involved approximately 2 hours of fine-tuning the Llama-3-8b model on an NVIDIA A100 GPU with 80 GB memory, costing about \$3 on the cloud computing provider AutoDL.
+We also note that after fine-tuning the model, several cost-efficient techniques are available for deploying it for inference, significantly reducing computational costs and hardware requirements.
+For instance, llama.cpp allows LLM models to be deployed on machines without a GPU, making implementation more affordable.
+Moreover, model compression techniques such as quantization, sparsification, and knowledge distillation can drastically cut inference computation costs while maintaining, or only slightly affecting, model performance.
+For a comprehensive survey on efficient LLM deployment and inference techniques, please refer to (zhou2024survey).
+In summary, we recommend using either embedding-based methods or fine-tuning methods, as they achieve similar performance (at least in our setting), superior to prompt-based methods, and are cost-effective.
+These methods offer different levels of transparency: prompt-based methods using OpenAI’s API are entirely black-box; embedding-based methods use OpenAI’s black-box embedding model combined with a user-controlled classification model; while fine-t tuning open-source models provides full transparency.
+Fine-tuning methods require more significant engineering effort to fine-tune and deploy an LLM than embedding-based methods.
+Appendix F LLM-based CTR Prediction Models
+We compare two approaches for CTR prediction: an OpenAI-embedding-based model and a LoRA fine-tuned Llama-3-8b model.
+In summary, we find the LoRA approach has better performance and we finally choose this model in the first step of our LOLA algorithm.
+The OpenAI-embedding-based CTR prediction model is similar to the one used for binary classification, as shown in Figure 3, with three key modifications: changing the loss function from log loss to MSE, using the text embedding of a single headline (3072 dimensions) instead of the concatenated embedding from headline pairs, and switching from binary output to continuous CTR prediction.
+To fine-tune the Llama-3-8b model for predicting CTR, we use a slightly different model structure compared to the one used for the pairwise comparison task shown in Figure 5.
+Specifically, as illustrated in Figure A11, we still use the Llama-3-8b model as the base LLM.
+The input is a single headline, and we apply a linear transformation to the output vector of the final token from the last Transformer layer.
+This transformation produces a single value, representing the predicted CTR.
+The model is trained to minimize the MSE between the real CTR and the predicted CTR.
+All other hyperparameters, such as batch size, number of training epochs, learning rate and its scheduling, and LoRA-related parameters, remain unchanged as in §$\S$3.3.
+Figure A11: Model architecture for CTR prediction based on the Llama-3-8b model.
+Figure A12 presents the numerical results, showing both the training loss (MSE) and accuracy throughout the training process.
+Accuracy is computed as follows: for each dataset (training or test), we use the fine-tuned model to predict the CTR for each headline.
+For each news article, if the headline with the highest predicted CTR matches the headline with the highest actual CTR, we consider this prediction correct.
+The accuracy is then calculated as the proportion of news articles for which the model’s predictions are correct.
+Figure A12: Performance of fine-tuning the Llama-3-8b model for CTR prediction.
+The accuracy of the OpenAI embedding-based method and the random guess method are also reported.
+For the random guess method, the winner headline is uniformly and randomly selected for each news article.
+As shown in Figure A12, the fine-tuned Llama-3-8b model significantly outperforms the OpenAI embedding-based method in terms of accuracy.
+Regarding the R-squared values, the embedding-based approach achieves 0.241, whereas the LoRA fine-tuned method achieves a higher R-squared value of 0.315.
+In contrast, for the pairwise comparison task, the fine-tuned Llama-3-8b model shows only a marginal improvement over the OpenAI embedding-based method.
+CTR prediction on all headlines is more challenging than pairwise comparison on significant pairs because the CTR output has higher uncertainty compared to a binary label.
+We believe this increased difficulty necessitates a more complex model, such as Llama-3-8b, which results in a larger performance gap between the fine-tuned Llama-3-8b and the OpenAI embedding-based method.
+Additionally, note that the accuracy on the test set is higher than on the training set.
+In our experiment, many news articles have headlines with very similar CTRs, making it more difficult to distinguish between them, which lowers the accuracy.
+This result may be due to having more such news articles in the training set and fewer in the test set.
+It is possible that with a different random split of the training and test sets, the accuracy of the training set could be higher than that of the test set.
+Appendix G Details of LOLA Variants and Extensions
+G.1 LOLA with Best Arm Identification
+In this section, we showcase how the LOLA framework can be used to solve the best arm identification problem.
+The best arm identification problem is a variant of the multi-armed bandit problem, where the goal is to identify the arm (or arms) with the highest reward.
+This problem is particularly useful in practice when the objective is to identify the best arm as quickly as possible rather than optimizing cumulative reward.
+Best arm identification has many variants, and our LOLA framework can be effectively integrated into them.
+We present the general steps of such a best-arm identification algorithm and the integration of LOLA in Algorithm 2.
+Algorithm 2 LLM-Assisted Best Arm Identification (LLM-BAI)
+1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
+2:Hyperparameter Fine-Tuning: Use another subset of data to choose the LLM’s equivalent auxiliary sample size, $n^{\text{aux}}$.
+3:Online Learning Phase: Initialize the number of arms $K$, the number of initial pulls $T_{k}\leftarrow n^{\text{aux}}$, LLM-based CTR prediction $\bar{\mu}_{k}\leftarrow\mathcal{M}(x_{k})$ for all arms $k\in[K]$, and the candidate set as $\mathcal{C}\leftarrow[K]$.
+4:while $|\mathcal{C}|>1$ do
+5:     Pull each arm in $\mathcal{C}$ once, denote the observed reward for arm $k$ as $r_{k}$.
+6:     Update the sample average CTRs $\bar{\mu}_{k}\leftarrow(\bar{\mu}_{k}T_{k}+r_{k})/(T_{k}+1)$, update $T_{k}\leftarrow T_{k}+1$ for all arms $k\in\mathcal{C}$.
+7:     Update confidence interval for each arm as $\mathrm{CI}_{k}\leftarrow[\bar{\mu}_{k}-C(T_{k}),\bar{\mu}_{k}+C(T_{k})]$, for all arms $k\in\mathcal{C}$, where $C(T_{k})$ is a decreasing and non-negative function of $T_{k}$.
+8:     Eliminate suboptimal arms from $\mathcal{C}$, i.e., eliminate arms from $\mathcal{C}$ if their upper confidence bound is less than the lower confidence bound of some other arms.
+9:end while
+10:Return: The set of good arms.
+We compare the performance of our proposed LLM-BAI to the standard benchmarks, including the standard A/B test, pure BAI algorithm, and pure LLM.
+Similar to the numerical setting under the regret minimization framework, we treat the reported CTR in the Upworthy dataset as the true expected reward of each headline.
+Each bandit problem instance corresponds to one news item, and each news headline is viewed as one arm.
+The return of each arm follows a Bernoulli distribution with the expectation given as the reported CTR.
+We observe that many news items in the Upworthy dataset have headlines with similar CTRs, making the best arm identification problem challenging.
+In real-world practice, the benefit of choosing the best arm is quite small (or non-existent) when the difference between the best arm and other arms is small (or insignificant).
+Therefore, we select the $(\mathrm{ST})^{2}$ algorithm from (mason2020finding) as the pure BAI algorithm, which is tailored to scenarios where the difference between the best arm and the suboptimal arms could be small.
+[^24] Note that we can also consider a similar setting in the regret minimization framework, where the planner’s goal is to ensure that the regret is below a certain threshold rather than specifying it in comparison to the best arm, as in Equation (1); see feng2024satisficing for details.
+Specifically, if we denote the real expected return of each arm as $\mu_{k}$ and the expected return of the best arm as $\mu^{*}=\max_{k\in[K]}\mu_{k}$, and given thresholds $\epsilon$ and $\gamma$, and failure rate $\delta$, this algorithm returns a set of arms containing all arms with expected returns better than $\mu^{*}-\epsilon$, and no arms with expected returns worse than $\mu^{*}-\epsilon-\gamma$ with high probability.
+In our case, we set $\epsilon=0$, $\gamma=0.005$, and $\delta=0.2$.
+It means, with at least $0.8$ probability, the algorithm returns a good set, which does not contain any bad arm with CTR lower than $\mu^{*}-0.5\%$.
+We make slight modifications to the original algorithm for numerical stability; for details, please see our code at LLM News Github Repository.
+In the implementation, we adopt the same data splitting rule as described in §$\S$4.3.
+Specifically, we used 70% data to fine-tune the Llama-3-8b CTR prediction model (see details in Web Appendix §$\S$ F), 10% data to fine-tune the $n^{\text{aux}}$ parameter, and tested the performance of the LLM-BAI algorithm on the remaining 20% test dataset.
+To fine-tune the $n^{\text{aux}}$ parameter, we run the LLM-BAI on the algorithm tuning dataset with different $n^{\text{aux}}$ values in $\{20,50,100,200,300,400,500,700,1000\}$, and choose the one with highest success rate in BAI.
+We finally choose $n^{\text{aux}}=300$ as the best parameter since it has the highest success rate and the number of pulls is not significantly more than others.
+For the natural benchmark pure LLM, we use the same CTR prediction model as LOLA.
+We also implement the plain BAI without using LLM-based predicted CTR as priors.
+For the last benchmark, the A/B test method, we allocate a fixed number of pulls to each headline, pulling all headlines an equal number of times.
+We then calculate the empirical CTR for each headline and select the one with the highest empirical CTR.
+To evaluate the performance of the A/B test method, we run experiments using several different numbers of pulls per headline, specifically $\{200,300,400,500,600\}$.
+Figure A13: Performance of pure LLM, LOLA, BAI, and A/B test on the test dataset.
+The X-axis represents the average number of pulls per test (news).
+The Y-axis represents the success rate.
+A test is considered successful if all selected headlines (both LOLA and BAI could return multiple good arms) have CTRs above $\mu^{*}-0.5\%$.
+The right figure zooms in on the area boxed by the gray rectangle in the left figure to highlight the performance improvement of our LOLA method.
+We report the performance of LLM-BAI and the standard benchmarks in Figure A13.
+Note that the success rate for pure LLM here is higher than the accuracy (46.55%) reported in Figure A12, as we use a more lenient definition of success.
+Our results indicate that our LOLA algorithm outperforms both the pure BAI method and the A/B test method in terms of success rate when using a similar number of pulls per test.
+Specifically, LOLA improves the success rate by an absolute 2.39% and saves 1.47% of pulls compared to the A/B test.
+Compared to the plain BAI method, LOLA improves the success rate by an absolute 2.94% and saves 1.19% of pulls.
+G.2 LOLA with Thompson Sampling
+We present a version of LOLA (LLM-TS) that combines LLM and Thompson Sampling and compare its numerical performance with standard Thompson Sampling and LLM-2UCBs.
+LLM-TS extends the standard TS algorithm by initializing an LLM-based prior.
+See Algorithm 3 for details.
+Similar to LLM-2UCBs, here also we need to fine-tune the hyperparameter, $n^{\text{a}ux}_{k}$ in this case, which is the LLM’s equivalent auxiliary sample size.
+As in the main analysis, we fine-tune this parameter on the fine-tuning dataset and choose from the set In our numerical experiment, we consider the $n^{\text{a}ux}_{k}\in\{600,800,1000,1200,1400\}$ and find that $n^{\text{a}ux}_{k}=1200$ generates the highest clicks in the algorithm tuning dataset (which we then use in all our numerical experiments).
+Notice that this number is quite close to 1000, which was the number used in LLM-2UCBs.
+There is no significant difference using either $1000$ or $1200$ in clicks on the tuning dataset.
+The Beta prior $(\alpha^{0}_{k},\beta^{0}_{k})$ implies a CTR mean of $\alpha^{0}_{k}/(\alpha^{0}_{k}+\beta^{0}_{k})$ and variance of $\alpha^{0}_{k}\beta^{0}_{k}/((\alpha^{0}_{k}+\beta^{0}_{k})^{2}(\alpha^{0}_{k}+\beta^{0}_{k}+1))$.
+We use the mean and variance from the training data to initialize the priors to $\alpha^{0}_{k}=1.38$ and $\beta^{0}_{k}=96.11$ for all $k\in[K]$ in our experiment.
+Algorithm 3 LLM-Assisted Thompson Sampling (LLM-TS)
+1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
+2:Hyperparameter Fine-Tuning: Use another subset of data to select the LLM’s equivalent auxiliary sample size, $n^{\text{a}ux}$ and the standard TS prior $(\alpha^{0},\beta^{0})$.
+3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$, LLM-based CTR prediciton $\bar{\mu}^{\text{a}ux}_{k}\leftarrow\mathcal{M}(x_{k})$, the parameters in Beta priors $(\alpha^{1}_{k},\beta^{1}_{k})\leftarrow(\alpha^{0},\beta^{0})+(n^{\text{a}ux}\bar{\mu}^{\text{a}ux}_{k},n^{\text{a}ux}(1-\bar{\mu}^{\text{a}ux}_{k}))$ for all arms $k\in[K]$.
+4:for $t=1$ to $T$ do
+5:     Sample $\theta_{k}^{t}\sim\text{beta}(\alpha^{t}_{k},\beta^{t}_{k})$ for all arms $k\in[K]$.
+6:     Play the arm $a_{t}=\arg\max_{k\in[K]}\theta_{k}^{t}$.
+7:     Observe the payoff $r_{a_{t}}$.
+8:     Update $(\alpha^{t+1}_{k},\beta^{t+1}_{k})\leftarrow(\alpha^{t}_{k},\beta^{t}_{k})+(r_{a_{t}},1-r_{a_{t}})$ if $a_{t}=k$; update $(\alpha^{t+1}_{k},\beta^{t+1}_{k})\leftarrow(\alpha^{t}_{k},\beta^{t}_{k})$ if $a_{t}\neq k$.
+9:end for
+We now present the results on the performance of both LLM-TS and the standard TS algorithm on the test dataset in Figure A14.
+For comparison, we also show the performance of LLM-2UCBs and the other benchmarks.
+We also document the pairwise comparison of different algorithms and report the relative percentage reward improvement and the p-value from the t-tests in Table A3.
+We find that UCB-based algorithms perform better than TS-based algorithms in both the LLM-assisted version and the standard version.
+This observation is not uncommon because it is widely known that TS suffers from over-exploration (min2020policy); in other words, UCB’s upper confidence shrinks faster, leading it to explore more effectively).
+As a result, while it is possible to implement a Bayesian bandit such as an LLM-Assisted Thompson Sampling within our LOLA framework, at least in our setting, UCB-based algorithms perform better.
+Figure A14: Average clicks per experiment per period under different time horizon multipliers.
+Note that the Y-axis captures the average clicks per test per period.
+For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
+The Y value is simply the average of this number $0.03$ over all tests.
+This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
+Parameter $\tau$ LLM-2UCBs vs. UCB LLM-2UCBs vs. Pure LLM LLM-2UCBs vs. E&C LLM-2UCBs vs. LLM-TS LLM-2UCBs vs. TS UCB vs. Pure LLM UCB vs. E&C UCB vs. LLM-TS
+50 $6.95$∗∗∗∗ $0.09$ $8.69$∗∗∗∗ $1.42$∗∗ $8.54$∗∗∗∗ $-6.41$∗∗∗∗ $1.62$ $-5.17$∗∗∗∗
+100 $6.38$∗∗∗∗ $1.03$∗∗ $7.72$∗∗∗∗ $1.94$∗∗∗∗ $8.15$∗∗∗∗ $-5.02$∗∗∗∗ $1.26$ $-4.17$∗∗∗∗
+200 $4.98$∗∗∗∗ $1.78$∗∗∗∗ $8.16$∗∗∗∗ $2.37$∗∗∗∗ $6.88$∗∗∗∗ $-3.05$∗∗∗∗ $3.03$∗∗ $-2.49$∗∗∗∗
+400 $3.96$∗∗∗∗ $3.02$∗∗∗∗ $5.76$∗∗∗∗ $2.10$∗∗∗∗ $5.45$∗∗∗∗ $-0.90$ $1.74$ $-1.79$∗∗∗∗
+600 $3.04$∗∗∗∗ $3.49$∗∗∗∗ $4.83$∗∗∗∗ $1.86$∗∗∗∗ $4.66$∗∗∗∗ $0.43$ $1.73$∗∗ $-1.15$∗∗∗
+800 $2.35$∗∗∗∗ $4.20$∗∗∗∗ $5.36$∗∗∗∗ $1.91$∗∗∗∗ $4.24$∗∗∗∗ $1.81$∗∗∗ $2.94$∗∗∗∗ $-0.44$
+1000 $2.23$∗∗∗∗ $5.02$∗∗∗∗ $5.08$∗∗∗∗ $2.00$∗∗∗∗ $4.07$∗∗∗∗ $2.73$∗∗∗∗ $2.79$∗∗∗∗ $-0.23$
+Parameter $\tau$ UCB vs. TS Pure LLM vs. E&C Pure LLM vs. LLM-TS Pure LLM vs. TS E&C vs. LLM-TS E&C vs. TS LLM-TS vs. TS
+50 $1.48$ $8.59$∗∗∗∗ $1.33$∗∗ $8.44$∗∗∗∗ $-6.68$∗∗∗∗ $-0.14$ $7.02$∗∗∗∗
+100 $1.67$∗∗∗ $6.61$∗∗∗∗ $0.90$∗∗ $7.04$∗∗∗∗ $-5.36$∗∗∗∗ $0.40$ $6.09$∗∗∗∗
+200 $1.80$∗∗∗∗ $6.27$∗∗∗∗ $0.58$ $5.01$∗∗∗∗ $-5.36$∗∗∗∗ $-1.19$ $4.41$∗∗∗∗
+400 $1.43$∗∗∗∗ $2.66$∗∗∗ $-0.89$∗∗ $2.36$∗∗∗∗ $-3.47$∗∗∗∗ $-0.30$ $3.28$∗∗∗∗
+600 $1.57$∗∗∗∗ $1.30$ $-1.58$∗∗∗∗ $1.13$ $-2.84$∗∗∗∗ $-0.16$ $2.76$∗∗∗∗
+800 $1.84$∗∗∗∗ $1.12$ $-2.20$∗∗∗∗ $0.04$ $-3.28$∗∗∗∗ $-1.07$ $2.29$∗∗∗∗
+1000 $1.80$∗∗∗∗ $0.05$ $-2.88$∗∗∗∗ $-0.90$ $-2.93$∗∗∗∗ $-0.96$ $2.03$∗∗∗∗
+Table A3: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001.
+For example, under the impression per headline equal to 50, the average click per test of LOLA and UCB is 0.053760 and 0.050267, respectively; so the relative improvement is calculated as ${0.053760}/{0.050267}-1=6.95\%$.
+Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
+G.3 LOLA with Contextual Information
+We now introduce two additional LOLA algorithms that leverage contextual (i.e., user-specific features) information.
+For linear contextual bandits, we assume that at each round, we observe the context vector $x_{t,k}\in\mathbb{R}^{d}$ for each arm $k\in[K]$.
+This context information can include LLM embedding vectors and user-specific features, among other features.
+The random reward of arm $k$ at round $t$ is denoted by $r_{t,k}$.
+After pulling one arm $a_{t}$ at each round, only the reward of that arm, $r_{a_{t}}$, is revealed.
 We make the linear realizability assumption that the expected reward is linear in the context vectors, which means $\mathbb{E}[r_{t,k}|x_{t,k}]=x_{t,k}^{\top}\theta^{*}$.
@@
 Additionally, we assume access to a historical dataset of size $M$, denoted by $\{(x_{t,k},r_{t})\}_{t=-M}^{-1}$.
-With these definitions, we are ready to present our proposed Algorithm 2, named LLM-Assisted Linear Contextual Bandit.
+With these definitions, we now present our proposed Algorithm 4, named LLM-Assisted Linear Contextual Bandit.
 The LLM-LinUCB algorithm extends LinUCB (chu2011contextual).
 Under the linear realizability assumption, this algorithm achieves asymptotically optimal regret, matching the lower bound up to logarithmic factors.
+Note that this algorithm does not have an auxiliary sample hyperparameter $n^{\text{a}ux}$ because the historical sample size $M$ can automatically adjust the confidence bound via Gram matrix $A$, assuming the same linear reward function during the online learning phase.
 Algorithm 4 LLM-Assisted Linear Contextual Bandit (LLM-LinUCB)
-1:LLM Training Phase: Calculate $A\leftarrow I_{d}+\sum_{t=-M}^{-1}x_{t,a_{t}}x_{t,a_{t}}^{\top}$ (The $d$-by-$d$ identity matrix), and $b\leftarrow\sum_{t=-M}^{-1}x_{t,a_{t}}r_{t}$
+1:Embedding-based Classifier Training Phase: Calculate $A\leftarrow I_{d}+\sum_{t=-M}^{-1}x_{t,a_{t}}x_{t,a_{t}}^{\top}$, where $I_{d}$ is the $d$-by-$d$ identity matrix to avoid singular $A$, and $b\leftarrow\sum_{t=-M}^{-1}x_{t,a_{t}}r_{t}$.
 2:Hyperparameter Fine Tune: $\alpha\in\mathbb{R}^{+}$ for controlling the upper bound.
@@
 5:     Get weight estimator: $\theta_{t}\leftarrow A^{-1}b$
-6:     Observe $K$ context vectors, $x_{t,1},x_{t,2},\dots,x_{t,K}\in\mathbb{R}^{d}$, generated through an LLM-based embedding model concatenated with user context information.
+6:     Observe $K$ context vectors, $x_{t,1},x_{t,2},\dots,x_{t,K}\in\mathbb{R}^{d}$, including text embeddings, user-specific feature.
 7:     for $k=1,2,\dots,K$ do
@@
 10:     Choose action $a_{t}=\arg\max_{k\in[K]}\hat{r}_{t,k}$.
-11:     Observe reward $r_{t}$.
+11:     Observe reward $r_{a_{t}}$.
 12:     $A\leftarrow A+x_{t,a_{t}}x_{t,a_{t}}^{\top}$.
-13:     $b\leftarrow b+x_{t,a_{t}}r_{t}$.
+13:     $b\leftarrow b+x_{t,a_{t}}r_{a_{t}}$.
 14:end for
-The algorithm proceeds as follows: During the LLM training phase, we prepare the matrix $A$ and vector $b$ to estimate the linear weight vector.
-We may need an LLM-based embedding model to extract embedding vectors from the original text information.
-Similar to LLM-2UCBs, we fine-tune the upper bound control parameter $\alpha$ using another fold of data and a bisection search to minimize regret on this fine-tuning dataset.
+The algorithm proceeds as follows: During the classifier training phase, leveraging historical data, we prepare the matrix $A$ and vector $b$ to estimate the linear weight vector.
+We need an LLM-based embedding model to extract embedding vectors from the original text information.
+The hyperparameter fine-tuning is similar to LLM-2UCBs but with only the upper bound control parameter $\alpha$.
+Specifically, one can use the bisection search algorithm to find the optimal $\alpha$ that maximizes the total reward on this fine-tuning dataset.
 The online learning phase remains the same as the standard LinUCB.
-The advantages of LLM-LinUCB are its transparency due to the linear assumption, ease of execution with only one hyperparameter, and theoretical soundness with optimal regret.
-However, it is restricted to linear forms.
-There are two issues with this assumption.
+The advantages of LLM-LinUCB are its transparency due to the linearity assumption and ease of execution with only one hyperparameter, as well as the theoretical soundness of its optimal regret results.
+However, there are two issues with the linearity assumption.
 First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over embedding-based classification.
-Second, once concatenated with user contexts, it is unclear whether the linear form is a good approximation, necessitating further numerical study.
-To alleviate the limitations of the linear assumption, we introduce another LLM-assisted contextual bandit algorithm (Algorithm 3), which extends the FAst Least-squares-regression-oracle CONtextual bandits (FALCON) algorithm proposed by (simchi2022bypassing).
-This type of contextual bandit algorithm generalizes the linear realizability assumption to a more flexible framework, assuming the true prediction model $f^{*}\in\mathcal{F}$.
+Second, once concatenated with user contexts, it is unclear whether the linear function is a good approximation of the rewards model (and this requires further numerical study).
+In addition, LLM-LinUCB cannot accommodate LoRA fine-tuned open-source LLM, which is a complex nonlinear reward model.
+To alleviate the limitations of the linear reward assumption and allow general LoRA fine-tuned open-source LLM with user-specific features, we introduce another general LLM-assisted contextual bandit algorithm (Algorithm 5), which extends the FAst Least-squares-regression-oracle CONtextual bandits (FALCON) algorithm proposed by (simchi2022bypassing).
+This contextual bandit algorithm generalizes the linear realizability assumption to a more flexible framework, assuming that the true prediction model $f^{*}\in\mathcal{F}$.
 The functional class $\mathcal{F}$ can be very general and is not limited to linear models.
@@
 $p_{t}(k)=\begin{cases}\frac{1}{K+\gamma_{m}(\hat{f}_{m}(x_{t,\hat{a}_{t}},\hat{a}_{t})-\hat{f}_{m}(x_{t,k},k))},&\forall k\neq\hat{a}_{t},\\ 1-\sum_{k\neq\hat{a}_{t}}p_{t}(k),&k=\hat{a}_{t}.\end{cases}$
-11:         Sample $a_{t}\sim p_{t}(\cdot)$ and observe reward $r_{t}$.
+11:         Sample $a_{t}\sim p_{t}(\cdot)$ and observe reward $r_{a_{t}}$.
 12:     end for
@@
 After getting an LLM trained on new data, then the key step in this algorithm is using the “inverse proportional to the gap” rule to balance exploration and exploitation.
-At each round, we first identify the empirically best arm $\hat{a}_{t}$ and then compute the gap in expected reward between this arm and the others.
-The sampling probability $p_{t}(k)$ is inversely proportional to this gap, meaning that a larger estimated reward $\hat{f}_{m}(x_{t,k},k)$ results in a smaller gap and a higher probability of being pulled.
-For a more detailed discussion of this algorithm, please refer to (simchi2022bypassing).
-Our purpose here is to adapt this algorithm in the LLM setting and discuss its pros and cons for completeness.
-Appendix C LLM-Assisted Thompson Sampling
-For completeness, we also present another version of LOLA that combines LLM and Thompson Sampling. We will update its numerical results in the next version of the manuscript.
-Algorithm 4 LLM-Assisted Thompson Sampling (LLM-TS)
-1:LLM Training Phase: Train a LLM-based prediction model $\mathcal{M}(x)$ for CTRs using historical data samples, where $x$ is the contextual information for arms.
-2:Hyperparameter Fine-Tuning: Use another subset of data to select the LLM’s equivalent auxiliary sample size, $n^{\text{a}ux}$ and the standard TS prior $(\alpha^{0},\beta^{0})$.
-3:Online Learning Phase: Initialize the number of periods $T$, number of arms $K$, LLM-based CTR prediciton $\bar{\mu}^{\text{a}ux}_{k}\leftarrow\mathcal{M}(x_{k})$, the parameters in Beta priors $(\alpha^{1}_{k},\beta^{1}_{k})\leftarrow(\alpha^{0},\beta^{0})+(n^{\text{a}ux}\bar{\mu}^{\text{a}ux}_{k},n^{\text{a}ux}(1-\bar{\mu}^{\text{a}ux}_{k}))$ for all arms $k\in[K]$.
-4:for $t=1$ to $T$ do
-5:     Sample $\theta_{k}^{t}\sim\text{beta}(\alpha^{t}_{k},\beta^{t}_{k})$ for all arms $k\in[K]$.
-6:     Play the arm $a_{t}=\arg\max_{k\in[K]}\theta_{k}^{t}$.
-7:     Observe the payoff $r_{a_{t}}$.
-8:     Update $(\alpha^{t+1}_{k},\beta^{t+1}_{k})\leftarrow(\alpha^{t}_{k},\beta^{t}_{k})+(r_{a_{t}},1-r_{a_{t}})$ if $a_{t}=k$; update $(\alpha^{t+1}_{k},\beta^{t+1}_{k})\leftarrow(\alpha^{t}_{k},\beta^{t}_{k})$ if $a_{t}\neq k$.
-9:end for
-Appendix D Survey Results
-To illustrate the difficulty of the headline selection task and compare it with LLM-based methods, we conducted a survey among students, asking respondents which headline was catchier.
-Specifically, we conducted the survey study among undergraduate students, acknowledging that these participants were less experienced in editing than Upworthy editors and writers.
-Despite this limitation, we aim to highlight the challenges of this task for humans.
-It is important to note that all candidate headlines were chosen by Upworthy editors, who were uncertain which one was catchier, prompting the need for further A/B testing.
-We designed two similar surveys using Qualtrics.
-The first survey randomly sampled 3,000 pairs from 39,158 headline pairs, differentiated by CTRs at a 0.05 significance level, along with 500 pairs showing no significant differences in CTRs.
-The second survey utilized the same 3,000 significant samples as the first.
-In both surveys, students were asked to select the catchier headline from each pair.
-The introductory text for the surveys was:
-“Please help us understand what makes a news headline appealing to readers.
-In this survey, you will be presented with 20 pairs of headlines.
-Please choose the one that captures your attention more effectively or feels more engaging in each pair, assuming both headlines relate to the same news content.
-There are no right or wrong answers; we are simply interested in your opinions.”
-After providing identification information, including the last four digits of their phone and student ID, respondents proceeded to a new page featuring 20 headline comparison questions with the prompt, “Please select the headline that you find more attractive,” with only one permissible response.
-Notably, the first survey included an additional response option: “Both headlines seem equally good.” At the survey’s conclusion, we asked, “How often do you read the news?” with options “Never”, “1-2 days per week”, “3-4 days per week”, “5-6 days per week”, “Everyday”.
-Detailed information and survey results are recorded in Table A1.
-Because some students did not finish the survey or skipped some comparisons, one can see there are 3,809 responses out of $202\times 20=4,040$.
-To calculate the accuracy metric, we only focused on the significant pairs, and in the first survey, we ruled out the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys.
-“Accuracy” measure here is consistent with the LLM tests.
-We viewed the headline with a higher CTR as the correct answer, and calculated the proportion of correct answers as accuracy.
-By doing so, one should expect 50% accuracy by random guess since we randomly swapped the order of the headlines to make sure the displayed order does not correlate with the correct answer.
-Detailed information and survey results are documented in Table A1.
-Some students did not complete the survey or skipped some comparisons, resulting in 3,809 responses out of a potential 4,040 ($202\times 20$).
-To calculate the accuracy metric, we focused solely on the significant pairs, and in the first survey, we excluded the uncertain answer “Both headlines seem equally good” to facilitate a fair comparison between the two surveys.
-Accuracy here aligns with the LLM tests.
-We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy.
-Accordingly, one might expect a 50% accuracy rate from random guessing, as we randomized the order of the headlines to ensure that the display order did not correlate with the correct answer.
-The survey results indicate that respondents’ ability to identify catchier headlines is comparable to random guessing, with no accuracies significantly deviating from 50% at the 0.05 significance level.
-Furthermore, accuracies did not vary significantly across different groups based on news reading frequency.
-We also incorporated two attention checks in the second survey.
-These findings suggest that the human ability to discern catchier headlines is limited, highlighting the intrinsic difficulty of the task.
-Survey 1 Survey 2
-conducted time 4/14 - 4/21 4/29 - 5/17
-# of respondents 202 339
-headline pair pool 3000 significant + 500 insignificant pairs 3000 significant pairs
-# of sampled pairs in survey 20 20
-# total received responses 3,809 4,571
-overall accuracy % 45.22 51.58
-Results based on different answer to “How often do you read the news?”
-Never proportion % 21.78 16.22
-accuracy % 45.76 51.62
-1-2 days per week proportion % 48.51 49.56
-accuracy % 45.36 52.54
-3-4 days per week proportion % 18.81 18.29
-accuracy % 46.83 51.39
-5-6 days per week proportion % 4.95 5.31
-accuracy % 39.05 50.15
-Everyday proportion % 5.94 8.26
-accuracy % 42.35 47.48
-Table A1: Survey Response Analysis.
-No accuracies in this table are significantly lower or higher than 50% at 0.05 significant level.
-8 respondents did not finish Survey 2 in the limited given time, so the summation of proportion from different subgroups is smaller than 100% in Survey 2.
+This rule is not new (foster2020beyond; simchi2022bypassing) and can effectively reduce a contextual bandit problem to a regression problem.
+At each round, it first identifies the empirically best arm $\hat{a}_{t}$ and then computes the gap in expected reward between this arm and the others.
+The sampling probability $p_{t}(k)$ is inversely proportional to this gap, meaning that a larger predicted reward $\hat{f}_{m}(x_{t,k},k)$ results in a higher probability of being pulled/exploited.
+The gradually increasing parameter $\gamma_{m}$ (as a sequential exploration and exploitation balancer) guarantees the algorithm to explore more in the beginning rounds and exploits more in later rounds.

### v3

@@
 Abstract
-In the rapidly evolving digital content landscape, media firms and news publishers require automated and efficient methods to enhance user engagement.
-This paper introduces the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery.
-Leveraging a large-scale dataset from Upworthy, which includes 17,681 headline A/B tests, we first investigate three pure-LLM approaches: prompt-based methods, embedding-based classification models, and fine-tuned open-source LLMs.
-We find that prompt-based approaches perform poorly, achieving no more than 65% accuracy in identifying the catchier headline.
-In contrast, both OpenAI-embedding-based classification models and fine-tuned Llama-3 with 8 billion parameters achieve an accuracy of around 82-84%.
-We then introduce LOLA, which combines the best pure-LLM approach with the Upper Confidence Bound algorithm to allocate traffic and maximize clicks adaptively.
+Modern media firms require automated and efficient methods to identify content that is most engaging and appealing to users.
+Leveraging a large-scale dataset from Upworthy (a news publisher), which includes 17,681 headline A/B tests, we first investigate the ability of three pure-LLM approaches to identify the catchiest headline: prompt-based methods, embedding-based methods, and fine-tuned open-source LLMs.
+Prompt-based approaches perform poorly, while both OpenAI-embedding-based models and the fine-tuned Llama-3-8B achieve marginally higher accuracy than random predictions.
+In sum, none of the pure-LLM-based methods can predict the best-performing headline with high accuracy.
+We then introduce the LLM-Assisted Online Learning Algorithm (LOLA), a novel framework that integrates Large Language Models (LLMs) with adaptive experimentation to optimize content delivery.
+LOLA combines the best pure-LLM approach with the Upper Confidence Bound algorithm to allocate traffic and maximize clicks adaptively.
 Our numerical experiments on Upworthy data show that LOLA outperforms the standard A/B test method (the current status quo at Upworthy), pure bandit algorithms, and pure-LLM approaches, particularly in scenarios with limited experimental traffic.
@@
 This approach could significantly reduce the costs associated with experimentation while simplifying the content delivery processes.
-To address this question, we leverage a large dataset of content experiments by the publishing firm Upworthy conducted in the 2013–2015 timeframe.
+To address this question, we employ a large dataset of content experiments by the publishing firm Upworthy conducted in the 2013–2015 timeframe.
 The dataset consists of 17,618 headline tests spanning 77,245 headlines, 277 million impressions, and 3.7 million clicks.
@@
 In the first part of the paper, we use this dataset to examine the extent to which pure-LLM-based approaches can accurately predict headline attractiveness.
-For this exercise, we focus on pairs of headlines (within a given A/B test) whose CTRs are significantly different (based on the observed data).
-This gives us a total of 39,158 pairs of headlines and allows us to recast the prediction task as a binary classification task, where a random guess gives 50% accuracy, and the perfect classifier reaches 100% accuracy.
-Thus, this serves as a preliminary testbed to evaluate whether pure-LLM-based approaches can predict which of the two headlines is catchier (or clickable).
+For this exercise, we test the different approaches’ accuracies, defined as the success rate of identifying the headline with the highest click-through rate (CTR) in a given A/B test.
 We consider the three most widely used LLM-based approaches for this analysis—(1) Prompt-based approaches, (2) LLM text-embedding approaches, and (3) Fine-tuning approaches.
@@
 In-context learning involves providing the LLM with a few demonstrations of similar tasks before generating responses for a specific task (min2022rethinking; xie2021explanation).
-We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches, with accuracy hovering around 50%.
-GPT-4 performs marginally better—it achieves an accuracy of 57.26% for zero-shot prompting and 60-65% for in-context learning.
+We find that GPT-3.5 performs as poorly as a random guess for both prompt-based approaches.
+GPT-4 performs marginally better—it achieves an accuracy of 37.85% for zero-shot prompting and 38-40% for in-context learning.
 Overall, this suggests that prompt-based approaches cannot aid/replace content experiments in any meaningful way.
@@
 Specifically, we transform headlines into embedding vectors using the most recent embedding model from OpenAI.
-We then concatenate two embeddings as inputs, along with the winner with higher CTR as the output, to train binary classification models, including logistic regression and multilayer perceptron (a type of neural network).
-We find that OpenAI’s embeddings combined with simple logistic regression can achieve around 82.50% accuracy, while more complex neural networks do not improve performance over the simple logistic model.
+We then use the embedding as the input, along with the CTR as the output, to train CTR prediction models, including linear regression and multilayer perceptron (a type of neural network).
+After getting predicted CTRs, we choose the headline with the highest predicted CTR as the winner.
+We find that OpenAI’s embeddings combined with simple linear regression can achieve around 46.28% accuracy, while more complex neural networks do not improve performance over the simple linear model.
 Finally, we fine-tune a state-of-the-art open-source LLM — Llama-3 with 8 billion parameters, using low-rank adaptation (LoRA) (hu2021lora).
-This approach performs the best, with 83.40% accuracy, which is a small improvement over the embedding-based approach.
-However, even the best-fine-tuned LLMs are unable to perfectly generate/predict which types of content are most appealing to users.
+This approach performs the best, with 46.86% accuracy, which is a small improvement over the embedding-based approach.
+However, even the best-fine-tuned LLMs are unable to perfectly predict which types of content are most appealing to users.
 As such, none of the LLM-based approaches can match the accuracy of experimentation-based approaches.
 Therefore, in the second part of the paper, we propose and evaluate a novel framework for content experiments that combines the strengths of LLM-based approaches with the advantages of online learning algorithms, termed LOLA (LLM-Assisted Online Learning Algorithm).
-Our key insight is to use predictions from an LLM model as priors in online algorithms to optimize experimentation.
-The approach involves two steps.
+Our main insight is to use predictions from an LLM model as priors in online algorithms to optimize experimentation.
+While the general idea that priors can impact regret and rewards has existed for some time (bubeck2013prior; russo2014learning), these papers typically assume that the quality and distribution of priors are known.
+However, a key challenge in our setting is that the LLM predictions cannot be naively treated as priors since we do not know their prediction error or theoretical properties.
+Therefore, we build on the recently proposed 2-UCBs algorithm by gur2022adaptive to overcome this challenge.
+The original 2-UCBs algorithm was designed for a setting where the experimenter has access to auxiliary data and knows both the size of this auxiliary sample ($n^{aux}$) and knowledge of the prediction error in this sample.
+Our innovation is to extend this algorithm to accommodate LLM predictions by treating these predictions as if they came from a “pseudo-sample” and then to fine-tune the size of this pseudo-sample for a given application based on prior data using hyper-parameter tuning.
+Thus, we are able to combine LLM predictions with online experimentation without making strong assumptions on the theoretical properties of the LLM-based predictors.
+LOLA involves two steps.
 First, we train an LLM-based CTR prediction model using LoRA fine-tuning.
@@
 We provide variants of the standard LOLA (and results on the empirical performance of these variants when applicable) for a few natural extensions, e.g., for Bayesian bandit settings where the second stage algorithm can be Thompson Sampling instead of UCB, for the best arm identification problem where the goal is to identify the best content/headline rather than regret minimization, and for settings where the firm has access to user/contextual features that can be used to personalize content delivery.
-In summary, our paper makes two key contributions to the literature.
-First, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings.
+In summary, our paper makes three key contributions to the literature.
+First, from an empirical perspective, the paper provides evidence that LLMs (or, more precisely, the current vintage of LLMs) cannot replace experimental approaches in the media and digital marketing industry.
+Second, from a methodological perspective, we develop a novel framework, LOLA, to leverage LLMs for improving content experimentation in digital settings.
 To the best of our knowledge, this is the first paper that proposes combining LLMs with adaptive experimentation techniques.
 We demonstrate the value of our approach using a large-scale dataset from Upworthy when compared to baseline methods.
-Second, from a managerial perspective, LOLA can be used in a broad range of settings where firms need to decide which content to show.
-While we focus on the news/publishing industry given our empirical context, the LOLA framework is general and can easily extend to other settings, including digital advertising, email marketing, and website design.
+Our work thus provides a foundation for future research into more advanced algorithms and offers insights for real-world validation in industry settings.
+Third, from a managerial perspective, LOLA can be used in a broad range of settings where firms need to decide which content to show.
+While we focus on the news/publishing industry, given our empirical context, the LOLA framework is general and can easily extend to other settings, including digital advertising, email marketing, and website design.
 Further, the method is cost-effective, can be built on open-source LLMs (alleviating data privacy concerns), and is easily deployable across different domains once fine-tuned on relevant datasets.
@@
 # of clicks in tests 3,741,517
-# of headline pairs 140,655
-# of headline pairs with different CTRs at 0.05 significant level 39,158
-# of headline pairs with different CTRs at 0.01 significant level 23,682
 # of impressions (both test and non-test impressions) 2,351,171,402
@@
 Table 3: Distribution of number of headlines tested, across A/B tests.
-To examine whether these A/B tests were informative, or not, we generate pairs of headlines and conduct pairwise t-tests on the CTRs of the headlines.
-For example, if four headlines were tested for one article, we generate $4\times 3/2=6$ pairs of headlines, and test whether each pair is significantly different from each other; [^3] The issue of multiple testing may arise when conducting a large number of pairwise comparisons, as it increases the likelihood of Type I errors, or false positives.
-In our dataset, tests involving 10 pairs or less (equivalent to no more than 5 packages) account for over 80% of all tests.
-The risk of multiple testing is relatively low when conducting fewer than 10 tests, and conventional significance levels, e.g., $\alpha=0.05$ are generally adequate to account for this risk.
-Ultimately, our objective is to develop a more effective decision rule for assigning experimental traffic, and in §$\S$4, we utilize all available data without concern for multiple testing.
-see Table 4 for examples of such tests for the two A/B tests highlighted earlier in Table 1.
-After generating pairs, we obtain 140,655 comparable headline pairs across all tests, of which 39,158 pairs significantly differ in CTRs at a 0.05 level and 23,682 at a 0.01 level.
-In summary, we find that only 28% of the tests were significant at the 0.05 level, and a mere 17% are significant at the 0.01 level.
-Overall, these results suggest that most pairs of headlines tend to perform similarly, and the task of learning which headline is catchier is non-trivial.
-To further illustrate the difficulty of this task, we conduct a survey, where we give human users pairs of headlines from the set of headline pairs that are significantly different from each other and ask them to identify the catchier headline.
+To illustrate that learning which headline is best is a non-trivial task, we conduct a survey where we give human users pairs of headlines from the set of headline pairs that are significantly different from each other and ask them to identify the catchier headline.
 Even within this set of headline pairs, we find that the respondents’ accuracy in the survey was not significantly better than random guessing.
@@
 See Web Appendix §$\S$§A for more details of the survey and its results.
-Test ID Headline 1 Headline 2 CTR 1 CTR 2 p value Winner
-1 How YOU Can Help… Why Yoko Ono Is… 0.0071 0.0124 0.054 2
-1 Why Yoko Ono Is… New York’s Last… 0.0124 0.0056 0.009 1
-1 How YOU Can Help… New York’s Last… 0.0071 0.0056 0.496 1
-2 If You Know Anyone Who… Hey Dude. If… 0.0289 0.010 0.000 1
-Table 4: Example of pairwise comparisons for two A/B tests from the Upworthy data.
+[^3] Note that just because an average reader cannot discern/identify the most engaging or catchiest headline among a pair/set of headlines, it does not follow that such differences do not exist.
+Indeed, the whole premise of the experimentation culture in digital firms relies on the idea that while managers/lay users cannot correctly predict which version of an ad creative/promotion/website layout will be more engaging, such differences exist and that identifying these differences and implementing the best-performing arm is important and will lead to better consumer engagement and business outcomes.
 3 Pure-LLM-Based Methods
-In this section, we consider the 39,158 headline pairs that significantly differ from each other at the 0.05 significance level.
-We focus on these significant pairs for three reasons: First, insignificant pairs suggest that decisions in final headline selection do not substantially affect users’ click behavior, implying that accuracy in predicting outcomes for “equally good” samples is not critically important from a business perspective.
-Second, the primary objective here is to evaluate various LLM-based classification methods to determine the most effective approach for the algorithm design in §$\S$4.
-As such, relative comparison across methods is more important as long as we use the same dataset for all methods.
-Third, as demonstrated in Web Appendix §$\S$§D.2.1, our robustness checks confirm that including insignificant pairs in the training phase does not affect the overall performance of the methods considered.
-Furthermore, we note that all methods discussed in this section can be readily adapted from binary to multi-class classification tasks.
-Indeed, in §$\S$4 where we introduce our LOLA approach, we incorporate practical elements such as multiple headlines and insignificant pairs to ensure our experimental setup closely mirrors real-world conditions and that our conclusions are directly applicable to business scenarios.
-In the rest of this section, we investigate the performance of three widely used LLM-based approaches—(1) GPT prompt-based approaches, including zero-shot prompting and in-context learning, (2) embedding-based classification models, and (3) fine-tuning of LLMs.
+We divided the 17,681 headline tests into three folds: 70% for training, 10% for hyperparameter fine-tuning of the algorithms (as discussed in §$\S$4), and 20% for testing.
+This sample split is used consistently throughout the remainder of the paper unless stated otherwise.
+Specifically, in this section, we have 12,376 headline tests as a training set and 3,263 headline tests as a test set.
+[^4] Note this test set is slightly smaller than 20% of headline tests.
+This is because, to avoid information leakage, we deleted the duplicated tests that had the same headlines in the training dataset.
+This happens because Upworthy allowed additional experiments for the same headline under a new experiment ID, and the editor may require an extra test to make an informed decision.
+In the rest of this section, we investigate the performance of three widely used LLM-based approaches—(1) GPT prompt-based approaches, including zero-shot prompting and in-context learning, (2) embedding-based models, and (3) fine-tuning of LLMs.
 3.1 Prompt-based Approaches
-Prompt-based approaches use GPT prompts to learn which headline is catchier, and there is no modeling or explicit fine-tuning involved.
+Prompt-based approaches use GPT prompts to select which headline is the catchiest, and there is no modeling or explicit fine-tuning involved.
 Prompt-based approaches have been the standard use-case for AI models in the marketing literature so far and have been used to examine if and when LLMs can simulate consumer behavior in market research studies (gui2023challenge; brand2023using; li2024frontiers).
@@
 3.1.1 Zero-shot Prompting
-Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly trained on specific examples.
-We evaluate the performance of headline pair comparison using zero-shot prompting based on OpenAI’s GPT models.
+Zero-shot prompting is a technique in which an LLM generates responses or performs tasks without being explicitly provided by any specific examples.
+We evaluate the performance using zero-shot prompting based on OpenAI’s GPT models.
 We do this through OpenAI’s API in Python, which offers various parameters to customize and control the model’s behavior and responses.
@@
 The user’s input or query is contained in the user parameter, while the assistant parameter holds the model’s previous responses to maintain context in an ongoing conversation.
-Additionally, we set the default temperature parameter of 1.444The temperature parameter scales and controls the randomness of the model’s responses.
+Additionally, we set the temperature parameter to be 0, which indicates there is no randomness in the GPT’s response for reproducibility.
+[^5] The temperature parameter scales and controls the randomness of the model’s responses.
 A larger value results in a more diverse and creative output, while a lower temperature makes the output more deterministic and focused.
-For simplicity, we use the default setting with temperature as 1.0.
-We also tested a lower temperature of 0.1 to generate more deterministic outcomes, but we did not observe any significant difference in accuracy.
+The default setting with temperature is 1.0.
+We also tested the default temperature, and we did not observe any significant difference in accuracy.
 Figure 1: Zero-Shot Prompting for Headline Selection.
-We design the zero-shot prompting to ask GPT which headline is catchier, as shown in Figure 1.
-We test this prompt using three OpenAI large language models, [^5] As recommended by OpenAI, see this page, accessed on May 24, 2024, one can default to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o.
+We design the zero-shot prompting to ask GPT which headline is the catchiest, as shown in Figure 1.
+We tell ChatGPT that it is an expert in marketing and specializing in optimizing headline selection.
+In this figure, we use letters (A,B,C,...) to mark different headlines.
+(brucks2023prompt) shows that different marker types and the order of options provided to LLM might affect the model’s response, in the sense that LLM might prefer to select the first option or select a certain marker.
+To mitigate this issue, we also consider other markers, numbers (1,2,3,...) and symbols (!,@,#,...), and we randomize the order of the options.
+Further, we adopt the common practice of LLM prompting for a specialized task by detailing the task for the LLM and clearly defining the meaning of a catchy headline.
+Notably, the first paragraph of the prompt assigns the role of a digital media expert to the LLM.
+This technique, commonly referred to as role prompting, involves explicitly assigning a specific role to the LLM to guide its responses.
+Role prompting has been shown to be an effective strategy in the broader survey of prompt engineering techniques (schulhoff2024prompt), and we find that it works well in our setting, too.
+We test this prompt using three OpenAI large language models, [^6] As of May 24, 2024, OpenAI recommends defaulting to GPT-3.5-turbo, GPT-4-turbo, or GPT-4o.
 OpenAI notes that GPT-4-turbo and GPT-4o offer similar levels of intelligence.
@@
 We design in-context learning prompts as shown in Figure 2.
-Our prompts include demonstrations of catchier headlines and ask the model to determine which headline is catchier for a subsequent pair of headlines.
-We use the headline with the higher CTR from the test, as indicated in the “Winner” column of Table 4, as the correct answer in our demonstrations.
+Our prompts include demonstrations of best-performing headlines and ask the model to evaluate the subsequent headlines.
+The demonstrations are selected from the training set, and for different prompt inquiries, we use the same demonstration to maintain comparability.
 A few demonstrations are typically sufficient to guide the model.
@@
 An interesting aspect of in-context learning is the model’s robustness to incorrect labels in the demonstrations.
-(min2022rethinking) show that randomized labels in demonstrations barely affect performance and highlight other critical aspects like the label space, input text distribution, and sequence format.
+(min2022rethinking) show that providing randomly selected labels as the answers in demonstrations barely affects performance and highlight other critical aspects like the label space, input text distribution, and sequence format.
 This is likely because the model relies more on the distribution of input text, label space, and format rather than specific input-label mappings.
 That is, in-context learning seems to benefit from recognizing general patterns and structures in the data, which the model infers from its pre-training rather than the actual labels.
-To examine if this is the case in our setting as well, we also consider experiments where we intentionally flip the labels in the demonstrations provided to the model.
+To examine if this is the case in our setting as well, we also consider experiments where we intentionally provide incorrect labels in the demonstrations.
 3.1.3 Performance Analysis of Prompt-based Approaches
-Before presenting the results, we make note of two points.
-First, we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity.
-Second, the GPT models can produce different answers even when given the exact same prompt.
-To mitigate this variability, we randomly select 1,000 significant pairs of headlines and conduct all our experiments on this same test set.
-Similarly, all the demonstrations are also randomly sampled and fixed for all experiments to reduce the variance of prompt outputs.
+Before presenting the results, we first make a note that we do not find any significant difference between GPT-4 and GPT-4o; therefore, we only report the results from GPT-3.5 and GPT-4 for simplicity.
 The results of all the prompt-based experiments are shown in Table 4.
-A more detailed pairwise comparison and t-test analysis of all the different experiments are available in Web Appendix §$\S$§B.1.
+A more detailed pairwise comparison and t-test analysis of all the different prompts are available in Web Appendix §$\S$§B.1.
 The second column of Table 4, $n_{demo}$, represents the number of demonstrations used for in-context learning, with $n_{demo}=0$ indicating a zero-shot prompting.
-Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with flipped outcome labels.
-The third column, “Flipped Label” indicates whether the labels were flipped or not.
-GPT Model $n_{demo}$ Flipped Label Accuracy [0.025 0.975]
-GPT-3.5 0 0 54.20% 51.10% 57.10%
-GPT-3.5 2 0 48.70% 45.50% 51.50%
-GPT-3.5 2 1 50.85% 47.60% 53.70%
-GPT-3.5 5 0 51.30% 48.20% 54.30%
-GPT-3.5 5 1 50.10% 46.80% 53.20%
-GPT-4 0 0 57.26% 54.20% 60.00%
-GPT-4 2 0 63.60% 60.50% 66.40%
-GPT-4 2 1 62.90% 59.70% 65.70%
-GPT-4 5 0 60.80% 57.70% 63.40%
-GPT-4 5 1 61.80% 58.50% 64.80%
-Table 5: Results for prompt-based approaches with different GPT models, the number of demonstrations, and whether the demo’s label is flipped or not.
-Accuracy is defined as the ratio of correct answers divided by the number of prompts (1000 in our analysis).
-The 95% confidence interval is calculated by performing 5,000 bootstrap resamples of the entire 1,000 pairs.
-To save on the monetary cost of prompts and not introduce additional variance, we did not rerun the GPT prompts for the same samples; instead, we used the responses/predictions from the original run.
-We find that prompt-based approaches using GPT-3.5, except for the zero-shot prompting, are not significantly better than random predictions (accuracy rates hover around 50%).
-Neither the number of demonstrations used nor the label type have any significant impact on the accuracy of the GPT-3.5 predictions.
-Interestingly, zero-shot prompting using GPT-3.5 is slightly better than random prediction and also marginally outperforms in-context learning.
-In comparison, experiments with GPT-4 have accuracy rates around 60-65%, which is a significant improvement in performance compared to GPT-3.5.
-For GPT-4, in-context learning significantly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance.
-There is also no significant difference between using correct labels and incorrect labels in the demonstrations for in-context learning.
+Additionally, for all the in-context learning experiments, we consider two scenarios—one where the demonstrations are shown with ground truth labels and another where they are shown with incorrect labels.
+The third column, “Incorrect Label” indicates whether the labels were correct (“Incorrect Label” = 0) or incorrect (“Incorrect Label” = 1).
+Model $n_{demo}$ Incorrect Label Accuracy on the test set
+Random Guess - - 33.02%
+GPT-3.5 0 - 29.76%
+GPT-3.5 2 0 34.23%
+GPT-3.5 2 1 31.60%
+GPT-3.5 5 0 30.95%
+GPT-3.5 5 1 30.59%
+GPT-4 0 - 37.85%
+GPT-4 2 0 39.96%
+GPT-4 2 1 38.77%
+GPT-4 5 0 39.17%
+GPT-4 5 1 38.28%
+Table 4: Results for prompt-based approaches with different GPT models, the number of demonstrations, and whether we intentionally provide incorrect solutions in the demonstration or not.
+Accuracy is defined as the ratio of correct answers over the test set size.
+The best performance on all test data is boldfaced.
+We find that prompt-based approaches using GPT-3.5, are not significantly better than random guess or even slightly worse.
+In comparison, experiments with GPT-4 have accuracy rates around 38-40%, which is a significant improvement in performance compared to GPT-3.5, and all settings for GPT-4 outperform the random guess.
+For GPT-4, in-context learning with two demonstrations slightly improves performance over zero-shot prompting in GPT-4, though increasing the number of demonstrations from two to five does not result in a significant difference in performance.
+The difference between using correct labels and incorrect labels in the demonstrations is also not significant for in-context learning.
 This confirms that in-context learning helps GPT understand the task better through the input-output examples, but it fails to leverage the information in mapping from inputs to outputs effectively.
-In summary, we find that prompt-based methods have low accuracy (no more than 65%), high randomness (different answer outputs even under the same prompt inputs), slow running time (compared to other methods introduced later in this section), and high monetary cost (see Web Appendix §$\S$§E for a detailed discussion of costs and running time).
+In summary, we find that prompt-based methods have low accuracy (no more than 40%), slow running time (compared to other methods introduced later in this section), and high monetary cost (see Web Appendix §$\S$§E for a detailed discussion of costs and running time).
 All of these drawbacks make them unsuitable for real-time business applications, i.e., these methods cannot replace experimentation for content selection in digital platforms.
-3.2 Classification Models with LLM-based Text Embeddings
-We now consider another LLM-based method for the prediction task.
-Specifically, we leverage the text embeddings from OpenAI and utilize classification models, including Logistic Regression (Logit) and Multilayer Perceptron (MLP), to predict which of the headlines in a pair is likely to generate a higher CTR.
+3.2 CTR Prediction Models with LLM-based Text Embeddings
+We now consider another LLM-based method for the headline selection task.
+Specifically, we leverage the text embeddings from OpenAI and utilize regression models, including Linear Regression (Linear) and Multilayer Perceptron (MLP), to predict the CTRs of headlines for the same article and then select the one with the highest predicted CTR as the winning headline.
 Text embedding techniques transform large chunks of text, such as sentences, paragraphs, or documents, into numerical vectors.
 These embedding vectors can then be used in a variety of applications, including text classification, where texts are grouped into categories; semantic search, which improves search accuracy by understanding the meaning behind the search queries; and sentiment analysis, which is used to determine the emotional tone of a piece of text.
-[^7] Embedding models and standard prompt-based LLMs used share some common features, e.g., the underlying Transformer architectures and extensive pretraining on large text corpora.
+[^8] Embedding models and standard prompt-based LLMs share some common features, e.g., the underlying Transformer architectures and extensive pretraining on large text corpora.
 However, they are optimized for different tasks.
@@
 See patil2023survey for a more detailed survey of the most recent text embedding models and their applications.
-An interesting question is why text embeddings obtained from LLM models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for downstream classification tasks.
-(saunshi2020mathematical) provide an explanation by demonstrating that classification tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text.
+An interesting question is why text embeddings obtained from LLM models trained for the next word/token prediction task with the cross-entropy loss are able to produce informative features for downstream tasks.
+(saunshi2020mathematical) provide an explanation by demonstrating that typical downstream tasks can be reformulated as sentence completion tasks, which can be further solved linearly using the conditional distribution over words following an input text.
 Thus, the embeddings, which can be roughly viewed as the next token probability generated from a next-word prediction task, inherently capture the contextual information and relationships between words, making them useful for various downstream tasks such as sentiment analysis and headline selection (as our experiments illustrate).
-3.2.1 Implementation Details
-Figure 3 illustrates the conceptual framework of our approach here.
-We concatenate the embeddings of two headline embeddings888 Because embeddings approximate the next token probability, one can also use the difference (subtraction) of two embedding vectors instead of the concatenation as the input to the downstream classification model.
-In our task, these two input types have no significantly different impact on the final classification accuracy.
-and apply the downstream classification models to predict the probability that the first headline will generate a higher CTR.
-We consider three embedding models:
-  • $\bullet$
-    OpenAI embeddings: At the time of this study, the latest and best embedding model available from OpenAI was text-embedding-3-large.
-These embeddings are obtained using the OpenAI embedding API in Python.
-This embedding model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021.
-A notable feature of this model is its flexibility – it allows users to tailor the embedding dimension to their specific needs.
-[^9] Other models capable of producing text embeddings, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained.
-These dimensions cannot be changed or resized for different applications.
-However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka).
-This allows a single embedding to adapt to the computational constraints of downstream tasks.
-More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used.
-In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream classification models.
-[^10] A higher dimension does not necessarily mean better performance.
-While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples.
-This sparsity can increase the possibility of over-fitting.
-  • $\bullet$
-    Llama-3-8b: Meta’s Llama-3 (metallama3) is regarded as one of the most advanced open-source LLMs to date (llama3modelcard).
-We obtain embeddings from Llama-3 with 8 billion parameters (Llama-3-8b).
-We fine-tune this exact model using LoRA in §$\S$3.3; therefore, using the embeddings from this model allows for a direct comparison of fine-tuning vs.
-embedding-based classification methods (see §$\S$3.4 for comparisons).
-Llama-3-8b does not produce embeddings explicitly designed to represent the semantic meaning of entire input sequences.
-To address this issue, we apply the technique described in (behnamghader2024llm2vec), which transforms the Llama-3-8b model into a more suitable text embedding model, outputting embeddings with a dimension of 4,096.
-  • $\bullet$
-    Word2Vec embeddings: Word2Vec, proposed by mikolov_etal_2013a, was the leading embedding model before the widespread adoption of Transformers in the Natural Language Processing (NLP) literature.
-Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information.
-Embeddings from Word2Vec have been used in the recent marketing literature in a variety of use cases, including product attribute modeling (timoshenko_hauser_2019; wang_etal_2022), the effect of creativity on product popularity (sozuer_etal_2024).
-We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec.
-To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
-Next, to assess the impact of model complexity on performance, we consider two classification models:
-  • $\bullet$
-    Logit model with linear features: This is one of the simplest binary classification models and predicts the probability that a given input belongs to a particular class by applying the logistic function to a linear combination of the input variables.
-  • $\bullet$
-    Multilayer perceptron: MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data.
-It employs a fully connected neural network with ReLU activation for hidden layers and Sigmoid transformation for the output probability, with dropout layers added to prevent over-fitting.
-We refer interested readers to Chapter 5 in (zhang2021dive) for more details.
-We train the MLP using the stochastic gradient descent algorithm to minimize binary cross-entropy and perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration.
-Please see Web Appendix §$\S$§C.1 for additional details of the fine-tuning process and the architectures used in the final implementation.
-Figure 3: The pipeline of the classification with the LLM text embeddings. Headline 1 and headline 2 are natural language sentences while Embedding 1 and Embedding 2 are numerical vectors.
-Finally, in terms of the data, as discussed earlier, we choose all headline pairs that are significant at the 0.05 level.
-Note that the same headline can appear in multiple pairs when there are more than two headlines tested in one experiment.
-To avoid information leakage, we first randomly split the data into 80% training and 20% test sets based on the test ID (the experiment ID for Upworthy) to ensure all headline pairs under the same experiment are either in the training set or the test set.
-Notice that Upworthy allowed additional experiments for the same headline under a new experiment ID because the editor may require an extra test to make an informed decision.
-It means we may have the same headlines in both the training and test sets, even if they were in different experiments.
-In such cases, we discard headlines that have appeared in the training set from the test set to avoid information leakage.
-[^11] There are alternative ways to do the train-test split without discarding any samples.
-One approach is assigning one headline pair at one time to a training or test set and then adding all pairs with the same headline in the same group until all samples are allocated.
-However, after conducting several experiments in our study, we found that the splitting procedure has minimal impact on the final results.
-Hence, to maintain clarity and simplicity in our exposition, we simply discard around 25% of test headlines that have appeared in the training set.
+Figure 3 illustrates the conceptual framework of our approach.
+We first use the LLM-based embedding model to transform the original headlines into embeddings.
+Then, we use the embeddings as inputs and CTRs as outputs to train the downstream CTR prediction model.
+For the test set, we leverage the trained regression model to predict the CTR of one headline at a time.
+After getting all CTR predictions of headlines in an A/B test, we rank the headlines with the highest CTR as the winning headline.
+We consider four embedding models, including the OpenAI embedding model, Llama-3-8b, Word2Vec, and BERT.
+Note that we do not train the embedding model and only train the downstream CTR prediction model using the training dataset, containing 12,376 headline A/B tests.
+A detailed introduction to embedding models and the implementation specifics are provided in Web Appendix §$\S$§C.1.
+Figure 3: The pipeline of the headline selection using LLM text embeddings.
+We use an A/B test with three headlines for illustration.
+Headlines 1, 2, and 3 are natural language sentences, while Embedding 1, 2, and 3 are numerical vectors.
 3.2.1 Performance Analysis of Embedding-based Approaches
-In Table 6, we present the performance of various embedding-based classification models.
-Embedding Model Classification Model Accuracy
-OpenAI-256E Logit 77.31%
-MLP 77.62%
-OpenAI-3072E Logit 82.50%
-MLP 82.26%
-Llama-3-8b-4096E Logit 74.03%
-MLP 79.40%
-Word2Vec-256E Logit 67.89%
-MLP 67.32%
-Word2Vec-3072E Logit 67.55%
-MLP 67.02%
-Table 6: The performance of different embedding and classification models on test data.
+Embedding Model CTR Prediction Accuracy on the test data
+OpenAI-256E Linear 46.28%
+MLP 44.38%
+OpenAI-3072E Linear 43.06%
+MLP 45.17%
+Table 5: The performance of OpenAI embeddings and CTR prediction models on the test data.
 All models use the same training data, and accuracy is evaluated on the test data.
-Note that the Logit model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents.
-Therefore, we rerun MLP training three times with different sample paths (training set shuffling) and report the average accuracy from three reruns.
-At a high level, we see that all the embedding-based classification models perform better than prompt-based approaches.
+Note that the Linear model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents.
+Therefore, we rerun MLP training three times with different random seeds and report the average accuracy from three reruns.
+In Table 5, we present the performance of CTR prediction models based on OpenAI’s embeddings, including the 256-dimensional embedding (OpenAI-256E) and the 3072-dimensional embedding (OpenAI-3072E, which is the maximum size).
+The full performance results, including Llama-3-8b, Word2Vec, and BERT, can be found in Web Appendix §$\S$§C.2.
+At a high level, we see that all the embedding-based regression models obtain better accuracy than prompt-based approaches.
 This performance can be attributed to the fact that these models are trained on a customized dataset, allowing them to learn decision rules tailored to the headline selection task.
@@
 In contrast, in-context learning relies on the LLM’s ability to understand and generalize from only a few demonstrations.
-Furthermore, the task of choosing the catchier headline from two options may not be a common scenario in OpenAI’s training corpus.
+Furthermore, the task of choosing the best headline from multiple candidates may not be a common scenario in OpenAI’s training corpus.
 Consequently, GPT models may struggle to generalize effectively from the limited examples provided in prompts, leading to less accurate predictions.
-Next, we compare performance under different embedding models.
-We see that OpenAI embeddings outperform both Llama-3-8b and Word2Vec-based models.
-This is understandable since OpenAI’s LLMs are state-of-the-art models that naturally perform better than the classic Word2Vec embeddings.
-Also, the performance of embeddings from open-source Llama-3-8b falls between that of the proprietary OpenAI model and the older Word2Vec that does not use the transformer architecture.
-We also see that for OpenAI, higher-dimensional embeddings give higher-quality encoding and, thus, higher prediction accuracy.
-However, there is some limit beyond which there are diminishing returns.
-In our experiment, the 3072-dimensional embedding (OpenAI-3072E) improves over the 256-dimensional embedding (OpenAI-256E) by around 5%.
-A more striking observation is that classification model complexity does not improve the prediction for OpenAI embeddings.
-Note that there is no significant difference in accuracy between MLP and Logit for the following embedding models – OpenAI-256E, OpenAI-3072E, Word2Vec-256E, and Word2Vec-3072E.
-But for Llama-3-8b embeddings, the MLP classification model outperforms Logit classification model by approximately 5%.
-We hypothesize that with more dimensions in text embeddings, the separation between different classes becomes more complex (e.g., higher sparsity and higher risk of over-fitting), requiring a more sophisticated model for classification.
-Overall, our findings suggest that a Logit classifier with linear features on top of fixed text embeddings from OpenAI can effectively capture the attractiveness of content.
-We now investigate a few other aspects of these models and summarize our findings below.
-For brevity, we simply discuss the findings here and refer interested readers to the Web Appendix for details.
-Note that in all these analyses, we focus on OpenAI embeddings (since it offers the best performance) and the logistic regression with linear features as our classification model (since the MLP does not offer any significant improvement over the logit model for OpenAI embeddings and is easier to train).
+We also find that, for OpenAI, higher-dimensional embeddings do not lead to higher prediction accuracy.
+Given that the 256-dimensional embedding stores only essential information, while the 3072-dimensional embedding also captures finer details, this performance gap is possibly due to the overfitting with the limited size of the training set.
+Predictions using higher-dimensional embeddings may cause the prediction model to focus too much on details and overfit the training set.
+A more striking observation is that increasing the complexity of the CTR prediction model does not necessarily improve predictions for OpenAI embeddings.
+We hypothesize that with more dimensions in text embeddings, the CTR prediction task requires a more sophisticated model due to the higher risk of overfitting.
+Overall, our findings suggest that a Linear CTR prediction model on top of fixed text embeddings from OpenAI can effectively capture the attractiveness of content.
+We now investigate three other aspects of these models and summarize our findings below. For brevity, we simply discuss the findings here and refer interested readers to the Web Appendix for details.
   • $\bullet$
-    First, we examine the conditions under which these approaches fail.
-We use cosine similarity measures to quantify the similarity between the two headlines in a pair and find that these approaches tend to perform worse when the headlines are similar.
-This is understandable since text embeddings may not be able to capture the differences between two headlines when they are similar.
-As a result, it is hard for a classification model to differentiate between them.
-See Web Appendix §$\S$§C.2 for details.
+    We compare performance under different embedding models. We observe that OpenAI embeddings outperform Llama-3-8b, Word2Vec, and BERT-based models. See Web Appendix §$\S$§C.2 for details.
   • $\bullet$
-    Second, we quantify the role of the number of training samples on predictive accuracy.
-For the 256E-dimensional embedding, we find that approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance.
-However, when we use the 3072E-dimensional embedding, there is evidence of overfitting even when we use 95% of the data for training.
-This suggests that for larger embedding vectors, we need much larger training datasets.
-See Web Appendix §$\S$§C.3 for details.
+    We quantify the impact of training sample size on predictive accuracy.
+In Web Appendix §$\S$§C.3, we train the model using varying proportions of the training data and evaluate its performance.
+We find that larger training datasets lead to higher accuracy rates, although the improvement is gradual.
+For instance, increasing the training size from 20% to 70% of the total dataset raises accuracy from 44% to 46% on the test data.
   • $\bullet$
-    Third, we examine whether we can improve the performance and alleviate any overfitting issues by adopting dimensionality reduction techniques (PCA) and/or regularization (adding $L^{1}$/$L^{2}$ norms).
-However, we don’t see any significant improvement with these methods.
-See Web Appendix §$\S$§C.4 for details.
-  • $\bullet$
-    Finally, there are two potential threats to the inference from our analysis.
+    There are two potential threats to the inference from our analysis.
 First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate the results.
@@
 3.3 Fine-Tuning Open-Source LLMs with LoRA
-In §$\S$3.2, we trained classification models using text embeddings as fixed inputs, bypassing the need to train the LLM itself.
-The best-performing embedding-based model reaches an accuracy of 82.50% without modifying any parameters in the LLM.
+In §$\S$3.2, we trained prediction models using text embeddings as fixed inputs, bypassing the need to train the LLM itself.
+The best-performing embedding-based model reaches an accuracy of 46.28% on the test set without modifying any parameters in the LLM.
 This raises a natural question: if we can train LLMs directly, can we achieve better performance?
@@
 Note that to fine-tune LLMs, we need to be able to access the underlying model parameters.
-This is only feasible for open-source LLMs, like BERT and Llama.
+This is only feasible for open-source LLMs like BERT and Llama.
 [^9] Some proprietary LLMs offer black-box fine-tuning.
-For instance, OpenAI allows fine-tuning GPT-3.5, GPT-4/GPT-4o (with limited experimental access) through its API in a black-box manner.
-Users provide the training and test datasets and select the base GPT model, but are not provided any details of the fine-tuning process or the underlying model parameter.
+For instance, OpenAI allows fine-tuning GPT models through its API in a black-box manner.
+Users provide the training and test datasets and select the base GPT model, but are not provided any details of the fine-tuning process or the underlying model parameters.
 This approach also has a number of other drawbacks – the fine-tuning API is quite expensive compared to the open-source fine-tuning approaches, the fine-tuned model cannot be hosted on a local machine, data privacy concerns since the media firm’s proprietary data will need to be submitted to OpenAI.
 Given these drawbacks, it is preferable for firms to fine-tune open-source LLMs, as the process and data are fully under the control of the firm.
-However, for the completeness of comparison, we fine-tuned the ChatGPT-3.5 model on the training set using this API, with the default configuration provided by OpenAI, and the same training-test dataset split.
-We obtained an accuracy of only 79.80% on the test set.
+However, for the completeness of comparison, we fine-tuned the GPT-4o (gpt-4o-2024-08-06 version) model on the training set using this API, with the default configuration provided by OpenAI and the same training and test datasets.
+Input to the model is several headlines of one article, and the target output is the index of the catchiest headline.
+We obtained an accuracy of 48.82% on the test set, but this result comes with a high cost of \$71.77 for the test set alone.
 The goal of fine-tuning is to adapt a general-purpose LLM for specific downstream tasks through a process typically involving supervised learning using labeled datasets.
@@
 Llama-3 offers two versions: one with 8 billion parameters (Llama-3-8b) and another with 70 billion parameters (Llama-3-70b).
-We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see §$\S$3.3.3 for details).
+We opted for the smaller version since it is easier to fine-tune and, as we will see, even with this smaller model, we see evidence of over-fitting (see §$\S$3.3.1 for details).
 Hence, we avoid fine-tuning the 70-billion-parameter model, where over-fitting problems are likely to be exacerbated.
-The rest of this section is organized as follows.
-We present the LoRA method for efficient fine-tuning in §$\S$3.3.1.
-The goal is not to delve into every technical detail but to provide the necessary background information for a high-level understanding of fine-tuning.
-Then we detail our implementation in §$\S$3.3.2 and report our results in §$\S$3.3.3.
-Background information related to Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 are given in Web Appendix §$\S$§D.1.
-3.3.1 LoRA Fine-Tuning Technique
+Background information related to the Transformer block, the fundamental component of most LLMs, including Llama-3, and the basic architecture of Llama-3 is given in Web Appendix §$\S$§D.1.
 Traditional fine-tuning of LLMs involves updating a substantial number of parameters, which can be computationally expensive and memory-intensive.
@@
 This approach leverages the low-rank decomposition, which significantly reduces both memory usage and computational time through a significantly reduced number of trainable parameters.
-We illustrate LoRA in Figure 4.
-We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$.
-Specifically, $W_{Q}$ is the weight matrix for the query vectors, $W_{K}$ is for the key vectors, and $W_{V}$ is for the value vectors in the attention mechanism of the Transformer architecture.
-These matrices store most of the information and knowledge learned by LLM, and are crucial for calculating the attention scores and subsequently determining the relevance of different tokens in the input sequence.
-For more technical background, we refer to Web Appendix §$\S$§D.1.
-We use $d\times d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices.
-We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
-$W^{\prime}=W+\Delta W,\quad\mathrm{with}\quad\Delta W=AB,\quad\mathrm{where}\quad A\in\mathbb{R}^{d\times r},\quad B\in\mathbb{R}^{r\times d_{W}}.$
-Figure 4: Illustration of LoRA Fine-Tuning.
-LoRA essentially introduces the low-rank matrices $A$ and $B$ to approximate the weight matrix change $\Delta W$ after the fine-tuning.
-These low-rank matrices are updated during the fine-tuning process, while the original weight matrix $W$ remains frozen.
-This strategy significantly reduces the number of trainable parameters from $d\times d_{W}$ to $r\times(d+d_{W})$ because the rank $r$ is typically set to a small value, such as 4 or 8, to balance the trade-off between model capacity and computational efficiency.
-3.3.2 Implementation Details
-We use exactly the same training and test sets as in §$\S$3.2, but the inputs here are the original headlines with the text information.
-To help Llama-3-8b understand the boundary between two headlines, we concatenate them using the default end of sequence (EOS) token, <|begin_of_text|>, placed between them.
-For example, for the headline pair The First Headline and The Second Headline, the combined sequence would be, <|begin_of_text|>The First Headline<|begin_of_text|>The Second Headline.
-We add a classification head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure 5.
-Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence.
-Thus, the last token in the sequence, being informed by all prior tokens, is used for classification.
-This embedding vector of the last token (which is 4096-dimensional), output by the last Transformer layer, is processed through a linear layer producing a two-dimensional logit vector, denoted by $Z=(Z_{1},Z_{2})$.
-The $Z$ vector is then transformed into probabilities by the softmax function, indicating the likelihood of each headline being catchier, i.e.,
-$\mathbb{P}\{\text{i-th headline is catchier}\}=\frac{e^{Z_{i}}}{e^{Z_{1}}+e^{Z_{2}}}\mathrm{~{}for~{}}i=1,2.$
-Figure 5: The classification pipeline using the Llama model.
-Headlines are first tokenized and each token is converted into an embedding vector.
-For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration.
-The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens.
-Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity.
-The output from the last token is fed into a classification head, which computes the probability that one headline is more impactful than the other.
-During fine-tuning, our objective is to minimize the cross-entropy loss on the training set.
-We implement LoRA fine-tuning using the PEFT library developed by Hugging Face.
-We set the dimension of the low-rank matrices as $r=4$.
-Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora).
-Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$.
-To mitigate over-fitting issues, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices.
-[^14] Furthermore, we enable 16-bit mixed precision training.
-Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly.
-This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
-With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which translates to only 0.023% of the total 7,506,636,800 parameters.
-We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training.
-A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process.
-This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
+More details about LoRA fine-tuning and the implementation are given in Web Appendix §$\S$§D.2 and §$\S$§D.3 respectively.
 3.3.1 Performance Analysis of Fine-tuning-based Approaches
 Figure 4: Loss curve and accuracy curve as the number of training epochs increases.
-We repeat the fine-tuning process three times and display the average loss and accuracy.
-For each repetition, we consistently used the same training and test sets, ensuring that any randomness in the results came solely from the fine-tuning process, not from variations in the training and test set splits.
-The final accuracies of the three replicates on the test set are 83.59%, 82.70%, and 83.93% respectively, and 83.40% on average.
-We add the best performance (red dash line) obtained using OpenAI’s embedding in §$\S$3.2 for comparison.
 We report the loss curve and accuracy curve during the LoRA fine-tuning process in Figure 4.
-We observe an accuracy of 83.40% accuracy on the test data, which is marginally better than the best result (82.50% accuracy) using OpenAI embeddings in §$\S$3.2.
-However, such marginal improvements should not undermine the utility of fine-tuning.
-Comparing the fine-tuning of Llama-3-8b with the use of OpenAI embeddings is not straightforward due to inherent differences in models’ capability.
-To highlight the benefits of fine-tuning, we observe that the MLP classification model using embeddings derived from Llama-3-8b (79.40% accuracy in Table 6) performs worse than LoRA fine-tuned Llama-3-8b (83.40% accuracy).
-This demonstrates that fine-tuning Llama-3-8b can enhance performance compared to Llama-3-8b embeddings.
-Similar to the case of the embedding-based classification model, we notice that the fine-tuning approach also suffers from some overfitting issues; see Figure 6.
-This is because Llama-3-8b is inherently complex; even though LoRA reduces the number of trainable parameters to 0.023% of the total parameters, the model remains complex.
-Common approaches to address over-fitting include reducing model complexity, adding regularization, and increasing the dataset size.
-We experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting.
-However, these modifications generally reduced performance on the training set without yielding significant improvements on the test set.
-We also consider two other versions of the LoRA fine-tuned model to document their performance:
-  • $\bullet$
-    Larger dataset: One way to improve performance and alleviate overfitting is to fine-tune the model on a larger dataset that also includes data on headline pairs that are not significantly different from each other.
-In our setting, including insignificant pairs of headlines expands the dataset from 39,158 pairs to 134,941 pairs.
-However, even with this expanded training data, accuracy on the test set continues to hover around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset.
-We refer interested readers to Web Appendix §$\S$§D.2.1 for details of this analysis and numerical results.
-  • $\bullet$
-    DistilBERT Fine-tuning: Before the widespread adoption of the term “LLM”, BERT was considered one of the best language models available, and it is still widely used as a benchmark in many NLP tasks.
-To provide a comparison with these earlier models, we therefore fine-tune DistilBERT – a distilled version of BERT that retains over 95% of BERT’s performance while containing 40% fewer parameters (sanh2019distilbert; devlin2018bert).
-Fine-tuned DistilBERT provides an accuracy of about 75% on the test dataset, which is good, but still lower than that of the fine-tuned Llama-3-8b model.
-Please see Web Appendix §$\S$§D.2.2 for details on the implementation and performance of the fine-tuned DistilBERT.
-In sum, we find that using more data and/or different base models does not lead to significant improvement in performance in our setting.
+We observe an accuracy of 46.86% on the test data, which is marginally better than the best performance of 46.28% using OpenAI embeddings in §$\S$3.2.
+However, such similar performance should not undermine the utility of fine-tuning.
+Comparing the fine-tuning of Llama-3-8b with the use of OpenAI embeddings is not straightforward due to inherent differences in the models’ capabilities.
+To highlight the benefits of fine-tuning, we observe that the Linear CTR prediction model using embeddings derived from Llama-3-8b (42.78% on the test data, reported in Web Appendix §$\S$§C.2 Table A2) performs worse than the LoRA fine-tuned Llama-3-8b.
+This demonstrates that fine-tuning Llama-3-8b can enhance performance compared to using Llama-3-8b embeddings alone.
+In Figure 4, we observe that the accuracy on the test set (46.86%) is higher than on the training set (42.09%), which is uncommon in machine learning tasks.
+The reason for this discrepancy might be that the split of the training and test datasets results in some easily distinguishable headlines being included in the test set, making the training set more challenging than the test set.
+To verify our hypothesis, in Web Appendix §$\S$§D.4, we try several additional random splits of the training and test datasets and evaluate their performance on them.
+We find that in these other splits, the test performance is slightly worse than the training performance.
 3.4 Summary of Performance of Pure-LLM Based Methods
 We summarize the performance of various pure-LLM-based methods in Table 6.
-We also compare the accuracy of human respondents on the same task.
+We also report the results of human respondents on a similar task but with paired headline comparison.
 Details of the survey questions and the accuracy analysis of human respondents are shown in Web Appendix §$\S$§A.
-Method Accuracy Notes
-Human respondents 51.58% On 4,571 human labeled pairs
-Prompt-based approach $<$65% GPT-4 with in-context learning, evaluated on 1,000 random samples
-Classification model with LLM text embeddings 82.50% OpenAI-3072E embedding model with logistic regression
-Fine-tuned LLM model 83.40% LoRA fine-tuned Llama-3-8b
-Table 7: Summarization of performance of various pure-LLM-based methods and comparison with survey results.
-All results in this table have been reported before.
-We only report the best performance of each method.
+Method Accuracy on test set Notes
+Random guess 33.02%
+Human respondents - Similar to random guess as tested on 4,571 human responses
+Prompt 39.96% GPT-4 with in-context learning
+Embedding 46.28% OpenAI-256E embedding model with linear regression
+Fine-tuning 46.86% LoRA fine-tuned Llama-3-8b
+Table 6: Summarization of performance of various pure-LLM-based methods. All results in this table have been reported before. We only report the best performance of each method.
 There are four main takeaways from these comparisons.
@@
 This is consistent with the fact that media firms (and even expert editors) rely on experimentation to decide which headlines and articles to display to users.
-Second, while simple prompt-based approaches like in-context learning with examples are marginally better than purely random guesses or human predictions, they nevertheless do not provide sufficiently high accuracy.
-Third, both embedding-based models and fine-tuning perform equally well (at least in our setting), and provide equally good performance with around 83% accuracy.
+Second, while simple prompt-based approaches like in-context learning with examples are better than purely random guesses or human predictions, they nevertheless do not provide sufficiently high accuracy.
+Third, both embedding-based models and fine-tuning perform equally well (at least in our setting), and provide equally good performance with around 46-47% accuracy on the test set.
 They each have their pros and cons.
-The main advantage of the logit classification model with OpenAI embeddings is that it is extremely easy to implement and use in business settings since it does not require any significant model training/deployment.
+The main advantage of the linear CTR prediction model with OpenAI embeddings is that it is easy to implement and use in business settings since it does not require any significant model training/deployment.
 However, embeddings from the OpenAI models are proprietary, and as such, we lack visibility into the exact nature of the embeddings.
 In contrast, fine-tuning approaches use open-source models such as Llama-3-8b, and we have access to the underlying parameters of the LLM.
-However, this transparency comes with significant costs – fine-tuning LLMs can be computationally expensive and requires non-trivial training and deployment skills that not all businesses may have or want to invest in.
+However, this transparency comes with costs – fine-tuning LLMs requires non-trivial training and deployment skills that not all businesses may have or want to invest in.
 The fourth and most important takeaway is that while pure-LLM-based approaches are informative, they are unable to reach the accuracy of experimental approaches.
 That is, firms cannot replace content experimentation with LLM-based predictions.
+An important caveat here is consumer preferences may have changed significantly since the time period of the experiment (2013–2015).
+As such, the prediction numbers shown here may not represent LLMs’ ability to predict the preferences of consumers today.
+[^11] This concern is somewhat assuaged by Web Appendix §$\S$§C.4, which shows that consumer preferences did not change significantly during the 2013–2015 timeframe (at least within the prediction capabilities of the LLMs).
+However, we cannot empirically comment on the period beyond 2015.
+As such, the performance of LLM-based methods in §$\S$3 should be mainly used for relative comparisons with each other.
+More broadly, both LLM capabilities and consumer preferences will continue to evolve.
+Therefore, if a firm/manager is concerned about the generalizability/accuracy of a specific LLM-based prediction model to their current business environment, the simplest solution is to retrain their LLM-based models more frequently and collect up-to-date labeled data to maintain relevance.
 Finally, we discuss the monetary costs associated with three LLM-based approaches.
-The cost of prompt-based, embedding-based, and LoRA fine-tuning in our analysis are around \$20, \$0.18, \$3, respectively.
+The cost of prompt-based, embedding-based, and LoRA fine-tuning in our analysis are around \$15.87, \$0.29, \$3, respectively.
 The cost of prompt-based and embedding-based comes from the access to OpenAI’s API in order to get responses and text embeddings.
@@
 4 Our Approach: LLM-Assisted Online Learning Algorithm
-So far, we have seen that pure-LLM-based approaches can help with headline selection with reasonably high accuracy (around 83% for the embedding-based and LoRA fine-tuning approaches).
-The main advantage of pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is better and go with it.
+So far, we have seen that pure-LLM-based approaches can help identify the best headline with 46.86% accuracy using the LoRA fine-tuning approach.
+The main advantage of pure LLM-based approaches is that they can be deployed without any experimentation, i.e., the manager can simply use the LLM-based model to predict which headline is the best and go with it.
 Thus, there is no need to use any traffic for experimentation.
-The downside, of course, is that these models are not fully accurate – in a pairwise comparison, in around 17% of the cases, they are likely to mispredict which headline is better.
-This can, of course, lead to lower clicks and revenues for the firm.
-[^15] It should be noted that even this accuracy is calculated under the assumption that the results from the A/B tests are the ground truth.
-However, even if we were to reach 100% accuracy on this dataset, we still may not reach 100% accuracy in practice because of Type I errors in the creation of significant samples.
-Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
+The downside, of course, is that these models are not fully accurate – in some cases, they are likely to mispredict which headline is the best.
+This can, of course, lead to lower clicks and revenues for the firm.Thus, it is not clear that relying on pure-LLM-based methods will lead to better overall outcomes for the firm.
 On the other hand, we have the current status quo, which is the use of experimentation-based approaches to choose headlines/content.
@@
 In the first step, we use LLMs to get an initial prediction of each headline’s CTR.
-Note the prediction task here is somewhat different from that in §$\S$3.
-In the previous section, we used LLM-based approaches to perform pairwise comparisons of headlines and predict which headline amongst a pair is more engaging.
-However, as discussed earlier, that approach does not work when we have more than two headlines (which is the case for over 90% of the A/B tests in our data).
-Further, we need actual CTR predictions for each headline to feed into the second step of our algorithm.
-Therefore, we adapt the LLM-based classification models from §$\S$3 for predicting CTRs by making the following modifications: changing the loss function from log loss to mean squared error (MSE), using single headline inputs instead of headline pairs, and switching from binary output to continuous CTR prediction.
-The training and fine-tuning procedures remain unchanged.
-[^16] Using CTR as the outcome measure does not adversely affect the performance.
-For instance, using linear regression with 256-dimensional text embeddings from OpenAI to predict CTR for significant pairs and labeling the headline with the higher predicted CTR as the winner yields an accuracy of around 75%, which is very close to the 77.31% accuracy obtained with logistic regression.
-This minor difference is likely due to the inherently skewed nature of the CTR distribution and the sensitivity of MSE to outliers.
-We compare two approaches for CTR prediction: an OpenAI-embedding-based method and a LoRA fine-tuned Llama-3-8b model.
-Our numerical results indicate that the R-squared values for the embedding-based approach and the LoRA fine-tuned method are 0.241 and 0.315, respectively.
-When selecting the headlines with the highest predicted CTRs, the accuracy—defined as the percentage of times the chosen headline indeed has the highest actual CTR—is 42.66% for the embedding-based method and 46.55% for the LoRA fine-tuned method.
-Note that this accuracy is significantly lower than the binary classification accuracy discussed in §$\S$3 because we can have more than two headlines in a test and many headlines have no significant difference in CTRs.
-Based on these results, we select the LoRA fine-tuned Llama-3-8b model as the best CTR prediction model in the first step.
-Additional details on the CTR prediction methods and their results can be found in Web Appendix §$\S$§F.
+Based on numerical results in §$\S$3, we select the best-performing approach, LoRA fine-tuned Llama-3-8b model, as the best CTR prediction model in the first step.
 In the second step, we build on the bandit framework from §$\S$4.1 to design an online algorithm that incorporates the LLM-based CTR predictions from the first step.
@@
 After playing the arm and observing the outcome, we update the impressions, clicks, and mean reward accordingly, following standard bandit algorithm procedures.
-The 2-UCBs algorithm offers three advantages.
+The intuition behind using two UCBs instead of a single UCB incorporating LLM predictions $U_{k}^{2}$ is as follows: if we rely solely on $U_{k}^{2}$, and the LLM’s CTR predictions are significantly incorrect, i.e., overestimating the CTR of a poorly performing headline (underestimating the CTR of the best headline is of less risk because there are more suboptimal headlines than the one and only best headline), it would take many rounds to correct this error, especially when the number of auxiliary samples, $n^{\text{aux}}$ is large.
+On the other hand, the first UCB $U_{k}^{1}$, though unbiased, suffers from large stochastic noise, requiring large $n_{k}^{t}$ to get an accurate estimator.
+By choosing the minimum of the two UCBs, we balance these opposing risks effectively.
+However, the formalism for this approach is only partially addressed in prior literature (gur2022adaptive), which operates under a different setting without LLMs.
+In particular, (gur2022adaptive) shows 2-UCBs is optimal under the assumption that the auxiliary data is generated from a linear function class, including the standard UCB setting with exogenous auxiliary samples.
+The formalism in our context would require additional assumptions about the errors in the LLM’s predictions, which remains an open research question with limited understanding.
+We leave this formalization for future work.
+The LOLA algorithm offers three advantages.
 First, it is easy to implement, requiring the tuning of only two hyperparameters.
@@
 Because it can easily accommodate the edge cases, e.g., one where the manager has no auxiliary information or one where the manager wants to fully rely on the LLM predictions, there is no need for extra engineering effort to set up separate systems for different edge cases.
-This adaptability is extremely valuable given the rapid advances in LLM technology.
+This adaptability is valuable given the rapid advances in LLM technology.
 When the capability of the pure-LLM method continues to advance, the manager can easily fine-tune and increase $n^{\text{aux}}$ to take more advantage of LLM.
+As a final note, our algorithm does not update the LLM during the online learning phase.
+Instead, we use the LLM fine-tuned on historical data solely for initialization purposes.
+That being said, the algorithm can be readily modified to fine-tune LLMs with online data collected during the learning phase.
+While such fine-tuning could potentially improve LOLA’s performance, the improvement is expected to be limited in our current dataset.
+As shown in Figure A3 in Appendix §C.3, achieving a significant performance boost in LLMs would require a much larger dataset than is currently available to us.
+Discussion: We now provide a broader discussion on LOLA.
+We note that the general idea that priors can impact regret and rewards is not new and has existed for some time; see bubeck2013prior and russo2014learning.
+These earlier papers primarily focus on theoretical analyses and often rely on assumptions about the quality of priors.
+However, since we do not have a theoretical understanding of LLM predictions, it is unclear how these predictions can be translated to priors or the nature of such priors.
+Separately, the 2-UCBs algorithm by gur2022adaptive was designed for a setting where the experimenter has access to auxiliary data.
+However, their algorithm assumes that the researcher has access to both the size of this auxiliary sample ($n^{aux}$) and knowledge of the prediction error in this sample.
+Thus, neither of these earlier approaches is directly applicable to our setting.
+A key challenge in our setting is that the LLM predictions cannot be naively treated as priors since we do not know their prediction error or theoretical properties.
+The main novelty of LOLA comes from the idea that we can treat the LLM predictions as coming from a “pseudo-sample” and that we can fine-tune the size of this pseudo-sample ($n^{aux}$) for a given application based on prior data.
+Thus, we build on the 2UCBs algorithm but modify it to accommodate LLM predictions – that do not come from a real sample and whose quality and theoretical properties are unknown.
 4.3 Benchmarks and Implementation Details
@@
 We use the training data to learn a model that predicts the CTR of a headline based on an LLM (either through text embeddings or LoRA fine-tuning).
-In our numerical study, we use the best-performing pure-LLM-based approach based on our experiments (see Web Appendix §$\S$§F), i.e., the LoRA fine-tuned Llama-3-8b with MSE loss, which gives a 46.55% accuracy.
+In our numerical study, we use the best-performing pure-LLM-based approach based on our experiments, i.e., the LoRA fine-tuned Llama-3-8b, which gives a 46.86% accuracy.
 [^12] Note that since this is a pure prediction-based greedy approach, there are no hyper-parameters associated with the algorithm, and hence we do not use the fine-tuning data.
@@
 Based on fine-tuning, we choose $n^{\text{aux}}=1000$ and $\alpha=0.08$ for our setting.
-    Note that the former two methods can be viewed as special cases of our proposed LOLA algorithm.
+We refer interested readers to Web Appendix §$\S$§F for a detailed sensitivity analysis of $n^{\text{aux}}$.
+    Note that the first two methods can be viewed as special cases of our proposed LOLA algorithm.
 Both our algorithm and the pure LLM-based approach use the exact same CTR prediction algorithm.
@@
 Again, note that we use the same hyper-parameter $\alpha=0.08$ across both algorithms, which ensures that our algorithm defaults to the standard UCB with $\alpha=0.08$ when we ignore the CTR predictions from the LLM.
+  • $\bullet$
+    UCB with LLM priors: To demonstrate the effectiveness of the 2-UCBs policy $\min\{U^{1},U^{2}\}$ compared to using only the second UCB $U^{2}$, we also run the UCB with LLM as priors, i.e., play the arm $a_{t}=\arg\max_{k\in[K]}U_{k}^{2}$.
+This comparison highlights how selecting the minimum of the two UCBs improves performance by mitigating the risk of overestimated CTRs from the LLM predictions.
 After the fine-tuning of algorithms, we evaluate all algorithms on the test dataset, which has never been used for either CTR training or algorithm tuning.
@@
 Note that the Y-axis captures the average clicks per test per period.
-For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
+For instance, if there is a test with two headlines receiving 1 and 2 clicks, respectively, under $\tau=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
 The Y value is simply the average of this number $0.03$ over all tests.
 This measure scales well with the platform’s total clicks in tests because headlines in different tests with different numbers of headlines take the same weight in this measure.
-We visualize the results from our numerical simulations in Figure 7 and document the pairwise comparison of different algorithms, along with the relative percentage reward improvement and the p-values from the t-tests in Table 8.
-We find that, on average, LOLA performs the best among all algorithms across all time horizons.
-We document a few additional patterns of interest.
+We visualize the results from our numerical simulations in Figure 5131313The Y metric in Figure 5 is essentially the accumulated reward divided by $\tau$, i.e., the value of the X-axis.
+We choose to report this Y metric instead of the actual regret for two reasons: (1) the ground truth arm is not known in practice, making it impossible to compute regret; and (2) for better visualization.
+Specifically, the accumulated reward (unscaled by $\tau$) grows very rapidly as $\tau$ increases, which compresses the visual differences between algorithms, making them difficult to distinguish with the naked eye.
+With this being said, the results in Table 7 can still be interpreted as the percentage improvement of the accumulated reward because all algorithms are measured at the fixed $\tau$ and the accumulated reward $=$ Y ×$\times\tau$., and document the pairwise comparison of different algorithms, along with the relative percentage reward improvement and the p-values from the t-tests in Table 7.
+We find that, on average, LOLA performs the best among all algorithms across all time horizons.141414 Note that we cannot translate the click improvement to revenues or other business outcomes of interest since we do not have data on how clicks/time spent on the website relate to downstream metrics like advertising and subscription revenues.
+Therefore, we do not make additional claims on the business impact of these numbers.
+That said, if the firm has access to these numbers, it should be relatively straightforward to make that connection.
 Parameter $\tau$ LOLA vs UCB LOLA vs Pure LLM LOLA vs E&C UCB vs Pure LLM UCB vs E&C Pure LLM vs E&C
@@
 1000 $2.23$∗∗∗∗ $5.02$∗∗∗∗ $5.08$∗∗∗∗ $2.73$∗∗∗∗ $2.79$∗∗∗∗ $0.05$
-Table 8: Percentage % reward improvement between two algorithms with significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001.
+Table 7: Percentage improvement in reward (total clicks) between the algorithms LOLA (LLM-2UCBs), UCB, Pure LLM, and E&C (we omit UCB with LLM priors here due to space constraints, but report its performance relative to LOLA in Figure 5).
+Significance levels: $p\leq$ 0.05, ∗∗ $p\leq$ 0.01, ∗∗∗ $p\leq$ 0.001, ∗∗∗∗ $p\leq$ 0.0001.
 For example, under the impression per headline equal to 50, the average click per test of LOLA and UCB is 0.053760 and 0.050267, respectively; so the relative improvement is calculated as ${0.053760}/{0.050267}-1=6.95\%$.
 Parameter $\tau$ represents the average impression per headline to scale the time horizon $T=\tau\times K$.
+We now document a few additional patterns of interest.
 First, the Pure LLM approach exhibits consistent performance across different $\tau$ values because it is not an adaptive algorithm.
@@
 However, as the horizon increases, LOLA will start showing significant gains.
-Next, comparing LOLA with regular UCB, we see that it always outperforms UCB.
+Second, comparing LOLA with regular UCB, we see that it always outperforms UCB.
 Interestingly, here we see that, as the time horizon increases, the relative improvement from LOLA diminishes.
@@
 While E&C shows better performance as the horizon increases, even when $\tau=1000$, LOLA is still better than E&C by 5.08%, which is a large improvement in the media and publishing industry.
-Finally, comparing the benchmark algorithms with each other, we see that when the horizon is short, the Pure LLM does better, while UCB is better when the horizon is longer.
+Third, comparing the benchmark algorithms with each other, we see that when the horizon is short, the Pure LLM does better, while UCB is better when the horizon is longer.
 In general, this reflects the intuition that active learning is not very useful when the time horizons are very short since the algorithm has no time to explore and learn the CTRs effectively (and then exploit this learning).
@@
 This is because E&C neither exploits LLM information and nor does it actively learn over time.
+Finally, we will discuss the performance of UCB with LLM priors.
+Comparing LOLA (2-UCBs) with UCB with LLM priors, the key difference lies in whether the regular UCB $U^{1}$ is included in the algorithm.
+The core trade-off here is between the large potential error of $U^{2}$ when the LLM’s predictions are incorrect and the stochastic noise in $U^{1}$, especially in the early stages.
+Our observations confirm this trade-off.
+In short time horizons $\tau=50$, LOLA underperforms compared to UCB with LLM priors, by a significant margin of $-0.74\%$.
+This is because the error in the LLM-based CTR predictions is smaller than the stochastic noise of the regular CTR estimator at the start, and including the noisier $U^{1}$ reduces performance.
+However, as the time horizon increases (e.g., $\tau=800$ and $\tau=1000$), LOLA significantly outperforms UCB with LLM priors by $0.37\%$ and $0.47\%$, respectively.
+This improvement occurs because, over longer periods, the error in LLM-based CTR predictions becomes more significant than the stochastic noise of the regular estimator, which can also be observed by comparing the performance between the standard UCB and Pure LLM methods.
+Incorporating the unbiased and less noisy $U^{1}$ helps enhance performance in these longer horizons.
+This observation suggests that it is especially beneficial to use 2-UCBs in longer horizons.
 In summary, we find that LOLA outperforms existing approaches for experimentation by leveraging the strengths of LLMs and combining them with the benefits of active/online learning.
-In particular, our results demonstrate the value of LOLA as the time horizon increases, and highlight the limitations of relying solely on pure LLM predictions or the naive E&C approach.
+In particular, our results demonstrate the value of LOLA as the time horizon increases, highlighting the limitations of relying solely on pure LLM predictions or the naive E&C approach.
 4.5 LOLA Variants and Extensions
@@
 However, click-based metrics have raised concerns about promoting clickbait headlines, which may reduce user engagement or spread negativity.
-As a result, the news and media industry is now evolving to focus more on content quality and humanity, although clicks still remain an important measure; see (Upworthy2023) for a more detailed discussion.
+As a result, the news and media industry is now evolving to focus more on content quality and longer-term user engagement and retention metrics, although clicks still remain an important measure; see (Upworthy2023) for a more detailed discussion.
 The LOLA framework can be easily adapted to alternative reward measures that align with new business goals, e.g., time spent on the platform during a session, or time spent on the article (and not just clicks on the headline).
@@
 Specifically, we ask, “How often do you read the news?” with options “Never”, “1-2 days per week”, “3-4 days per week”, “5-6 days per week”, “Everyday”.
-Detailed information on the surveys and the results from both surveys are shown in Table A1.
-Note that the total responses received is less than the number of respondents ×$\times 20$ because some students did not finish the survey or skipped some comparisons.
-For example, for the first survey, the total number of responses 3,809 is less than $202\times 20=4,040$.
-To calculate the accuracy metric, we only focus on the significant pairs, and in the first survey, we discard the uncertain answer “Both headlines seem equally good” for a relatively fair comparison between the two surveys.
-“Accuracy” measure here is consistent with the LLM tests.
-We treat the headline with a higher CTR as the correct answer and calculate the proportion of correct answers as accuracy.
-By doing so, one should expect 50% accuracy by random guess since
 Detailed information and survey results are documented in Table A1.
@@
 We regarded the headline with the higher CTR as the correct answer and calculated the proportion of correct responses as accuracy.
-As such, we should expect a 50% accuracy rate from random guessing.
+As such, we should expect a 50% accuracy rate from random guessing in this paired comparison.
 The survey results indicate that respondents’ ability to identify catchier headlines is comparable to random guessing – the accuracy rate in both surveys is not significantly different from 50% at the 0.05 significance level.
@@
 B.1 Pairwise Comparison Between GPT Prompts
-To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on 1,000 correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other.
+To determine whether one prompt is significantly better than the other, we record a binary indicator of the correct/incorrect answer for each prompt on each sample and then conduct a paired t-test on correct/incorrect answer indicators to see whether the answers from one prompt are significantly more accurate than the other.
 The mean differences and p values are documented in Figure A1 and Figure A2.
-The random guess model simply predicts the first headline wins with a half probability.
 The nomenclature for GPT prompts in these two figures follows a specific format.
-The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are flipped.
+The first component is the name of the GPT model, the second component indicates the number of demonstrations, and the third component specifies whether the outputs in the demonstrations are correct or not.
 A value of 0 indicates the correct label, while a value of 1 indicates the wrong label.
@@
 Figure A2: The number in the heatmap presents the p-value from the paired t-test between different GPT prompts.
-Appendix C Appendix to Embedding-based Classification Approaches
-C.1 Implementation Details of MLP Classification Models
+Appendix C Appendix to Embedding-based Approaches
+C.1 Implementation Details
+We start by introducing three embedding models used:
+  • $\bullet$
+    OpenAI embeddings: At the time of this study, the latest and best embedding model available from OpenAI was text-embedding-3-large.
+These embeddings are obtained using the OpenAI embedding API in Python.
+This embedding model generates embeddings with up to 3,072 dimensions and supports a maximum token limit of 8,191, with its knowledge base extending up to September 2021.
+A notable feature of this model is its flexibility – it allows users to tailor the embedding dimension to their specific needs.
+[^16] Other models capable of producing text embeddings, like BERT (devlin2018bert) and Sentence-BERT (reimers2019sentence), have fixed embedding vector dimensions that remain constant once the model is trained.
+These dimensions cannot be changed or resized for different applications.
+However, OpenAI’s embedding model uses a technique that, while producing outputs of fixed length, encodes information at different granularities (kusupati2022matryoshka).
+This allows a single embedding to adapt to the computational constraints of downstream tasks.
+More specifically, downstream tasks can use only the first several dimensions of the original vector, and the information retained decreases smoothly as a shorter part of the vector is used.
+In our experiments, we consider both 3,072- and 256-dimensional embeddings to examine whether the size of the embedding vector size affects the performance of the downstream CTR prediction models.
+[^17] A higher dimension does not necessarily mean better performance.
+While a higher dimension can retain more information for each sample, it also means that the samples are spread more sparsely in a higher-dimensional space given a limited number of training samples.
+This sparsity can increase the possibility of over-fitting.
+  • $\bullet$
+    Llama-3-8b: Meta’s Llama-3 (metallama3) is regarded as one of the most advanced open-source LLMs to date (llama3modelcard).
+We obtain embeddings from Llama-3 with 8 billion parameters (Llama-3-8b).
+We fine-tune this exact model using LoRA in §$\S$3.3; therefore, using the embeddings from this model allows for a direct comparison of fine-tuning vs.
+embedding-based methods (see §$\S$3.4 for comparisons).
+Llama-3-8b does not produce embeddings explicitly designed to represent the semantic meaning of entire input sequences.
+To address this issue, we apply the technique described in (behnamghader2024llm2vec), which transforms the Llama-3-8b model into a more suitable text embedding model, outputting embeddings with a dimension of 4,096.
+  • $\bullet$
+    Word2Vec embeddings: Word2Vec, proposed by mikolov_etal_2013a, was the leading embedding model before the widespread adoption of Transformers in the Natural Language Processing (NLP) literature.
+Unlike LLM-based text embeddings, Word2Vec does not use Transformers and cannot pay attention to surrounding token information.
+Embeddings from Word2Vec have been used in the recent marketing literature in a variety of use cases, including product attribute modeling (timoshenko_hauser_2019; wang_etal_2022), the effect of creativity on product popularity (sozuer_etal_2024).
+We compute the Word2Vec embedding vector for each word in the headline and average them to get the headline embedding, a common practice in Word2Vec.
+To ensure consistency with the dimensions used for OpenAI embeddings, we consider Word2Vec embeddings with 256 and 3,072 dimensions.
+  • $\bullet$
+    BERT embeddings: BERT (Bidirectional Encoder Representations from Transformers), introduced by devlin2018bert, is a seminal model in natural language processing that generates context-sensitive embeddings for text.
+Unlike earlier models that produce static word embeddings, BERT captures the nuances of word meanings based on their context within a sentence, resulting in more accurate representations.
+The standard BERT model produces embeddings with 768 dimensions, which are fixed once the model is trained.
+This fixed dimensionality can be a limitation when adapting to tasks with varying computational constraints.
+Despite this, BERT’s ability to understand context has led to its widespread adoption in various NLP applications, including sentiment analysis, question answering, and named entity recognition.
+To assess the impact of model complexity on performance, we consider using a simple linear model and a multilayer perceptron (MLP) model.
+MLP is a powerful neural network architecture consisting of multiple linear and non-linear hidden layers that allow the model to learn complex patterns in the data.
+We use a fully connected neural network with ReLU activation for hidden layers, with dropout layers added to prevent over-fitting.
+We refer interested readers to Chapter 5 in (zhang2021dive) for more details.
+We train the MLP using the stochastic gradient descent algorithm to minimize binary cross-entropy and perform hyper-parameter tuning with Bayesian Optimization to determine the optimal configuration, with the training process and the architecture details below.
 Hyperparameters of the network structure.
-We employ a fully connected neural network with ReLU activation functions for the hidden layers and a sigmoid activation function for the output layer.
-This configuration is widely used in practice.
-To determine the number of layers and the hidden nodes of each layer, we choose the number of layers within the range of $\{1,2,3,4,5\}$, and set the number of nodes in the first fully connected layer from the set $\{128,256,512,1024,2048\}$.
-Besides, we adjust the previous-post ratio, which is the proportion of nodes in each layer relative to those in the subsequent layer.
-This ratio here is in $\{2,4,8\}$.
-Note that the previous-post ratio becomes irrelevant to the results when there’s only one hidden layer.
-This is because the final output layer has only one dimension for ranking the most attractive headline.
-We incorporate the dropout after each hidden layer to mitigate the over-fitting.
-Dropout layers randomly deactivate neurons based on a specified dropout rate, while scaling up the remaining neurons to maintain the same mean of the layer.
-In this work, we set the dropout rate in $\{0.3,0.5\}$.
-All hyperparameters are automatically selected through hyperparameter optimization, which is introduced below.
-Hyperparameters of the training procedure.
-The MLP is implemented in PyTorch, with data split into 70% for training, 10% for validation, and 20% for testing.
+We employ a fully connected neural network with ReLU activation functions for the hidden layers, a configuration widely used in practice.
+The network consists of two hidden layers, each with 128 nodes.
+To mitigate overfitting, we incorporate dropout after each hidden layer.
+Dropout layers randomly deactivate neurons based on a specified dropout rate while scaling up the remaining neurons to maintain the same mean of the layer.
+We set the dropout rate to 0.5.
+We did not further test other hyperparameter settings, as this configuration already delivers sufficiently good performance.
+However, more careful hyperparameter optimization could potentially improve results.
+Training Procedure Hyperparameters.
+The MLP is implemented in PyTorch, with the data split into 70% for training, 10% for validation, and 20% for testing, as introduced in §$\S$3 .
 To prevent the network from memorizing the order of samples, we shuffle the training set at the start of each epoch.
-Determining the appropriate batch size is crucial, as large batch sizes may lead to convergence to local optima, while small batch sizes may hinder convergence.
+Determining the appropriate batch size is crucial, as large batch sizes can lead to convergence to local optima, while small batch sizes may slow down convergence.
 We balance accuracy and training speed by setting the batch size to 10 and the learning rate to $0.01$.
-Additionally, to mitigate over-fitting, we apply $L^{2}$ regularization, with the regularization weight chosen from $\{1e-6,1e-4,1e-2\}$.
-We use binary cross-entropy as the loss function and employ stochastic gradient descent for training.
-We implement early stopping by monitoring the validation set loss, halting the training process if the reduction in loss over 60 epochs falls below a tolerance threshold of $10^{-4}$.
-Hyperparameter optimization.
-We optimize the parameters mentioned above through Bayesian Optimization (snoek2012practical).
-This Bayesian technique uses Bayesian inference to construct a probabilistic model of the objective function, where the parameters are updated iteratively with new observations.
-This iterative process balances exploration of the parameter space to discover promising regions and exploitation of known high-performing areas.
-In this work, we consider the accuracy of the validation set as the objective and optimize hyperparameters selected from the specified parameter space, as delineated by the ranges mentioned previously.
-We choose the Gaussian process as the estimator with the automatic optimizer, which is commonly adopted by the community.
-To ensure a judicious trade-off between efficacy and efficiency, we repeat the optimization process 50 times for input data of 256 dimensions and 30 times for data of 3,072 dimensions.
-C.2 Effect of Headline Similarity on Performance
-We now examine the conditions under which these approaches fail.
-, Specifically, we investigate whether headline similarity affects the performance of embedding-based classification models.
-We first calculate the cosine similarity based on the text embeddings for each pair of headlines.
-This metric captures the similarity between two vectors in inner product space and is defined as the dot product of the vectors divided by the product of their lengths, i.e., ${e_{1}\cdot e_{2}}/(\|e_{1}\|\|e_{2}\|)$, where $e_{1}\cdot e_{2}$ is the dot product of the embedding vectors, and $\|e_{1}\|$ and $\|e_{2}\|$ are the their Euclidean norms.
-Figure 3(a) shows the distribution of cosine similarity across all headline pairs, whereas Figure 3(b) shows the distributions of cosine similarity separately for two cases: (1) cases where our embedding-based classification correctly predicts which headline is catchier, and (2) cases where our embedding-based classification incorrectly predicts which headline is catchier.
-We observe a rightward shift in the distribution of incorrect predictions (compared to the distribution of correct predictions).
-This suggests that incorrect predictions are more common when the headlines are very similar to each other (i.e., higher cosine similarity).
-There are two possible reasons for this.
-First, in cases where the headline pairs are very similar, there is a higher chance of Type 1 error (in the t-tests based on our A/B test data), i.e., the headlines may not be significantly different in fact.
-Second, text embeddings may not be able to capture the differences between two headlines when they are similar, i.e., the embeddings may also be very similar.
-As a result, it is hard for the classification model to differentiate between them.
-[^19] Note that it is unlikely that this phenomenon is due to problems/issues with the classification model because both Logit and MLP achieve nearly identical performance.
-(a) Cosine similarity distribution over all pairs in training and test sets
-(b) Cosine similarity over incorrect/correct predicted test pairs
-Figure A3: The left panel shows the distribution of cosine similarity for all pairs.
-The right panel shows the distributions of cosine similarity for both incorrect and correct predictions using the logit classification model with OpenAI-3072E embeddings.
+The loss is calculated as the mean squared error between the real CTR and the predicted CTR, and we use stochastic gradient descent for training.
+C.2 Performance Comparison of Different Embeddings
+Embedding Model CTR Prediction Accuracy on the test set
+OpenAI-256E Linear 46.28%
+MLP 44.38%
+OpenAI-3072E Linear 43.06%
+MLP 45.17%
+Llama-3-8b-4096E Linear 42.78%
+MLP 36.03%
+Word2Vec-256E Linear 41.07%
+MLP 36.96%
+Word2Vec-3072E Linear 40.27%
+MLP 37.27%
+BERT Linear 42.32%
+MLP 35.89%
+Table A2: The performance of different embedding and CTR prediction models on test data.
+All models use the same training data, and accuracy is evaluated on the test data.
+Note that the Linear model’s training is a deterministic process after fixing the data, while MLP training incurs randomness via stochastic gradient descents.
+Therefore, we rerun MLP training three times with different random seeds and report the average accuracy from three reruns.
+The comparison of different embeddings and CTR prediction models are listed in Table A2.
+We observe that OpenAI embeddings outperform Llama-3-8b, Word2Vec, and BERT-based models.
+This is because OpenAI’s LLMs are state-of-the-art models that naturally perform better than the other embeddings.
+Additionally, the performance of embeddings from the open-source Llama-3-8b falls between that of the proprietary OpenAI model and Word2Vec/BERT, which does not use the transformer architecture.
 C.3 Effect of Training and Test Data Size on Performance
-(a) Performance of OpenAI-256E
-(b) Performance of OpenAI-3072E
-Figure A4: Performance of Logit classification models with OpenAI embeddings under varying training dataset sizes and different embedding sizes.
-We now conduct experiments varying the proportion of data used for training vs.
-test.
-We use the Logit model for these experiments since it offers the best performance and is easy to train.
-The results from this exercise are shown in Figure A4.
-As we can see, when we use the 256-dimensional embeddings, approximately 65% of the training set, comprising approximately 25,000 pairs (with the remaining 35% used for the test set), is adequate to achieve high and stable performance.
-However, when we use the 3072E-dimensional embedding, there is a clear over-fitting [^20] Over-fitting refers to the phenomenon that the accuracy on the training set is consistently higher than the accuracy on the test set.
-It occurs when a machine learning model learns the details and noise in the training data to the extent that it negatively impacts the model’s performance on new, unseen data.
-This often happens because the model becomes too complex or because the training set is too small, causing the model to memorize specific examples and noise rather than general patterns.
-issue even if we use the simplest Logit model and 95% of the data for training.
-This suggests that for larger embedding vectors, we need much larger training datasets.
-[^21] In principle, overfitting can also be due to model complexity.
-However, this is unlikely to be the case in our setting since we are using a simple Logit model with linear features.
-C.4 Dimensionality Reduction and Regularization
-In this section, we examine whether employing dimensionality reduction techniques and/or regularization can improve performance and reduce overfitting.
-C.4.1 PCA for Embeddings
-One can observe that OpenAI-256E in Figure A4(a) does not suffer from the over-fitting in contrast to OpenAI-3072E as shown in Figure A4(b).
-Thus, one straightforward method one can try to attack the over-fitting issue is reducing the dimension of embeddings using PCA or directly using lower-dimensional OpenAI embeddings.
-However, as we have discussed that OpenAI-256E has a worse performance than OpenAI-3072E regarding the accuracy in the testing dataset, we turn to PCA to see if PCA-reduced embeddings can improve the performance by solving the over-fitting issue.
-Specifically, we first use PCA to reduce 3072-dimensional OpenAI embeddings.
-Then, we concatenate two reduced embeddings from healing pairs and use Logit for the downstream classification task.
-The result is displayed in Figure A5, which shows that as the dimension is reduced, although there is a smaller accuracy gap between the training and test set, the accuracy on both datasets decreases, indicating that PCA does not work for our task, which prefers higher testing accuracy.
-Figure A5: Accuracy of the Logit classification model with varying PCA-reduced OpenAI embeddings as inputs.
-C.4.2 $L^{1}$, $L^{2}$ Regularization
-[…]
-Figure A6: Accuracy of the Logit classification model with varying $L^{1}$ (left) and $L^{2}$ (right) regularization weights. X-axis is the regularization weight defined as $1/\texttt{C}$.
-In sum, we find that neither PCA nor regularization improves the performance accuracy on the test data, which suggests that over-fitting is unlikely to be due to learning noise in the training data.
-Our analysis suggests that we need more A/B tests to learn the relationship between a headline’s text embeddings and its attractiveness effectively.
-In summary, these comparisons suggest that data samples from Upworthy are insufficient to unlock the potential of LLMs fully.
-Nevertheless, diverting more and more traffic to A/B tests is costly, and this observation is one of the motivations for our approach later in §$\S$4.
-C.5 Potential Threats and Robustness Checks
+In this analysis, we focus on OpenAI-256E (as it provides the best performance) and use linear regression as our prediction model, since the MLP does not significantly outperform the linear model on OpenAI embeddings and is easier to train.
+Figure A3: Performance of CTR prediction model with OpenAI-256E under varying training dataset sizes
+To assess the effect of training sample size on accuracy, we vary the training data size and measure performance.
+Specifically, we set a training size ratio, randomly select a subset of articles based on this ratio from the full training set, and use this subset to train the model.
+As shown in Figure A3, increasing the amount of training data leads to a gradual improvement in test accuracy.
+For example, increasing the training size from 20% of the total dataset (about 30% of the training set) to 70% of the total dataset (100% of the training set) raises accuracy from 44% to 46% on the test data.
+This observation also indicates that, achieving a significant performance boost in LLMs would require a much larger dataset than is currently available.
+C.4 Potential Threats and Robustness Checks
+Note that in the following two analyses of threats, we still focus on OpenAI-256E with linear regression as our prediction model as it provides the best performance.
 Next, we discuss two potential threats to the inference from our analysis here and our robustness checks to alleviate concerns around these threats.
-First, one may be concerned as to whether Upworthy’s data has been used as part of OpenAI’s training corpus – which could potentially invalidate our analysis.
+First, one may be concerned about whether Upworthy’s data has been used as part of OpenAI’s training corpus, which could potentially invalidate our analysis.
 However, this is unlikely for several reasons.
 First, it is improbable that OpenAI would have specifically downloaded these Upworthy CSV files (Comma-Separated Values files) dataset from a subfolder in Upworthy Research Archive.
-Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than paired headlines with winner information.
-Therefore, during the training of OpenAI models, it is unlikely they were exposed to the winner information since the training objective is to predict the next word.
-Third, we conducted an experiment to determine if this dataset was indeed a part of OpenAI’s corpus.
+Second, the original data only contains impressions and clicks for each headline, as shown in Table 1, rather than CTR information.
+Therefore, during the training of OpenAI models, it is unlikely they were exposed to the CTR information since the training objective is to predict the next word.
+Third, we further conducted an experiment as the robustness check to determine if this dataset was indeed a part of OpenAI’s corpus.
 We used embeddings from two headlines with significantly different impressions (possibly from different tests) for the downstream task of predicting which headline received higher impressions.
 If the embedding model was trained on this corpus, its embeddings should reflect the number of impressions and perform well on this task.
-However, the observed accuracy was approximately 50%, indicating that the embeddings do not contain such information, confirming our assumption.
-Nevertheless, if a firm is concerned that its data is a part of public repositories or is concerned about privacy issues around submitting headlines to OpenAI to obtain embeddings, then instead of using embedding-based approaches, it can rely on fine-tuning based approaches; see §$\S$3.3 for details.
-Second, one may be concerned that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headline was created, making our classification task invalid.
-To address this concern, we conducted a robustness check by splitting the dataset based on the creation time of the headlines (e.g., the initial 80% of the data, in chronological order, as the training set, and the subsequent 20% as the test set).
-We found its accuracy was 81.50% (close to the best embedding accuracy of 82.50%), indicating that time shift preference is not a concern in our data.
+However, the observed accuracy was approximately 50%, indicating that the embeddings do not even contain such impression information, confirming our assumption.
+Second, one may be concerned that readers’ preferences changed over time and that the current sample splitting method fails to take into account when the headlines were created, potentially making our CTR prediction model invalid if users’ preferences shift.
+To address this concern, we conducted a robustness check by splitting the dataset based on the creation time of the headlines.
+Specifically, we used the initial 70% of the data, in chronological order, as the training set, and the latest 20% as the test set (the remaining 10% was reserved for algorithm hyperparameter fine-tuning).
+We found that the accuracy was 46.25% on the test data, which is close to the embedding performance on the randomly split training and test dataset, with an accuracy of 46.28%, as given in Table 5.
+This indicates that time-based preference shifts are not a concern in our data.
 Appendix D Appendix to LoRA Fine-Tuning
@@
 It is worth noting that LLMs do not necessarily operate in a conversational manner, nor do they always use language as their output.
-As we will demonstrate in §$\S$3.3.2, we employ the Llama model as a classifier, where the output is the distribution of predicted classes.
+As we will demonstrate in the following, we employ the Llama model as a CTR prediction model.
 In applications like conversational AI, exemplified by GPT, the contextualized embeddings from the last token of the last layer of the Transformer are fed into a linear layer followed by a softmax function, the configuration of which transforms these embeddings into a probability distribution over the vocabulary, enabling the model to predict the next token in the sequence.
-D.2 Additional LoRA Models
-D.2.1 Incorporating Insignificant Pairs for LoRA Fine-tuning
-Similar to the embedding-based logit model, we have experimented with various LoRA configurations, such as further reducing the number of trainable parameters, increasing the dropout ratio, and applying larger weight decay in an attempt to mitigate over-fitting.
-However, as we observed, these modifications generally reduce performance on the training set without yielding significant improvements on the test set.
-In the following, we test whether incorporating more data, including insignificant pairs, can help improve the performance.
-Notice that by incorporating more significant pairs, we indeed observed improvement in test accuracy (see Figure A4), the observation of which holds for both embedding-based models and LoRA fine-tuned models.
-Here, we only present the results from fine-tuning Llama-3-8b using different sizes of training data, and the result from the embedding-based model remains the same.
-Recall that in §$\S$3.3.3, we use only significant pairs (with p-values smaller than or equal to 0.05, which includes 39,158 pairs) and split them into training and test sets.
-We continue to use the same significant test set for performance evaluation.
-However, for the training set, we obtain the data in the following manner: first, we remove the fixed test set from all pairs, leaving a total of 134,941 pairs.
-Then, given a ratio (e.g., 60%), we select these remaining pairs’ first 60% (ranked by ascending p-value) as the training set, see Figure A8 for a diagram of our dataset construction.
-Figure A8: Dataset Construction by Incorporating Both Significant and Insignificant Pairs in Llama-3-8b LoRA Fine-tuning
-We set the ratio to 20%, 40%, 60%, 80%, and 100%.
-For each ratio and size of the training set, we finetune the Llama-3-8b model and evaluate the performance on the fixed test set.
-Except for the training set, all other training configurations are the same as in §$\S$3.3.2.
-The performance is reported in Figure A9.
-The accuracy of the training set decreases as the ratio gets larger since we are incorporating more insignificant pairs.
-The accuracy on the test set maintains around 83%, which indicates that incorporating more training data does not necessarily lead to better performance, at least for the current dataset.
-Notice that the decreasing trend of accuracy on the training set in Figure A4 is less steep because the pairs in that training set are all significant.
-However, the main goal here is to confirm that more insignificant pairs cannot improve the performance for either LoRA or the embedding method.
-Figure A9: Performance of fine-tuning Llama-3-8b under varying training dataset sizes.
-D.2.2 DistilBERT LoRA Fine-Tuning
-We now discuss One difference between DistilBERT and Llama-3 is that DistilBERT is a bidirectional Transformer model, which means that its attention mechanism allows the model to learn the context from both directions, while Llama-3 is a unidirectional model.
-An illustration of the structure of the DistilBERT model combined with a classification head is shown in Figure A10.
-By default, DistilBERT can handle two headlines together as input.
-The format of the input is: [CLS] Headline 1 [EOS] Headline 2 [EOS].
-The [CLS] token is a special token used to get the output of the model for classification, and the output of the last Transformer layer corresponding to this token is used as the input to the classification head.
-The [EOS] token is used to separate the two sentences.
-Figure A10: The classification pipeline using the DistilBERT model.
-Headlines are first tokenized, and each token is converted into an embedding vector.
-For illustrative purposes, only two Transformer layers are shown.
-The figure highlights the use of bidirectional attention in DistilBERT, whereby each token can incorporate information from any token.
+D.2 LoRA Fine-Tuning Technique
+We illustrate LoRA in Figure A5.
+We use $W$ to represent any pre-trained weight matrix, which can be either $W_{Q}$, $W_{V}$, or $W_{K}$.
+Specifically, $W_{Q}$ is the weight matrix for the query vectors, $W_{K}$ is for the key vectors, and $W_{V}$ is for the value vectors in the attention mechanism of the Transformer architecture.
+These matrices store most of the information and knowledge learned by LLM, and are crucial for calculating the attention scores and subsequently determining the relevance of different tokens in the input sequence.
+For more technical background, we refer to Web Appendix §$\S$§D.1.
+We use $d\times d_{W}$ to represent the dimension of the weight matrix and $r$ to represent the rank of the low-rank matrices.
+We denote the modified weight matrices after fine-tuning as $W^{\prime}$, which can be further expressed as:
+$W^{\prime}=W+\Delta W,\quad\mathrm{with}\quad\Delta W=AB,\quad\mathrm{where}\quad A\in\mathbb{R}^{d\times r},\quad B\in\mathbb{R}^{r\times d_{W}}.$
+Figure A5: Illustration of LoRA Fine-Tuning.
+LoRA essentially introduces the low-rank matrices $A$ and $B$ to approximate the weight matrix change $\Delta W$ after the fine-tuning.
+These low-rank matrices are updated during the fine-tuning process, while the original weight matrix $W$ remains frozen.
+This strategy significantly reduces the number of trainable parameters from $d\times d_{W}$ to $r\times(d+d_{W})$ because the rank $r$ is typically set to a small value, such as 4 or 8, to balance the trade-off between model capacity and computational efficiency.
+D.3 Implementation Details
+We use exactly the same training and test sets as in §$\S$3.2, but the inputs here are the original headlines with the text information.
+We add a regression head to the original Llama-3-8b base model, which includes 32 Transformer layers, as depicted in Figure A6.
+Llama-3-8b employs a causal attention mask, where the attention is unidirectional—each token can only attend to preceding tokens in the sequence.
+Thus, the last token in the sequence, being informed by all prior tokens, is used for predicting the CTR.
+This embedding vector of the last token (which is 4096-dimensional), output by the last Transformer layer, is processed through a linear layer producing a one-dimensional scalar that represents the predicted CTR.
+Figure A6: The regression pipeline using the Llama model.
+Headlines are first tokenized and each token is converted into an embedding vector.
+For illustrative purposes, only two Transformer layers are shown, although the actual model comprises 32 layers for the 8 billion parameter configuration.
+The figure highlights the use of uni-directional attention in Llama, whereby each token can only incorporate information from preceding tokens.
 Connections between Transformer blocks represent the attention mechanism, and only a subset of these links are displayed for simplicity.
-The output from the first token (CLS token) is fed into a classification head, which computes the probability that one headline is more impactful than the other.
-Another difference is that DistilBERT is a smaller model compared to Llama-3-8b.
-DistilBERT has around 66 million parameters, which is only 0.94% of the parameters of Llama-3-8b.
-Due to the smaller model size, we are able to conduct full parameter updates during the fine-tuning process without using LoRA.
-We use the same learning rate scheduling, number of training epochs, and train-test set split as in the experiment detailed in §$\S$3.3.2.
-The accuracy of DistilBERT LoRA fine-tuning on the same test dataset is around 75%.
-Thus, we find that the newer Llama-3-8b model offers better performance upon fine-tuning.
-We refer interested readers to our code repository for further details on the fine-tuning process.
+The output from the last token is fed into a regression head, whose output represents the predicted CTR.
+During fine-tuning, our objective is to minimize the mean square error on the training set.
+We implement LoRA fine-tuning using the PEFT library developed by Hugging Face.
+We set the dimension of the low-rank matrices as $r=4$.
+Although LoRA can be applied to any subset of weight matrices in LLMs, applying LoRA to query weight matrices $W_{Q}$ and value weight matrices $W_{V}$ is a common practice and typically gives better performance (hu2021lora).
+Following this practice, in our experiment, we specifically apply LoRA to the $W_{Q}\in\mathbb{R}^{d\times d_{q}}$ and $W_{V}\in\mathbb{R}^{d\times d_{v}}$ matrices within the Transformer blocks of Llama-3-8b, where $d=4096,d_{q}=4096,d_{v}=1024$.
+To mitigate over-fitting issues, we also apply a relatively large dropout rate of 0.3 to these low-rank matrices.
+[^19] Furthermore, we enable 16-bit mixed precision training.
+Although 16-bit training slightly reduces numerical precision, it accelerates operations on GPUs, which enhances computational efficiency significantly.
+This mode makes the training process faster and consumes less memory compared to the default 32-bit training, especially for large-scale models like Llama-3-8b.
+With this LoRA configuration on Llama-3-8b, the total number of trainable parameters in the model is 1,703,936, which translates to only 0.023% of the total 7,506,636,800 parameters.
+We fine-tune the model for 10 epochs, meaning each sample (pair of headlines) in the training set is used 10 times in the training.
+A linearly decreasing learning rate scheduler is employed, starting from an initial rate of 2e-5 and decreasing linearly to zero by the end of the fine-tuning process.
+This strategy helps stabilize the training by gradually reducing the rate of updates to the weights, thus smoothing the convergence and reducing the risk of overshooting minima in the loss landscape.
+D.4 Performance on Different Train-test Splitting
+Recall that we observe the accuracy on the test set being slightly higher than the accuracy on the training set in Figure 4.
+This phenomenon is not specific to the LoRA fine-tuning method; the same effect is observed with embeddings.
+It is likely due to the training-test split resulting in unbalanced samples in different sets.
+To investigate this issue, we re-run experiments on different splits and report the training and test performance to validate this explanation.
+To save computation costs, instead of rerunning LoRA fine-tuning on different splits, we use embeddings for this sensitivity check, specifically the OpenAI-256E embedding with the Linear CTR prediction model, as used previously in §$\S$3.2.
+The split used in the paper corresponds to random seed 42.
+We reported the results using seeds 1, 2, 3, 4, and 5 in Table A3.
+The accuracy on the training and test sets is similar across these splits.
+Seed 1 Seed 2 Seed 3 Seed 4 Seed 5
+Training Set Accuracy 39.98% 40.51% 40.64% 40.50% 40.24%
+Test Set Accuracy 40.85% 40.85% 39.84% 40.91% 40.33%
+Table A3: Training and test set accuracy for different random seeds.
 Appendix E Cost Comparison of Pure-LLM-Based Methods
@@
 Method Cost Runing time Effort
-Prompt \$21.24 (10,000 GPT prompts) 121.5 mins for 10,000 GPT prompts (82 prompt requests per min) Low
-Embedding \$0.18 (text-embedding-3-large) 26 mins for embedding all 77,245 headlines (3,000 embedding requests per min) Moderate
+Prompt \$15.87 (gpt-4o with two demonstrations) 5 mins for 3,263 GPT prompts Low
+Embedding \$0.29 (text-embedding-3-large) 8 mins for embedding all 12,039 headlines in test set Moderate
 LoRA \$3 (Llama-3-8b) 2 hours (Llama-3-8b) High
@@
 Note that the prompt and embedding response time can vary based on the API usage tiers, with our results specifically derived from Tier 1.
-For the prompt-based method utilizing OpenAI API, the cost was \$1.13 for 5,000 GPT-3.5 prompts and \$20.11 for 5,000 GPT-4 prompts, which is approximately 20 times the cost of GPT-3.5.
-The total run-time for the entire experiment was 121.5 minutes, involving the testing of 10,000 prompts.
-For the embedding-based methods, obtaining text embeddings via the OpenAI API is extremely time-efficient and more economical compared to prompt-based methods.
-When using the embedding API, data transmission involves smaller, fixed-size vectors instead of variable-length text, which reduces data overhead and accelerates processing speed.
-We spent approximately \$0.18 to obtain text embeddings for all 77,245 headlines, while the prompt experiment with only 10,000 prompts cost around \$20.232323At the time of this study (May 2024), GPT-4 (turbo version) prompt costs \$10 per 1M input tokens and \$30 per 1M output tokens.
-The embeddings used in our study, text-embedding-3-large, only cost \$0.13 per 1M input tokens and there is no output token cost associated with embeddings.
-Our entire dataset comprises 77,245 headlines, with an average of 19 tokens per headline, resulting in a total embedding cost of approximately \$0.18.
-Additionally, processing embeddings is much faster (the maximum Tier 1 rate limit—3,000 requests per minute reached) compared to prompt-based methods (around 82 pairs per minute).
+Since we only make prompt inquiries for test data, to maintain a fair comparison, we report the runtime for the embedding method only for headlines in the test data, although we also need to obtain embeddings for headlines in the training data.
+For the prompt-based method utilizing the OpenAI API, the cost was \$2.16 for gpt-3.5-turbo-0125 and \$15.87 for gpt-4, on the test dataset with 3,263 prompts when using two demonstrations.
+With gpt-4 approximately 7.5 times the cost of GPT-3.5.
+This cost is calculated based on the length (the number of tokens) of the prompt.
+Runtime for 3,263 prompts is about 5 minutes when using parallelization.
+There is an inquiry rate/speed limit set by OpenAI that depends on the account tier.
+Using parallelism could significantly speed up the runtime until the inquiry rate limit is reached.
+For embedding-based methods, obtaining text embeddings via the OpenAI API is more economical than using prompt-based methods.
+At the time of this study (Nov 2024), a gpt-4 prompt costs \$2.5 per 1M input tokens and \$10 per 1M output tokens.
+The embeddings used in our study, text-embedding-3-large, cost only \$0.13 per 1M input tokens, with no output token cost associated with embeddings.
+The embedding cost of our entire dataset with 77,245 headlines is \$0.29.
+The reported monetary cost for gpt-4 is based on our prompt design, which includes not only the headline but also a description of our problem.
+Additionally, processing embeddings is much faster (the maximum Tier 1 rate limit—3,000 requests per minute reached) compared to prompt-based methods (500 requests per minute).
 This efficiency can lead to lower operational costs and easier integration into real-time applications.
@@
 To illustrate the benefit of reduced GPU requirements, we compare two commonly used GPU models for LLM training.
-Without LoRA, the NVIDIA H100 GPU, which has nearly the largest GPU memory (80 GB), is insufficient for fine-tuning.
+Without LoRA, the NVIDIA A100 GPU, which has nearly the largest GPU memory (80 GB), is insufficient for fine-tuning.
 In contrast, with LoRA, a more affordable NVIDIA V100 GPU with 32 GB of memory is sufficient.
-The price of an NVIDIA H100 GPU is ten times higher than an NVIDIA V100 GPU as of May 2024.
+The price of an NVIDIA A100 GPU is five times higher than an NVIDIA V100 GPU as of May 2024.
 Thus, using LoRA significantly reduces costs, making fine-tuning more accessible to researchers and practitioners.
@@
 In summary, we recommend using either embedding-based methods or fine-tuning methods, as they achieve similar performance (at least in our setting), superior to prompt-based methods, and are cost-effective.
-These methods offer different levels of transparency: prompt-based methods using OpenAI’s API are entirely black-box; embedding-based methods use OpenAI’s black-box embedding model combined with a user-controlled classification model; while fine-t tuning open-source models provides full transparency.
+These methods offer different levels of transparency: prompt-based methods using OpenAI’s API are entirely black-box; embedding-based methods use OpenAI’s black-box embedding model combined with a user-controlled CTR prediction model; while fine-tuning open-source models provides full transparency.
 Fine-tuning methods require more significant engineering effort to fine-tune and deploy an LLM than embedding-based methods.
-Appendix F LLM-based CTR Prediction Models
-We compare two approaches for CTR prediction: an OpenAI-embedding-based model and a LoRA fine-tuned Llama-3-8b model.
-In summary, we find the LoRA approach has better performance and we finally choose this model in the first step of our LOLA algorithm.
-The OpenAI-embedding-based CTR prediction model is similar to the one used for binary classification, as shown in Figure 3, with three key modifications: changing the loss function from log loss to MSE, using the text embedding of a single headline (3072 dimensions) instead of the concatenated embedding from headline pairs, and switching from binary output to continuous CTR prediction.
-To fine-tune the Llama-3-8b model for predicting CTR, we use a slightly different model structure compared to the one used for the pairwise comparison task shown in Figure 5.
-Specifically, as illustrated in Figure A11, we still use the Llama-3-8b model as the base LLM.
-The input is a single headline, and we apply a linear transformation to the output vector of the final token from the last Transformer layer.
-This transformation produces a single value, representing the predicted CTR.
-The model is trained to minimize the MSE between the real CTR and the predicted CTR.
-All other hyperparameters, such as batch size, number of training epochs, learning rate and its scheduling, and LoRA-related parameters, remain unchanged as in §$\S$3.3.
-Figure A11: Model architecture for CTR prediction based on the Llama-3-8b model.
-Figure A12 presents the numerical results, showing both the training loss (MSE) and accuracy throughout the training process.
-Accuracy is computed as follows: for each dataset (training or test), we use the fine-tuned model to predict the CTR for each headline.
-For each news article, if the headline with the highest predicted CTR matches the headline with the highest actual CTR, we consider this prediction correct.
-The accuracy is then calculated as the proportion of news articles for which the model’s predictions are correct.
-Figure A12: Performance of fine-tuning the Llama-3-8b model for CTR prediction.
-The accuracy of the OpenAI embedding-based method and the random guess method are also reported.
-For the random guess method, the winner headline is uniformly and randomly selected for each news article.
-As shown in Figure A12, the fine-tuned Llama-3-8b model significantly outperforms the OpenAI embedding-based method in terms of accuracy.
-Regarding the R-squared values, the embedding-based approach achieves 0.241, whereas the LoRA fine-tuned method achieves a higher R-squared value of 0.315.
-In contrast, for the pairwise comparison task, the fine-tuned Llama-3-8b model shows only a marginal improvement over the OpenAI embedding-based method.
-CTR prediction on all headlines is more challenging than pairwise comparison on significant pairs because the CTR output has higher uncertainty compared to a binary label.
-We believe this increased difficulty necessitates a more complex model, such as Llama-3-8b, which results in a larger performance gap between the fine-tuned Llama-3-8b and the OpenAI embedding-based method.
-Additionally, note that the accuracy on the test set is higher than on the training set.
-In our experiment, many news articles have headlines with very similar CTRs, making it more difficult to distinguish between them, which lowers the accuracy.
-This result may be due to having more such news articles in the training set and fewer in the test set.
-It is possible that with a different random split of the training and test sets, the accuracy of the training set could be higher than that of the test set.
+Appendix F Sensitivity Analysis for Hyperparameter $n^{\text{aux}}$ in LOLA
+In this part, we document the sensitivity analysis results for the hyperparameter $n^{\text{aux}}$.
+We begin by fixing the other hyperparameter $\alpha=0.08$, which has been fine-tuned for our setting, and focus solely on $n^{\text{aux}}$.
+The sensitivity results are visualized in Figure A7.
+To investigate its effect on the final average clicks, we select a wide range of values for $n^{\text{aux}}\in\{0,200,1000,1800,3000,5000\}$.
+Note that this experiment is conducted on the test dataset.
+When $n^{\text{aux}}=0$, LOLA essentially reduces to the standard UCB, which performs the worst.
+We observe that LOLA performs best when $n^{\text{aux}}=1000$.
+This is why we fine-tuned $n^{\text{aux}}$ within the range $\{600,800,1000,1200,1400\}$ around 1000 in our algorithm.
+This analysis highlights that the performance of LOLA is robust to the choice of $n^{\text{aux}}$, as $n^{\text{aux}}=1800$ still yields results close to those at $n^{\text{aux}}=1000$.
+Figure A7: Result of Sensitivity Analysis for Hyperparameter $n^{\text{aux}}$
 Appendix G Details of LOLA Variants and Extensions
@@
 In the implementation, we adopt the same data splitting rule as described in §$\S$4.3.
-Specifically, we used 70% data to fine-tune the Llama-3-8b CTR prediction model (see details in Web Appendix §$\S$ F), 10% data to fine-tune the $n^{\text{aux}}$ parameter, and tested the performance of the LLM-BAI algorithm on the remaining 20% test dataset.
+Specifically, we used 70% data to LoRA fine-tune the Llama-3-8b CTR prediction model, 10% data to fine-tune the $n^{\text{aux}}$ parameter, and tested the performance of the LLM-BAI algorithm on the remaining 20% test dataset.
 To fine-tune the $n^{\text{aux}}$ parameter, we run the LLM-BAI on the algorithm tuning dataset with different $n^{\text{aux}}$ values in $\{20,50,100,200,300,400,500,700,1000\}$, and choose the one with highest success rate in BAI.
@@
 We report the performance of LLM-BAI and the standard benchmarks in Figure A8.
-Note that the success rate for pure LLM here is higher than the accuracy (46.55%) reported in Figure A12, as we use a more lenient definition of success.
+Note that the success rate for pure LLM here is higher than the accuracy (46.86%) reported in Figure 4, as we use a more lenient definition of success.
 Our results indicate that our LOLA algorithm outperforms both the pure BAI method and the A/B test method in terms of success rate when using a similar number of pulls per test.
@@
 Note that the Y-axis captures the average clicks per test per period.
-For instance, if there is a test with two headlines conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
+For instance, if there is a test with two headlines receiving 1 and 2 clicks, conducted under $x=100$, then the average click per period in this test is calculated as $(1+2)/100=0.03$.
 The Y value is simply the average of this number $0.03$ over all tests.
@@
 However, there are two issues with the linearity assumption.
-First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over embedding-based classification.
+First, although we have shown that linear functions almost fully leverage text embeddings from OpenAI to achieve high accuracy (there is almost no difference compared to MLP), the embedding itself sacrifices a lot of information, as demonstrated by the significant outperformance of LoRA fine-tuning over the embedding-based approach.
 Second, once concatenated with user contexts, it is unclear whether the linear function is a good approximation of the rewards model (and this requires further numerical study).